### v1

# Addressing Blind Guessing: Calibration of Selection Bias in Multiple-Choice Question Answering by Video Language Models

Abstract
Evaluating Video Language Models (VLMs) is a challenging task. Due to its transparency, Multiple-Choice Question Answering (MCQA) is widely used to measure the performance of these models through accuracy. However, existing MCQA benchmarks fail to capture the full reasoning capabilities of VLMs due to selection bias, when models disproportionately favor certain answer options based on positional patterns observed during training. In this work, we conduct a comprehensive empirical analysis of several VLM architectures across major datasets designed to assess complex video-focused reasoning. We identify where the bias is most pronounced and demonstrate to what extent model responses reflect genuine understanding of video content and related questions, as opposed to reliance on arbitrary patterns or superficial cues, such as answer position. By decomposing the MCQA task and adapting fairness bias metrics to VLMs, we introduce a post-processing calibration technique BOLD to balance this bias. Our results show that reducing selection bias improves not only debiasing metrics but also overall model performance, including Accuracy and F1 Mean score. Our method, by suppressing "blind guessing", offers a more cost- and time-effective approach to mitigating selection bias compared to existing techniques. This study represents the first focused investigation of selection bias in video-to-text LLM-powered models.
Addressing Blind Guessing: Calibration of Selection Bias in Multiple-Choice Question Answering by Video Language Models
Olga Loginova University of Trento olga.loginova@unitn.it                        Oleksandr Bezrukov Gran Sasso Science Institute oleksandr.bezrukov@gssi.it                        Alexey Kravets University of Bath ak3095@bath.ac.uk
1 Introduction
Multiple-choice question answering (MCQA) is a convenient method to assess language models’ performance thanks to its simplicity and transparency. Each MCQA instance presents a question with several answer options, and the task is to select the correct one. However, large language models (LLMs) often exhibit selection bias, favoring certain answer positions Zheng et al. (2024a); Wang et al. (2024a). This bias undermines the fairness, reliability, and reasoning capabilities of the models.
While LLMs process text, video-language models (VLMs) handle sequences of frames to capture both spatial and temporal features. Recent advances align pre-training LLMs with video encoders to enhance video understanding Bordes et al. (2024). Although VLMs inherit much from LLMs, they face unique challenges, including higher computational costs and the need for effective spatial and temporal mapping for dynamic inputs.
Despite their backbone shared with LLMs, VLMs’ specific characteristics require separate scrutiny of selection bias. Notably, no research has yet focused on bias in video-language models.
Figure 1: Decomposition approach: we separate the three key components of MCQA, reorganize them into pairs by removing one at a time, and apply probability debiasing to the aggregated data.
We address this gap by proposing the first study of MCQA selection bias in video datasets, focusing on temporal reasoning, spatial object tracking, and camera behavior rather than general plot understanding. This approach highlights the specific challenges of VLMs.
Traditionally, selection bias is mitigated through shuffling Zheng et al. (2024a); Liu et al. (2023), where answer options are permutated multiple times to yield more impartial results. In contrast, we propose treating it similarly to the fairness bias, where models exhibit preference toward specific groups, such as gender or race. We adapt fairness bias metrics for VLMs via the decomposition approach showed in Figure  1.
To evaluate our approach, we select three VLMs with different architectures and apply post-processing bias calibration techniques to four datasets. Our method, which adjusts predictions without retraining, improves interpretability and fairness with reduced computational costs. Even though there is no direct correlation between debiasing and performance, our approach generally improves both bias metrics and accuracy, as well as the F1 mean score, signaling improved recall.
In summary, our contributions are:
  • A comprehensive study of selection bias in VLMs using 11 dataset modifications focused on video-specific challenges.
  • Adaptation of debiasing metrics to assess VLMs’ behavior on video MCQA benchmarks.
  • An efficient calibration technique that improves the reliability of current MCQA datasets and the VLM performance assessment.
Our work aims to reveal deeper reasoning issues in contemporary VLMs through refined MCQA benchmarks.
2 Related Works
Selection bias in LLMs has been extensively studied Pezeshkpour and Hruschka (2023); Zheng et al. (2023); Wang et al. (2024a); Li et al. (2024); Liusie et al. (2024); Wang et al. (2024b). Fewer works are dedicated to image-to-text vision LLMs Zong et al. (2024); Zhang et al. (2024); Xue et al. (2024).
Zheng et al. (2024a) conducted an empirical analysis of both position and token biases and proposed PriDe, a debiasing method for LLMs that separates prior bias from unbiased results using shuffling.
We also treat the selection bias as subtraction from observed results in order to obtain the debiased outputs, but we move away from shuffling and regard bias as a vector projected onto three planes. Thus, our calibration method aligns more closely with Zhang et al. (2024), who cut off the visual input for the image-to-text MCQA. Differently, we operate on video-to-text, and, going further, decompose the bias vector into several projections rather than one. This way we eliminate the unimodal bias highlighted by Zhang et al. (2023b).
Wang et al. (2024a) and Liu et al. (2023) resolve the selection bias by augmenting datasets with re-ordered and additional options, increasing number of question-answer pairs in the dataset. We, in contrast, manipulate answer options, as well as videos and questions, not as a solution but as a means of analysis: we identify where the bias is most pronounced and use these modifications as a shortcut for debiasing.
Lastly, Pezeshkpour and Hruschka (2023) propose calibrating LLM predictions by redistributing probabilities to empirically determined options based on the model’s uncertainty. However, we find this approach insufficiently robust, as, according to our analysis, the bias depends on both the model and the dataset, and there is no consistent correlation between the model confidence and the presence of bias.
3 Empirical Analysis
3.1 Datasets
We selected four video MCQA datasets that challenge VLMs’ temporal and spatial reasoning abilities. These datasets represent everyday situations and span videos from a few seconds to over an hour, most of which last up to one minute.
Table 1: Summary of Question-Answer pairs, Videos, and Number of Options for each dataset.
Dataset QA pairs Videos Options
NExT-QA 8564 1000 5
NExT-GQA 4962 971 5
STAR 7098 914 4
Video-MME 2700 900 4
Perception Test 7656 3926 3
NExT-QA and NExT-GQA
NExT-QA Xiao et al. (2021) is one of the first datasets designed to assess a VLM’s ability to capture causal, temporal, and descriptive topics through what/when/why/how and before/after questions. We evaluate the models on the test split.
Table 2: Inference Accuracy (Acc) and F1 Mean Score by Models and Datasets.
Model NExT-QA STAR Perception Test Video-MME
Acc F1 Mean Acc F1 Mean Acc F1 Mean Acc F1 Mean
Video-LLaMA 40.85 40.85 36.59 31.86 41.59 37.19 32.67 28.15
Video-LLaVA 49.96 49.81 34.71 31.83 40.73 35.69 34.22 30.99
SeViLA 63.78 63.88 46.28 46.14 45.30 45.11 39.85 39.82
NExT-GQA is a subset of the NExT-QA test split annotated with timestamp to locate the answer moment in the video.
STAR
The question-answer (QA) pairs in STAR Wu and Yu (2021) are generated using functional programs based on situation hyper-graphs. They cover four categories of situational reasoning: Interaction, Sequence, Prediction, and Feasibility. We use the publicly available validation split.
Perception Test
The Perception Test dataset Puatruaucean et al. (2023) tests the memory, abstraction, physics, and semantics reasoning abilities of the models. We use a 40% split of the correct-answer-annotated validation set to save the resources. Sampling is stratified to maintain the original balance of question types and the correct answer distribution.
Video-MME
Video-MME Fu et al. (2024) is a comprehensive MCQA evaluation benchmark with key distinctions: (a) video duration range from 11 seconds to 1 hour, (b) it covers 6 primary visual domains with 30 subfields, (c) includes subtitles, and (d) features meticulously crafted annotations. The released test split has an imbalanced option distribution, so we equalized it to properly assess option bias.
The amount of QA pairs, videos and answer options of each dataset is shown in Table 1. Additional details on the datasets including QA examples are given in Appendix  A.
Figure 2: Video-LLaMA option distribution for all settings of NeXT-QA, NExT-GQA, STAR, Perception Test and Video-MME. Correct ai, Correct ai Shuffled and All ai represent Correct Answer, Correct Answer with Shuffling and All Correct Answers in each option, respectively.
3.2 Models
Three models were selected to represent different approaches to uniting video and language modalities Bordes et al. (2024). Despite architectural differences, SeViLA, Video-LLaMA, and Video-LLaVA share key characteristics: they sample frames to extract visual features, either by keyframe selection or uniform sampling, and align visual features with textual representations using pre-trained LLMs. Each model uses pre-trained visual encoders and fine-tunes pre-trained components with video-text datasets.
Video-LLaMA
Video-LLaMA Zhang et al. (2023a) aligns a BLIP-2-based visual encoder Li et al. (2023) with LLaMA Touvron et al. (2023). It incorporates audio signals using Imagebind Girdhar et al. (2023). Both the video and audio Q-formers are trained on the Webvid-2M dataset Bain et al. (2021) to align the modalities. LLaMA is fine-tuned with visual instructional data.
Video-LLaVA
Video-LLaVA (Lin et al., 2023a) aligns video and LLMs in a shared representation. It uses the pre-trained CLIP ViT-L/14 vision encoder Radford et al. (2021) and the Vicuna language model Chiang et al. (2023), improving multimodal instruction following.
SeViLA
SeViLA Yu et al. (2023) adopts a self-chaining approach with BLIP-2 Li et al. (2023) for keyframe localization and question answering. It includes a Localizer, which uniformly samples 32 frames and selects the 4 most relevant key ones, and an Answerer that generates responses based on those frames. The Localizer and Answerer both have Flan-T5  Chung et al. (2022) for LLM.
We provide the Accuracy and F1 Mean Score results of the zero-shot inference on the aforementioned datasets in Table 2. Further details on the model settings and prompts can be found in Appendix B.
3.3 Experimental Settings
Eleven dataset modifications are used to examine the role of each component in MCQA benchmarks: video, question, and answer options.
Video Modifications
alter the video input:
  • Correct Frames: Only frames containing the answer are provided. This tests if reducing extraneous visual information improves performance. The modification applies to NExT-QA (NExT-GQA) and STAR due to the availability of timestamp annotations.
  • Empty Frames: Black frames are passed instead of video input. This tests the model’s response to the absence of visual information within the multimodal setting.
Question Modifications
adjust the questions in two ways:
  • Rephrased Questions: Llama3 Dubey et al. (2024) rephrases the original questions. Five variations are generated, and one is randomly selected. [^1] A manual check of 10% of each rephrased dataset confirmed that meaning was retained, with minimal impact on accuracy due to meaning loss. Notably, very few hallucinations were spotted in the NExT-QA rephrased copy.
  • Empty Questions: Following Balepur et al. (2024), we replace the question with an empty string to test if models can infer the question from answer choices and video. [^2] The previous research for LLMs has shown surprisingly high accuracy in this scenario. We investigate if the video modality contributes to even more meaningful predictions.
Answer Option Modifications
examine the role of answer:
  • Answer Shuffling: Options are randomly shuffled to test robustness against bias through changing the option order.
  • Correct Answer in Each Option: The correct answer is placed in a fixed position; its content is swapped with another option.
  • Correct Answer with Shuffling: The correct answer is fixed, and the remaining options are shuffled.
  • Additional Empty Option: A new empty option is added in the end to check for overfitting to specific options.
  • All Identical Answers: All options are identical testing for deviation from uniform distribution.
  • All Correct Answers: Every option is correct testing for deviations in a setting where any answer is valid.
  • Empty Answers: All options are empty to check the model’s ability to choose a non-existent answer.
To reduce token bias, we use less common option names a0, a1, a2, etc. instead of the conventional A, B, C, etc.
Modifications can also be categorized as accuracy-relevant (e.g., correct frames, shuffling, or rephrased questions) and accuracy-irrelevant (e.g., identical answers or empty parts). Further details, including accuracy results and confusion matrices, are provided in the Appendix C.
Figure 3: Video-LLaVA option distribution for all settings of NeXT-QA, NExT-GQA, STAR, Perception Test and Video-MME. Correct ai, Correct ai Shuffled and All ai represent Correct Answer, Correct Answer with Shuffling and All Correct Answers in each option, respectively.
3.4 Observations
Figures  2– 4 display the answer distributions of each model across options and datasets. High bars concentrated in a specific option indicate bias, while consistently low bars suggest the model disregards that option.
Across all models and datasets, shuffling answer options – either by shifting the correct option or random permutation, results in similar option distributions. This indicates that models rely heavily on answer position rather than content. However, when the correct answer is moved to less favored positions, those positions show increased selection rates, as seen in peaks for Correct ai and Correct ai Shuffled. Additionally, rephrased questions yield distributions similar to the default, regardless of correctness.
Using all correct frames did not significantly alter the distributions, suggesting that the models’ localization capabilities are adequate. Remarkably, in scenarios where all answer options are identical (accuracy-irrelevant), the selection distribution is not uniform.
Video-LLaMA
Video-LLaMA exposes a significant bias toward option a1 (Figure 2): when the correct answer is in this position, accuracy exceeds 70% on NExT-QA and Perception Test. This bias is more evident across STAR, Perception Test, and Video-MME, especially with empty frames.
The model tends to ignore options beyond the third one. In Perception Test, which has only three choices, bias against the third option is already noticeable.
Notably, with empty answers, option a2 receives an unexpectedly high number of selections, even surpassing a1 in Video-MME.
Video-LLaVA
Video-LLaVA’s performance (Figure 3) is similar to Video-LLaMA’s but shows an even stronger bias toward option a1, with higher-numbered options being largely ignored. Unlike Video-LLaMA, Video-LLaVA overwhelmingly favors a1 when answers are empty. When questions are empty, option a0 starts competing with a1, especially in the Video-MME dataset. In NeXT-QA, moving the correct answer among options leads to a preference for the option containing it.
SeViLA
Figure 4: SeViLA option distribution for all settings of NeXT-QA, NExT-GQA, STAR, Perception Test and Video-MME. Correct ai, Correct ai Shuffled and All ai represent Correct Answer, Correct Answer with Shuffling and All Correct Answers in each option, respectively.
SeViLA is unique in becoming distracted when an empty option is added. This suggests SeViLA may have been trained on answer distributions with a fixed number of options. It shows reduced accuracy for all correct frames in STAR (see Table  13), echoing findings by Goyal et al. (2016) and Lin et al. (2023b), where "blind" language-only models could answer image questions.
Nevertheless, SeViLA demonstrates the greatest robustness on NExT-QA and STAR. Its unstable option distributions occur only when all correct answers are identical (preferring a0) and with empty answers (preferring the last option). Figure 4 shows slightly higher saturation for these options across all datasets.
4 Methodology
The bias calibration is based on the following requirements to the MSQA benchmark:
  • Balanced options: All answer options are presented equally without any inherent bias (e.g., one answer option is not more likely than others just by how it is presented).
  • Independent options: Each option for a particular question is independent of others (e.g., there are no inherent relationships between them that would influence selection).
  • Equal difficulty: All questions are equally difficult, and no question skews results toward a particular option group.
  • No logically better or worse options except the correct one: This ensures that other options are plausible distractors, but only one option is objectively correct.
4.1 Evaluation Criteria
Our framework both detects bias and evaluates performance. Bias metrics assess fairness by evaluating the model’s decision making on consistency in treating all answer options, while performance metrics evaluate accuracy and reliability.
We monitor the standard deviations of the F1 score (F1_std) and the recall (Recall_std). High F1_std indicates inconsistent performance across options, signaling bias. High Recall_std shows the model is better at correctly identifying answers for certain options over others.
For option probabilities, we use the standard deviation of the symmetric Jensen-Shannon distance (JS_std) to detect inconsistencies in how the predicted probability distributions differ from the true distributions. This metric, bounded between 0 and 1, is more interpretable than Kullback-Leibler (KL) divergence.
By tracking these metrics, we address group fairness, ensuring the model treats all options equitably. Large standard deviations in any metric suggest bias, which we aim to minimize.
4.2 Debiasing Approach
Figure 5: Bias vector and its projections on the decomposition planes.
Every MCQA benchmark comprises essential components that make the answering task feasible. In the case of video-MCQA, these components include the video, the question, and the answer options. We construct decompositions where the task becomes impossible by design, as one of the key components is removed. This allows us to examine the projections of model’s behavioral bias on each decomposition plane as shown in Figure 5. In a fair and unbiased model, projecting the dataset by omitting one of the key components would result in a uniform distribution. However, as shown in Section 3, this is far from the case. In the decomposition planes, the bias is more clearly manifested, and since we expect uniformity, the bias becomes easier to detect and mitigate.
4.2.1 Decomposition
Consider a task $T$ that we wish to decompose into $n$ non-overlapping components $C_{1},C_{2},\dots,C_{n}$. This decomposition can be expressed as:
$T=\bigcup_{i=1}^{n}C_{i},\quad\text{where}\quad C_{i}\cap C_{j}=\emptyset\quad\text{for all}\quad i\neq j$
The task is considered ill-defined if it lacks one of the key components. Additionally, there is a well-defining property, or correctness, $P(T)$ of the task. If any component $C_{i}$ is removed, the property $P(T)$ becomes ill defined. We assume that all tasks are initially well-defined.
A model $M$ takes a task $T$ as input and outputs an option ID corresponding to the correct answer for a video-based question: $M(T)=ID_{opt}$.
For an ill-defined task, a model that is both reasonable and unbiased should uniformly select any option ID. Formally, if $P(T)$ is ‘ill-defined’: $M(T)=\text{RandomChoice}(\{ID_{opt}\})$.
The key insight is that if the model has bias on a given dataset, this bias is likely to be more ubiquitous, hence evident, in the ill-defined versions of the tasks. Thus, the model bias can be decomposed as follows: $Bias_{M}\supseteq\bigcup Bias_{M}|_{\hat{D}}$, where $\hat{D}$ represents an ill-defined copy of the dataset $D$.
Let $A$ be an attack on the well-defined task that makes this task defined insufficiently. Subsequently, there are three types of attacks that make the task ill-defined: $A|_{v=0}$ for all zeroed frames in the video, $A|_{q=0}$ for all questions as an empty string, and $A|_{o=0}$ for all options as IDs only.
We investigate the importance of the following components of the bias: $Bias_{M}|_{A|_{v=0}(D)}$, $Bias_{M}|_{A|_{q=0}(D)}$, $Bias_{M}|_{A|_{o=0}(D)}$, and how efficiently we can handle the general bias by balancing through these components.
4.2.2 Debiasing
[…]
$P_{o}(d_{i}|T)=\frac{1}{Z_{T}}P_{p}(d_{i}|T)\times P_{d}(d_{i}|T)$ (2)
where $Z_{T}$ is a normalization factor.
The baseline method BOLD (Bias Optimisation Leveraging Decomposition) estimates the prior bias as following:
$\tilde{P}_{p}(d_{i}|T)=softmax\left(\sum_{j}{P_{p}(d_{i}|A_{j}(T)})\right)$
First, we estimate the model bias on the dataset after attack A. We derive the $P_{p}$ knowing that $P_{d}$ is uniformly distributed from Equation  2:
$P_{p}(d_{i}\mid A(T))=P_{o}(d_{i}\mid A(T))$
Table 3: Comparison of BOLD and Weighted_BOLD bias mitigation approaches across models and datasets for performance and bias monitoring metrics with $k$ = 0.5. Green arrows indicate improvements: upward for Accuracy and F1_mean, and downward for standard deviation metrics. Red arrows represent deterioration, respectively.
Model Dataset Configuration Performance Metrics Bias Monitoring Metrics
Accuracy F1_mean Recall_std F1_std JS_std
NExT-QA BOLD 45.88 (2.43%) 42.15 (3.18%) 22.98 (3.53%) 15.53 (2.11%) 15.55 (3.31%)
Weighted_BOLD 45.91 (2.51%) 42.2 (3.29%) 22.83 (4.19%) 15.51 (2.2%) 15.56 (3.27%)
Video-LLaMA STAR BOLD 37.19 (1.82%) 33.02 (3.65%) 22.29 (7.5%) 14.55 (5.75%) 14.58 (4.66%)
Weighted_BOLD 37.34 (2.24%) 33.5 (5.16%) 21.13 (12.3%) 14.05 (8.98%) 14.14 (7.54%)
Perception Test BOLD 41.9 (1.24%) 38.68 (4.01%) 24.05 (11.11%) 9.87 (13.45%) 14.45 (10.01%)
Weighted_BOLD 42.06 (1.62%) 39.36 (5.86%) 21.8 (19.45%) 9.11 (20.06%) 13.02 (18.94%)
Video-MME BOLD 32.73 (2.13%) 29.23 (3.85%) 19.29 (6.45%) 11.96 (4.46%) 13.1 (4.73%)
Weighted_BOLD 32.2 (0.47%) 28.98 (2.94%) 17.65 (14.38%) 11.69 (6.56%) 12.64 (8.04%)
NExT-QA BOLD 51.81 (3.72%) 51.71 (3.82%) 13.27 (18.63%) 2.67 (18.42%) 4.58 (15.06%)
Weighted_BOLD 52.15 (4.39%) 52.04 (4.49%) 12.72 (22.02%) 2.62 (19.83%) 4.49 (16.83%)
Video-LLaVA STAR BOLD 37.21 (7.01%) 35.76 (12.37%) 18.28 (26.85%) 4.97 (27.15%) 4.54 (21.39%)
Weighted_BOLD 37.53 (7.94%) 36.18 (13.69%) 17.45 (30.15%) 4.91 (28.03%) 4.26 (26.24%)
Perception Test BOLD 41.62 (2.22%) 38.69 (8.4%) 21.75 (20.9%) 10.92 (25.97%) 3.78 (28.4%)
Weighted_BOLD 41.95 (3.02%) 39.46 (10.58%) 19.73 (28.26%) 10.24 (30.53%) 3.4 (35.51%)
Video-MME BOLD 34.7 (1.19%) 32.79 (5.81%) 18.19 (24.51%) 6.16 (23.63%) 3.84 (19.93%)
Weighted_BOLD 34.63 (0.97%) 32.97 (6.38%) 16.36 (32.07%) 6.01 (25.43%) 3.43 (28.5%)
NExT-QA BOLD 63.92 (0.02%) 63.89 (0.02%) 2.03 (5.47%) 1.18 (10.4%) 1.99 (0.17%)
Weighted_BOLD 63.93 (0.04%) 63.91 (0.04%) 1.99 (7.64%) 1.19 (9.83%) 1.99 (0.04%)
SeViLA STAR BOLD 46.22 (0.12%) 46.1 (0.08%) 4.13 (7.46%) 2.26 (1.78%) 2.1 (6.64%)
Weighted_BOLD 46.2 (0.18%) 46.08 (0.13%) 4.01 (10.17%) 2.2 (4.36%) 2.07 (8.2%)
Perception Test BOLD 45.32 (0.06%) 45.18 (0.16%) 5.18 (16.61%) 2.95 (2.43%) 1.3 (18.42%)
Weighted_BOLD 45.31 (0.03%) 45.2 (0.21%) 4.56 (26.58%) 2.83 (6.43%) 1.08 (31.92%)
Video-MME BOLD 40.19 (0.84%) 40.17 (0.88%) 4.11 (12.03%) 1.41 (16.19%) 0.73 (12.52%)
Weighted_BOLD 40.04 (0.46%) 40.03 (0.54%) 3.78 (19.14%) 1.44 (14.6%) 0.69 (17.22%)
To estimate the bias under $A|_{v=0}$, $A|_{q=0}$ and $A|_{o=0}$ attacks, we first compute $P_{p}(d_{i}|A|_{v=0}(T))$, $P_{p}(d_{i}|A|_{q=0}(T))$ and $P_{p}(d_{i}|A|_{o=0})$, and then the prior estimation, respectively: $\tilde{P}_{p}(d_{i}|T)=softmax[P_{p}(d_{i}|A|_{v=0}(T))+P_{p}(d_{i}|A|_{q=0}(T))+P_{p}(d_{i}|A|_{o=0})]$.
According to Zheng et al. (2024b), biases related to option IDs may transfer effectively across different samples and domains. Therefore, finding the global prior with $K$ samples, we can mitigate the bias for the entire dataset with minimal efforts.
From the dataset $D$, we take $K=k*||D||$ samples, denoted as $D_{k}$, where $k$ is adjusted coefficient based on estimation budget. Each sample in $D_{k}$ is subjected to the set of attacks $A$. For each, we calculate the sample-specific prior $P_{p}(d_{i}|A(T))$ by Equation 4 and estimate the adjusted prior $\tilde{P}_{p}(d_{i}|T)$ by Equation 3. The global prior, $\tilde{P}_{\text{p}}(d_{i})$, is then obtained by averaging the priors for each sample.
Ultimately, the posterior debiased through the global prior is calculated as follows:
$P_{d}(d_{i}|T)=softmax\left(\log P_{o}(d_{i}|T)-\log\tilde{P}_{p}(d_{i})\right)$
In the extended version, Weighted_BOLD, we use weights for $P_{p}(d_{i}|A(T))$:
$\tilde{P}_{p}(d_{i}|T)=softmax\left(\sum_{j}{w_{j}P_{p}(d_{i}|A_{j}(T)})\right)$
This allows for a more optimal direction in the latent space of the priors, as we can prioritize certain priors by significance and refine the prior estimation $\tilde{P}_{p}(d_{i}|T)$.
To gain the optimal weights, we use a 5-fold cross-validation procedure. In each fold, we optimize weights $\{w_{i}\}$ using the COBYLA algorithm Powell (1994), subject to the constraints $0\leq w_{i}\leq 1$ or $|w_{i}|\leq 1$. While we minimize the standard deviation of Recall_std across the options on the test set, we monitor the same bias metrics on the validation set. After completing the cross-validation we average the estimated global priors across the folds and apply this averaged global prior.
Both algorithms, BOLD and Weighted_BOLD, are outlined in Appendix D.
5 Results
We conducted tests with the $k$ sample parameter set to 0.25, 0.5, 0.75, and 1 (the entire dataset) in combination with positive and negative weights. The configuration with $k$=0.5 and $w_{i}\geq 0$ provided the optimal balance between sample size and performance. As shown in Table 3 with results for positive weights and $k$=0.5, [^3] Additional results for $k$ values of 0.25, 0.75, and 1 combined with different weights are provided in Appendix D. the overall significant improvement in bias metrics, including Recall_std, indicates that our method effectively reduces option disparity.
Weighted_BOLD generally outperforms BOLD because the weights capture critical priors in the latent space. Thus, they prove to be effective for fine-tuning bias correction within the decomposed priors.
Both BOLD and Weighted_BOLD consistently improve performance metrics for Video-LLaMA and Video-LLaVA. SeViLA generally exhibits greater resilience to bias calibration, showing minimal effect on performance on NExT-QA and marginal deterioration on STAR. This may be attributed to the model’s training for statistically close to uniform distribution of results over a fixed number of options. Despite decreases in Accuracy and F1 Mean on STAR, the reduction in Recall_std across all datasets for this model suggests that its debiased responses are more truthful, reflecting a fairer distribution of correct answers across options.
Video-LLaVA, which exhibits the most pronounced selection bias, benefits the most from bias reduction via prior decomposition across all datasets, achieving a maximum accuracy gain of 7.94% on STAR with the weighted configuration.
6 Conclusion
We performed the first comprehensive analysis on modified MCQA datasets to localize selection bias in specific options across VLMs of various architectures. Our findings reveal that while all models exhibit common behavioral patterns in their answer distributions, each model also has unique characteristics both across and within datasets. Despite these differences, the bias vector consistently manifests more prominently in the projections on the decomposed planes of video-question, video-answer, and question-answer interactions with one MCQA key component removed.
Using bias metrics to detect inconsistencies between the predicted and true probability distributions of the answers and calculating the prior bias, we proposed an effective method to mitigate this bias. Using a unique global prior for each dataset we reinforced the robustness and generalizability. Our BOLD bias removal approach not only enhances fairness, but also proves effective in improving performance metrics. Using weights in Weighted_BOLD further optimizes the performance.
This easily scalable, post-processing, automatic, and unsupervised method allows for a sharper focus on the true reasoning capabilities of the models. Our study contributes to making MCQA benchmarks more transparent and robust against the ordering of options, thereby delivering fairer and more meaningful evaluation results for VLMs.
Furthermore, elevating bias suppression to an abstract level, where tasks are systematically decomposed and key components are removed to identify where bias is most pronounced, can extend our approach to address other bias-related challenges in machine learning models.
7 Limitations
A primary limitation of our approach is the assumption that models rely on indirect cues when lacking sufficient direct information to determine an answer. Our conclusion—that bias can be decomposed across empty videos, questions, and answers—is based on empirical observations where bias is most pronounced in these settings. However, this decomposition may not be exhaustive; there could be more optimal directions in the latent space of decomposed priors. For instance, models often perform markedly different results compared to other settings when all answers are correct. A supervised debiasing method using true labels, such as applying COBYLA optimization directly, might also yield better results. Nonetheless, our goal was to find an easy-to-implement unsupervised solution without using correct answers.
We also assume the validity of the uniform distribution: Jensen-Shannon Divergence operates under the assumption that, in the absence of prior information, the distribution should be uniform. If the test MCQA data or task suggest an uneven distribution of answer options, this may lead to underestimating the significance of certain factors.
8 Ethics Statement
We based our experiments on existing datasets, ensuring compliance with their respective copyrights. Our research uses open-source VLMs, which, while offering accessibility and flexibility, also carry the inherent risks associated with open-ended text generation. Upon acceptance of the paper, we will release our code and dataset to the public for research and reproducibility.
Acknowledgements
The first author thanks Amazon Alexa for supporting her research through a generous donation. We are deeply grateful to Prof. Ravi Shekhar and Giovanni Bonetta for their invaluable contributions at different stages of the project. In particular, we extend special thanks to Prof. Ravi Shekhar for developing the data preparation scripts that handled the movement of correct answers across options and the stratification of Perception Test, as well as for his consistent guidance and insightful consultations throughout the project.
Appendix A Datasets
A.1 NExT-QA and NExT-GQA
The NExT-QA dataset, released in 2021, comprises videos sourced from the VidOR dataset Shang et al. (2019), with an average length of 35.73 seconds. The test split primarily contains causal questions (52.57%) addressing why and how, temporal reasoning questions (31.03%) about preceeding, following, or concurrent actions, and descriptive tasks (16.04%). Videos are manually annotated with questions and correct answers. Distractors are selected from 50 candidates based on a cosine similarity of less than 0.9 to the correct answer, ensuring they are close but not identical. This approach sometimes allows questions to be meaningfully answered without viewing the video (see Figures  68).
Since our models were released after 2021, they might have included NExT-QA in their training or fine-tuning data.
NExT-GQA, released in 2024 as an extension of NExT-QA, focuses on evaluating the localization abilities of VLMs. It includes manually annotated timestamp labels that are key for answering the questions. Descriptive questions are excluded because they pertain to global video content or have answers available throughout the video with no specific temporal grounding. We also exclude QA pairs with multiple annotated time intervals. Consequently, the test set of NExT-GQA comprises 56.69% causal questions and 43.31% temporal questions.
Figure 6: Example of a QA pair from NExT-QA in the Causal category. The correct answer is in the box.
Figure 7: Example of a QA pair from NExT-QA in the Temporal category. The correct answer is in the box.
Figure 8: Example of a QA pair from NExT-QA in the Descriptive category. The correct answer is in the box.
A.2 STAR
The STAR dataset is derived from the Action Genome dataset Ji et al. (2019), which itself is based on Charades Yuan et al. (2017). Charades comprises indoor activity videos with an average length of 30 seconds. Action Genome provides annotations of objects and their relationships within these videos.
Situation hypergraphs are used to extract abstract representations from the videos. Multiple types of questions are then designed using functional programs to cover different levels of difficulty in situated reasoning. There are four types of questions focusing on occurred facts, temporal order, future probability, and feasibility in specific situations:
  • Sequence (50.53%): What did the person do before/after …?”
  • Interaction (33.78%): “What did a person do …?”
  • Prediction (8.79%): “What will the person do next with …?”
  • Feasibility (6.9%): “What is the person able to do?” or “Which object is possible to be …?”
All QA pairs are annotated with a timestamp label. To avoid answer distribution bias, each question type is balanced through resampling. To prevent reasoning shortcuts due to frequent action or entity co-occurrences and to ensure models cannot easily answer without genuine understanding, the compositionality of verbs and nouns is controlled.
Examples are shown in Figures  912.
Figure 9: Example of a QA pair from STAR in the Sequence category. The correct answer is in the box.
Figure 10: Example of a QA pair from STAR in the Interaction category. The correct answer is in the box.
Figure 11: Example of a QA pair from STAR in the Prediction category. The correct answer is in the box.
Figure 12: Example of a QA pair from STAR in the Feasibility category. The correct answer is in the box.
A.3 Perception Test
The Perception Test dataset consists of four categories designed to evaluate VLMs on different types of reasoning:
  • Physics (36.26%)
  • Semantics (30.54%)
  • Abstraction (29.28%)
  • Memory (3.92%)
These categories test descriptive, explanatory, predictive, and counterfactual reasoning abilities. The video scripts are based on human perception screening tests, with an average length of 23 seconds. All answers are crowd-sourced. Distractors are both manually created and automatically generated, particularly for standard closed yes/no questions.
The validation split we use contains approximately 15% closed questions, which poses an additional challenge for models that tend to exhibit affirmation bias Pezeshkpour and Hruschka (2023). Notably, for closed questions in this dataset, Video-LLaMA and Video-LLaVA often respond with "yes" or "no" rather than selecting an answer option.
Examples of QA pairs from the Perception Test are shown in Figures  1316.
Figure 13: Example of a QA pair from Perception Test in the Descriptive Physics category. The correct answer is in the box.
Figure 14: Example of a QA pair from Perception Test in the Descriptive Semantics category. The correct answer is in the box.
Figure 15: Example of a QA pair from Perception Test in the Counterfactual Abstraction category. The correct answer is in the box.
Figure 16: Example of a QA pair from Perception Test in the Explanatory Memory category. The correct answer is in the box.
A.4 Video-MME
Video-MME was developed to address the lack of diversity in video MCQA benchmarks. It consists of 900 YouTube videos ranging from 11 seconds to 1 hour in length. The dataset includes 12 task types covering questions on temporal and spatial perception, object and action reasoning, OCR, and counting problems. All QA pairs and distractors are manually created with three questions per video. Additionally, the dataset features videos in languages other than English.
Examples are shown in Figures  1720.
Figure 17: Example of a QA pair from Video-MME in the Temporal Reasoning category. The correct answer is in the box.
Figure 18: Example of a QA pair from Video-MME in the OCR Problems category. The correct answer is in the box.
Figure 19: Example of a QA pair from Video-MME in the Counting Problem category. The correct answer is in the box.
Figure 20: Example of a QA pair from Video-MME in the Spatial Perception category. The correct answer is in the box.
Appendix B Models’ Settings and Prompts
B.1 Video-LLaMA
Video-LLaMA is designed as a conversational model and was likely not specifically trained on MCQA, but rather on open-ended VQA datasets to generate captions, descriptions, and explanatory answers. Consequently, throughout our experiments with different prompt options, it tends to engage in dialogue: it repeats the question, responds with the general phrase "Hello, how can I help you today?", describes the video content, or even asks follow-up questions such as "What do you think about the video?" rather than providing a direct answer from a predefined list of options or option IDs. [^4] It is noteworthy that the model often uses the following template to explain its answer choice: ”The ID of the correct answer: $a_{i}$. Explanation: …” Making it elicit a clear and concise answer option is quite challenging.
To address this issue, we experimented with various prompts, temperature settings, and token limits, and developed the following strategy:
System prompt:
Analyse the video, choose the correct answer in multiple choice question answering. Give the ID of the correct answer option from the given list of options. Be concise.
User prompt:
<question> Options: <options>. The ID of the correct answer:
We set the number of beams to 2 and the temperature to 0.8. Additionally, we allowed the model up to 30 attempts to produce an answer that includes an answer ID. We used the 7B parameter model due to resource constraints. For answering, the model selects eight uniformly sampled frames.
B.2 Video-LLaVA
Video-LLaVA effectively delivers concise answers as just a specific options based on a short, general prompt. However, in some cases, the model outputs as an answer option only "A" or responds with "yes" or "no" to a closed question.
For this model, we used the following prompt:
<video_tag> Question: <question> Options: <options> The ID of the correct answer is: Respond with only the correct answer ID. ASSISTANT:
Notably, on the completely unfamiliar datasets Perception Test and Video-MME adding an empty option increases the number of responses that cannot be parsed as an answer ID (although this increase is statistically insignificant). Combined with the observation that the model tends to disregard options presented later in the prompt, this suggests that the length of the question affects the model’s performance, and tokens closer to the end of the prompt have a lower probability.
The model also selects eight frames from the video by uniform sampling. To conserve computational resources, we used Video-LLaVA-7B with half-precision set by Torch’s autocast. The temperature is equal to 1.
B.3 SeViLA
SeViLA does not have any unanswered or unparsable responses because it was specifically tailored for the multiple-choice VQA task: internally, it maps any answer option IDs to {A, B, C, D, E}. However, it is limited to observing at most six answer options. In a side experiment, we shifted all answers to options G, H, I, J, K leaving the first six options empty, and obtained identical results to our other experimental setting, Empty Answers. This indicates that the model is completely blind to options beyond F and is likely overfitted to the options {A, B, C, D, E}.
Furthermore, the shape of the confusion matrices for NExT-QA and STAR in Figures 21 and 22 shows an unusually accurate understanding of these datasets, unlike those for Perception Test and Video-MME. Since NExT-QA and STAR were released before SeViLA, the model might have been fine-tuned on them. On the other hand, this suggests that the positional bias in this model is much less pronounced.
Another vulnerability of SeViLA is that it uses argmax calculated on logit probabilities. However, logits without numerical stabilization can produce NaN values when activated via softmax. In such cases, according to argmax, the model’s answer defaults to $a_{0}$. Among the settings relevant for debiasing, such behavior was observed in 22 cases on NExT-QA.
For this model, we used the original prompts: Localizer prompt (when the model selects 32 frames):
Does the information within the frame provide necessary details to accurately answer the given question?
Answerer prompt (when the model selects the answer based on 4 key frames):
Considering information in frames, select the correct answer from the options.
The confusion matrices for all models in the default (unmodified) setting in Figures  2123 provide a quick insight into positional bias: darker blue cells indicate a greater accumulation of answers.
Figure 21: All Models’ Confusion Matrices for NeXT-QA.
Figure 22: All Models’ Confusion Matrices for STAR.
Figure 23: All Models’ Confusion Matrices for Video-MME.
Figure 24: All Models’ Confusion Matrices for Perception Test.
Appendix C Models’ Performance on All Settings
Tables 4 through 15 present the distribution of selections across answer options and the accuracy (where applicable) for all settings, along with the target distributions. Notably, as soon as the correct answer aligns with the model’s biased option, the performance boosts.
Table 4: Video-LLaMA Performance on All Settings of NExT-QA (including NExT-GQA). Each cell shows the count of selections for each option (a0a5), along with the percentage in parentheses. The Accuracy column represents the number of correct answers where applicable. The N/A column indicates the number of cases where the model did not output an answer option after 30 attempts.
Setting a0 a1 a2 a3 a4 a5 N/A Accuracy
Target 1,721 (20.10%) 1,671 (19.51%) 1,767 (20.63%) 1,719 (20.07%) 1,686 (19.69%) 0 (0.00%)
Default 1,430 (16.70%) 3,285 (38.37%) 2,727 (31.85%) 1,002 (11.70%) 117 (1.37%) 3 (0.04%) 3,837 (44.82%)
Answer Shuffling 1,443 (16.97%) 3,295 (38.75%) 2,654 (31.21%) 993 (11.68%) 119 (1.40%) 60 (0.70%) 3,799 (44.67%)
Correct Frames 869 (17.61%) 1,836 (37.20%) 1,588 (32.18%) 587 (11.89%) 55 (1.11%) 27 (0.54%) 2,155 (43.67%)
Rephrased Questions 1,305 (15.36%) 3,270 (38.50%) 2,704 (31.83%) 1,062 (12.50%) 153 (1.80%) 70 (0.82%) 3,762 (44.29%)
Additional Empty Option 1,432 (17.17%) 2,987 (35.82%) 2,067 (24.78%) 1,090 (13.07%) 725 (8.69%) 39 (0.47%) 224 (2.62%) 4,086 (48.99%)
All a1 3,537 (42.71%) 3,891 (46.98%) 737 (8.90%) 32 (0.39%) 85 (1.03%) 282 (3.29%)
All a2 3,614 (43.82%) 3,780 (45.83%) 748 (9.07%) 32 (0.39%) 73 (0.89%) 317 (3.70%)
All a3 3,585 (43.43%) 3,831 (46.41%) 722 (8.75%) 33 (0.40%) 83 (1.01%) 310 (3.62%)
All a4 3,590 (43.35%) 3,837 (46.33%) 733 (8.85%) 34 (0.41%) 87 (1.05%) 283 (3.30%)
Correct a0 4,136 (48.98%) 2,303 (27.27%) 1,489 (17.63%) 453 (5.36%) 64 (0.76%) 119 (1.39%) 4,136 (48.98%)
Correct a0 Shuffled 4,136 (49.00%) 2,304 (27.30%) 1,525 (18.07%) 415 (4.92%) 61 (0.72%) 123 (1.44%) 4,136 (49.00%)
Correct a1 730 (8.61%) 6,323 (74.61%) 1,102 (13.00%) 292 (3.45%) 28 (0.33%) 89 (1.04%) 6,323 (74.61%)
Correct a1 Shuffled 704 (8.30%) 6,348 (74.88%) 1,098 (12.95%) 297 (3.50%) 31 (0.37%) 86 (1.00%) 6,348 (74.88%)
Correct a2 622 (7.32%) 2,325 (27.35%) 5,240 (61.64%) 280 (3.29%) 34 (0.40%) 63 (0.74%) 5,240 (61.64%)
Correct a2 Shuffled 583 (6.88%) 2,356 (27.81%) 5,240 (61.84%) 254 (3.00%) 40 (0.47%) 91 (1.06%) 5,240 (61.84%)
Correct a3 762 (8.99%) 2,315 (27.32%) 2,321 (27.39%) 3,029 (35.74%) 47 (0.55%) 90 (1.05%) 3,029 (35.74%)
Correct a3 Shuffled 766 (9.07%) 2,335 (27.65%) 2,235 (26.47%) 3,049 (36.11%) 59 (0.70%) 120 (1.40%) 3,049 (36.11%)
Correct a4 1,194 (14.11%) 3,014 (35.63%) 2,857 (33.77%) 909 (10.74%) 486 (5.74%) 104 (1.21%) 486 (5.74%)
Correct a4 Shuffled 1,164 (13.77%) 2,912 (34.45%) 2,949 (34.88%) 912 (10.79%) 517 (6.12%) 110 (1.28%) 517 (6.12%)
All Correct Answers 5,247 (63.17%) 2,730 (32.87%) 295 (3.55%) 9 (0.11%) 25 (0.30%) 258 (3.01%)
Empty Frames 1,092 (12.77%) 3,494 (40.87%) 3,344 (39.11%) 544 (6.36%) 76 (0.89%) 14 (0.16%) 2,620 (30.64%)
Empty Questions 1,686 (20.31%) 3,975 (47.89%) 1,676 (20.19%) 786 (9.47%) 178 (2.14%) 263 (3.07%) 3,082 (37.13%)
Empty Answers 987 (11.66%) 4,543 (53.68%) 2,837 (33.52%) 37 (0.44%) 59 (0.70%) 101 (1.18%)
Table 5: Video-LLaMA Performance on All Settings of STAR. Each cell shows the count of selections for each option (a0a4), along with the percentage in parentheses. The Accuracy column represents the number of correct answers where applicable. The N/A column indicates the number of cases where the model did not output an answer option after 30 attempts.
Setting a0 a1 a2 a3 a4 N/A Accuracy
Target 1,755 (24.73%) 1,758 (24.77%) 1,742 (24.54%) 1,843 (25.97%) 0 (0.00%)
Default 1,357 (19.13%) 3,645 (51.38%) 1,982 (27.94%) 110 (1.55%) 4 (0.06%) 2,596 (36.59%)
Answer Shuffling 1,255 (18.72%) 3,458 (51.57%) 1,893 (28.23%) 99 (1.48%) 393 (5.54%) 2,467 (36.79%)
Correct Frames 1,322 (19.43%) 3,455 (50.79%) 1,899 (27.91%) 127 (1.87%) 295 (4.16%) 2,524 (37.10%)
Rephrased Questions 1,376 (20.03%) 3,462 (50.39%) 1,891 (27.53%) 141 (2.05%) 228 (3.21%) 2,520 (36.68%)
Additional Empty Option 1,314 (20.59%) 2,892 (45.31%) 1,494 (23.41%) 630 (9.87%) 52 (0.81%) 716 (10.09%) 2,528 (39.61%)
All a0 2,484 (38.04%) 3,949 (60.47%) 88 (1.35%) 9 (0.14%) 568 (8.00%)
All a1 2,695 (39.68%) 3,976 (58.54%) 104 (1.53%) 17 (0.25%) 306 (4.31%)
All a2 2,711 (39.93%) 3,957 (58.28%) 113 (1.66%) 9 (0.13%) 308 (4.34%)
All a3 2,512 (38.52%) 3,909 (59.94%) 94 (1.44%) 7 (0.11%) 576 (8.11%)
Correct a0 2,340 (34.53%) 2,996 (44.21%) 1,378 (20.33%) 63 (0.93%) 321 (4.52%) 2,340 (34.53%)
Correct a0 Shuffled 2,313 (34.36%) 2,960 (43.97%) 1,392 (20.68%) 67 (1.00%) 366 (5.16%) 2,313 (34.36%)
Correct a1 966 (14.16%) 4,677 (68.54%) 1,124 (16.47%) 57 (0.84%) 274 (3.86%) 4,677 (68.54%)
Correct a1 Shuffled 985 (14.48%) 4,622 (67.93%) 1,136 (16.70%) 61 (0.90%) 294 (4.14%) 4,622 (67.93%)
Correct a2 918 (13.53%) 2,913 (42.93%) 2,898 (42.71%) 56 (0.83%) 313 (4.41%) 2,898 (42.71%)
Correct a2 Shuffled 844 (12.57%) 2,908 (43.32%) 2,901 (43.21%) 60 (0.89%) 385 (5.42%) 2,901 (43.21%)
Correct a3 1,226 (18.03%) 3,298 (48.50%) 2,002 (29.44%) 274 (4.03%) 298 (4.20%) 274 (4.03%)
Correct a3 Shuffled 1,153 (17.21%) 3,190 (47.60%) 2,127 (31.74%) 231 (3.45%) 397 (5.59%) 231 (3.45%)
All Correct Answers 3,129 (47.60%) 3,376 (51.35%) 63 (0.96%) 6 (0.09%) 524 (7.38%)
Empty Frames 1,323 (18.64%) 3,721 (52.42%) 2,047 (28.84%) 7 (0.10%) 0 (0.00%) 1,854 (26.12%)
Empty Questions 2,365 (34.84%) 2,973 (43.80%) 1,275 (18.78%) 175 (2.58%) 310 (4.37%) 2,427 (35.75%)
Empty Answers 1,121 (15.84%) 3,093 (43.72%) 2,778 (39.27%) 83 (1.17%) 23 (0.32%)
Table 6: Video-LLaMA Performance on All Settings of Perception Test. Each cell shows the count of selections for each option (a0a3), along with the percentage in parentheses. The Accuracy column represents the number of correct answers where applicable. The N/A column indicates the number of cases where the model did not output an answer option after 30 attempts.
Setting a0 a1 a2 a3 N/A Accuracy
Target 2,549 (33.29%) 2,561 (33.45%) 2,546 (33.25%) 0 (0.00%)
Default 1,546 (20.34%) 5,281 (69.47%) 775 (10.19%) 54 (0.71%) 3,162 (41.59%)
Answer Shuffling 1,593 (21.10%) 5,165 (68.42%) 791 (10.48%) 107 (1.40%) 3,146 (41.67%)
Rephrased Questions 1,652 (21.83%) 5,004 (66.13%) 911 (12.04%) 89 (1.16%) 3,161 (41.77%)
Additional Empty Option 1,448 (19.56%) 4,356 (58.86%) 1,549 (20.93%) 48 (0.65%) 255 (3.33%) 3,004 (40.59%)
All a0 2,768 (36.60%) 4,414 (58.36%) 381 (5.04%) 93 (1.21%)
All a1 2,842 (37.56%) 4,334 (57.28%) 390 (5.15%) 90 (1.18%)
All a2 2,806 (37.15%) 4,326 (57.28%) 421 (5.57%) 103 (1.35%)
Correct a0 2,245 (29.76%) 4,675 (61.98%) 623 (8.26%) 113 (1.48%) 2,245 (29.76%)
Correct a0 Shuffled 2,271 (30.12%) 4,654 (61.73%) 614 (8.14%) 117 (1.53%) 2,271 (30.12%)
Correct a1 1,204 (15.95%) 5,833 (77.29%) 510 (6.76%) 109 (1.42%) 5,833 (77.29%)
Correct a1 Shuffled 1,233 (16.32%) 5,820 (77.06%) 500 (6.62%) 103 (1.35%) 5,820 (77.06%)
Correct a2 1,239 (16.40%) 5,077 (67.18%) 1,241 (16.42%) 99 (1.29%) 1,241 (16.42%)
Correct a2 Shuffled 1,235 (16.36%) 5,082 (67.30%) 1,234 (16.34%) 105 (1.37%) 1,234 (16.34%)
All Correct Answers 3,330 (43.87%) 4,053 (53.40%) 207 (2.73%) 66 (0.86%)
Empty Frames 990 (12.93%) 5,641 (73.68%) 1,025 (13.39%) 0 (0.00%) 3,005 (39.25%)
Empty Questions 2,564 (34.61%) 3,707 (50.03%) 1,138 (15.36%) 247 (3.23%) 3,021 (40.77%)
Empty Answers 2,969 (39.11%) 2,876 (37.88%) 1,747 (23.01%) 64 (0.84%)
Table 7: Video-LLaMA Performance on All Settings of Video-MME. Each cell shows the count of selections for each option (a0a4), along with the percentage in parentheses. The Accuracy column represents the number of correct answers where applicable. The N/A column indicates the number of cases where the model did not output an answer option after 30 attempts.
Setting a0 a1 a2 a3 a4 N/A Accuracy
Target 675 (25.00%) 675 (25.00%) 675 (25.00%) 675 (25.00%) 0 (0.00%)
Default 474 (17.92%) 1,344 (50.81%) 757 (28.62%) 70 (2.65%) 55 (2.04%) 864 (32.67%)
Answer Shuffling 451 (18.11%) 1,242 (49.86%) 733 (29.43%) 65 (2.61%) 209 (7.74%) 785 (31.51%)
Rephrased Questions 1,367 (50.65%) 979 (36.27%) 198 (7.34%) 155 (5.74%) 1 (0.04%) 873 (32.35%)
Additional Empty Option 391 (17.29%) 1,040 (45.98%) 624 (27.59%) 178 (7.87%) 29 (1.28%) 438 (16.22%) 746 (32.98%)
All a0 1,042 (42.81%) 1,244 (51.11%) 120 (4.93%) 28 (1.15%) 266 (9.85%)
All a1 1,027 (42.00%) 1,295 (52.97%) 91 (3.72%) 32 (1.31%) 255 (9.44%)
All a2 1,061 (43.36%) 1,232 (50.35%) 116 (4.74%) 38 (1.55%) 253 (9.37%)
All a3 1,001 (41.06%) 1,223 (50.16%) 149 (6.11%) 65 (2.67%) 262 (9.70%)
Correct a0 678 (27.19%) 1,141 (45.75%) 624 (25.02%) 51 (2.04%) 206 (7.63%) 678 (27.19%)
Correct a0 Shuffled 699 (27.90%) 1,085 (43.31%) 664 (26.51%) 57 (2.28%) 195 (7.22%) 699 (27.90%)
Correct a1 376 (14.98%) 1,532 (61.04%) 547 (21.79%) 55 (2.19%) 190 (7.04%) 1,532 (61.04%)
Correct a1 Shuffled 344 (13.78%) 1,519 (60.83%) 584 (23.39%) 50 (2.00%) 203 (7.52%) 1,519 (60.83%)
Correct a2 366 (14.63%) 1,147 (45.86%) 948 (37.90%) 40 (1.60%) 199 (7.37%) 948 (37.90%)
Correct a2 Shuffled 355 (14.27%) 1,144 (45.98%) 942 (37.86%) 47 (1.89%) 212 (7.85%) 942 (37.86%)
Correct a3 408 (16.47%) 1,178 (47.56%) 777 (31.37%) 114 (4.60%) 223 (8.26%) 114 (4.60%)
Correct a3 Shuffled 386 (15.55%) 1,199 (48.31%) 786 (31.67%) 111 (4.47%) 218 (8.07%) 111 (4.47%)
All Correct Answers 1,195 (46.46%) 1,226 (47.67%) 114 (4.43%) 37 (1.44%) 128 (4.74%)
Empty Frames 188 (6.98%) 1,612 (59.81%) 843 (31.28%) 52 (1.93%) 5 (0.19%) 776 (28.79%)
Empty Questions 753 (29.89%) 1,180 (46.84%) 527 (20.92%) 59 (2.34%) 181 (6.70%) 798 (31.68%)
Empty Answers 440 (17.07%) 725 (28.13%) 1,352 (52.46%) 60 (2.33%) 123 (4.56%)
Table 8: Video-LLaVA Performance on All the Settings of NExT-QA (including NExT-GQA). Each cell shows the count of selections for each option (a0a5) along with the percentage in parentheses. Accuracy represents the number of correct answers where applicable.
Setting a0 a1 a2 a3 a4 a5 N/A Accuracy
Target 1,721 (20.10%) 1,671 (19.51%) 1,767 (20.63%) 1,719 (20.07%) 1,686 (19.69%) 0 (0.00%)
Default 874 (10.21%) 3,733 (43.59%) 890 (10.39%) 1,505 (17.57%) 1,562 (18.24%) 0 (0.00%) 4,279 (49.96%)
Answer Shuffling 1,507 (17.60%) 4,366 (50.98%) 1,630 (19.03%) 1,061 (12.39%) 0 (0.00%) 0 (0.00%) 3,653 (42.66%)
Correct Frames 440 (8.87%) 2,194 (44.22%) 482 (9.71%) 791 (15.94%) 1,055 (21.26%) 0 (0.00%) 2,407 (48.51%)
Rephrased Questions 836 (9.79%) 3,531 (41.36%) 1,054 (12.34%) 1,613 (18.89%) 1,504 (17.62%) 26 (0.30%) 4,287 (50.21%)
Additional Empty Option 418 (4.88%) 3,449 (40.27%) 847 (9.89%) 1,208 (14.11%) 2,641 (30.84%) 1 (0.01%) 0 (0.00%) 4,264 (49.79%)
All a0 1,232 (14.39%) 5,839 (68.18%) 9 (0.11%) 175 (2.04%) 1,309 (15.28%) 0 (0.00%)
All a1 1,254 (14.64%) 5,839 (68.18%) 9 (0.11%) 187 (2.18%) 1,275 (14.89%) 0 (0.00%)
All a2 1,238 (14.46%) 5,755 (67.21%) 12 (0.14%) 158 (1.85%) 1,400 (16.35%) 1 (0.01%)
All a3 966 (11.28%) 7,063 (82.47%) 14 (0.16%) 178 (2.08%) 343 (4.01%) 0 (0.00%)
All a4 996 (11.63%) 7,021 (81.98%) 20 (0.23%) 176 (2.06%) 351 (4.10%) 0 (0.00%)
Correct a0 2,966 (34.87%) 3,054 (35.91%) 484 (5.69%) 975 (11.46%) 1,026 (12.06%) 59 (0.69%) 2,966 (34.87%)
Correct a0 Shuffled 2,702 (31.55%) 3,680 (42.97%) 341 (3.98%) 759 (8.86%) 1,082 (12.63%) 0 (0.00%) 2,702 (31.55%)
Correct a1 251 (2.93%) 6,993 (81.66%) 209 (2.44%) 431 (5.03%) 680 (7.94%) 0 (0.00%) 6,993 (81.66%)
Correct a1 Shuffled 270 (3.15%) 6,981 (81.52%) 203 (2.37%) 424 (4.95%) 686 (8.01%) 0 (0.00%) 6,981 (81.52%)
Correct a2 274 (3.20%) 3,619 (42.26%) 3,032 (35.41%) 669 (7.81%) 969 (11.32%) 1 (0.01%) 3,032 (35.41%)
Correct a2 Shuffled 282 (3.29%) 3,561 (41.58%) 3,056 (35.68%) 707 (8.26%) 958 (11.19%) 0 (0.00%) 3,056 (35.68%)
Correct a3 258 (3.01%) 2,765 (32.29%) 319 (3.73%) 4,161 (48.59%) 1,060 (12.38%) 1 (0.01%) 4,161 (48.59%)
Correct a3 Shuffled 278 (3.25%) 2,746 (32.06%) 271 (3.16%) 4,182 (48.83%) 1,087 (12.69%) 0 (0.00%) 4,182 (48.83%)
Correct a4 324 (3.78%) 2,374 (27.72%) 300 (3.50%) 557 (6.50%) 5,009 (58.49%) 0 (0.00%) 5,009 (58.49%)
Correct a4 Shuffled 336 (3.92%) 2,305 (26.91%) 313 (3.65%) 579 (6.76%) 5,031 (58.75%) 0 (0.00%) 5,031 (58.75%)
All Correct Answers 1,536 (17.94%) 6,531 (76.27%) 4 (0.05%) 59 (0.69%) 433 (5.06%) 1 (0.01%)
Empty Frames 634 (7.40%) 4,625 (54.01%) 740 (8.64%) 1,499 (17.50%) 1,066 (12.45%) 0 (0.00%) 3,025 (35.32%)
Empty Questions 1,976 (23.07%) 4,925 (57.51%) 228 (2.66%) 696 (8.13%) 739 (8.63%) 0 (0.00%) 2,979 (34.79%)
Empty Answers 45 (0.53%) 7,590 (88.63%) 12 (0.14%) 484 (5.65%) 433 (5.06%) 0 (0.00%)
Table 9: Video-LLaVA Performance on All the Settings of STAR. Each cell shows the count of selections for each option (a0a4) along with the percentage in parentheses. Accuracy represents the number of correct answers where applicable.
Setting a0 a1 a2 a3 a4 N/A Accuracy
Target 1,755 (24.73%) 1,758 (24.77%) 1,742 (24.54%) 1,843 (25.97%) 0 (0.00%)
Default 909 (12.81%) 4,748 (66.89%) 854 (12.03%) 587 (8.27%) 0 (0.00%) 2,464 (34.71%)
Answer Shuffling 825 (11.62%) 4,688 (66.05%) 675 (9.51%) 910 (12.82%) 0 (0.00%) 2,521 (35.52%)
Correct Frames 1,115 (15.71%) 4,611 (64.96%) 743 (10.47%) 629 (8.86%) 0 (0.00%) 1,881 (26.50%)
Rephrased Questions 759 (10.70%) 4,945 (69.70%) 799 (11.26%) 592 (8.34%) 3 (0.04%) 2,455 (34.60%)
Additional Empty Option 346 (4.87%) 4,472 (63.00%) 545 (7.68%) 1,541 (21.71%) 194 (2.73%) 0 (0.00%) 2,490 (35.08%)
All a0 499 (7.03%) 6,596 (92.93%) 1 (0.01%) 2 (0.03%) 0 (0.00%)
All a1 525 (7.40%) 6,573 (92.60%) 0 (0.00%) 0 (0.00%) 0 (0.00%)
All a2 523 (7.37%) 6,571 (92.58%) 1 (0.01%) 3 (0.04%) 0 (0.00%)
All a3 485 (6.83%) 6,610 (93.12%) 2 (0.03%) 1 (0.01%) 0 (0.00%)
Correct a0 1,438 (20.26%) 4,598 (64.78%) 459 (6.47%) 603 (8.50%) 0 (0.00%) 1,438 (20.26%)
Correct a0 Shuffled 1,424 (20.06%) 4,578 (64.51%) 453 (6.38%) 642 (9.05%) 1 (0.01%) 1,424 (20.06%)
Correct a1 533 (7.51%) 5,621 (79.19%) 426 (6.00%) 518 (7.30%) 0 (0.00%) 5,621 (79.19%)
Correct a1 Shuffled 566 (7.98%) 5,560 (78.34%) 449 (6.33%) 522 (7.36%) 1 (0.01%) 5,560 (78.34%)
Correct a2 1,055 (14.86%) 4,561 (64.26%) 849 (11.96%) 633 (8.92%) 0 (0.00%) 849 (11.96%)
Correct a2 Shuffled 586 (8.26%) 4,486 (63.20%) 1,381 (19.46%) 645 (9.09%) 0 (0.00%) 1,381 (19.46%)
Correct a3 588 (8.28%) 4,387 (61.81%) 458 (6.45%) 1,665 (23.46%) 0 (0.00%) 1,665 (23.46%)
Correct a3 Shuffled 1,061 (14.95%) 4,503 (63.44%) 741 (10.44%) 793 (11.17%) 0 (0.00%) 793 (11.17%)
All Correct Answers 374 (5.27%) 6,710 (94.53%) 1 (0.01%) 13 (0.18%) 0 (0.00%)
Empty Frames 709 (9.99%) 4,772 (67.23%) 1,048 (14.76%) 569 (8.02%) 0 (0.00%) 1,891 (26.64%)
Empty Questions 3,013 (42.45%) 3,748 (52.80%) 194 (2.73%) 143 (2.01%) 0 (0.00%) 2,119 (29.85%)
Empty Answers 63 (0.89%) 7,033 (99.08%) 0 (0.00%) 2 (0.03%) 0 (0.00%)
Table 10: Video-LLaVA Performance on All the Settings of Perception Test. Each cell shows the count of selections for each option (a0a3) along with the percentage in parentheses. Accuracy represents the number of correct answers where applicable.
Setting a0 a1 a2 a3 N/A Accuracy
Target 2,549 (33.29%) 2,561 (33.45%) 2,546 (33.25%) 0 (0.00%)
Default 1,944 (25.43%) 5,174 (67.68%) 527 (6.89%) 11 (0.14%) 3,114 (40.73%)
Answer Shuffling 2,164 (28.28%) 4,887 (63.86%) 602 (7.87%) 3 (0.04%) 3,259 (42.58%)
Rephrased Questions 1,836 (24.04%) 5,253 (68.79%) 547 (7.16%) 20 (0.26%) 3,154 (41.30%)
Additional Empty Option 1,133 (15.01%) 4,548 (60.24%) 1,634 (21.64%) 235 (3.11%) 106 (1.38%) 3,233 (42.82%)
All a0 2,942 (38.43%) 4,586 (59.90%) 128 (1.67%) 0 (0.00%)
All a1 3,001 (39.20%) 4,538 (59.27%) 117 (1.53%) 0 (0.00%)
All a2 3,684 (48.13%) 3,951 (51.61%) 20 (0.26%) 1 (0.01%)
Correct a0 3,086 (40.32%) 4,040 (52.79%) 527 (6.89%) 3 (0.04%) 3,086 (40.32%)
Correct a0 Shuffled 3,121 (40.78%) 4,038 (52.76%) 494 (6.45%) 3 (0.04%) 3,121 (40.78%)
Correct a1 1,524 (19.91%) 5,712 (74.62%) 419 (5.47%) 1 (0.01%) 5,712 (74.62%)
Correct a1 Shuffled 1,542 (20.15%) 5,712 (74.63%) 400 (5.23%) 2 (0.03%) 5,712 (74.63%)
Correct a2 3,085 (40.32%) 4,040 (52.80%) 527 (6.89%) 4 (0.05%) 527 (6.89%)
Correct a2 Shuffled 1,839 (24.03%) 4,944 (64.60%) 870 (11.37%) 3 (0.04%) 870 (11.37%)
All Correct Answers 3,493 (45.64%) 4,078 (53.28%) 83 (1.08%) 2 (0.03%)
Empty Frames 1,920 (25.08%) 5,265 (68.77%) 471 (6.15%) 0 (0.00%) 2,957 (38.62%)
Empty Questions 3,649 (47.66%) 3,767 (49.20%) 240 (3.13%) 0 (0.00%) 3,007 (39.28%)
Empty Answers 1,183 (15.45%) 6,284 (82.09%) 188 (2.46%) 1 (0.01%)
Table 11: Video-LLaVA Performance on All the Settings of Video-MME. Each cell shows the count of selections for each option (a0a4) along with the percentage in parentheses. Accuracy represents the number of correct answers where applicable.
Setting a0 a1 a2 a3 a4 N/A Accuracy
Target 675 (25.00%) 675 (25.00%) 675 (25.00%) 675 (25.00%) 0 (0.00%)
Default 476 (17.63%) 1,693 (62.70%) 353 (13.07%) 178 (6.59%) 0 (0.00%) 924 (34.22%)
Answer Shuffling 427 (15.81%) 1,728 (64.00%) 373 (13.81%) 172 (6.37%) 0 (0.00%) 841 (31.15%)
Rephrased Questions 478 (17.70%) 1,705 (63.15%) 352 (13.04%) 165 (6.11%) 0 (0.00%) 911 (33.74%)
Additional Empty Option 208 (7.84%) 1,833 (69.12%) 191 (7.20%) 420 (15.84%) 0 (0.00%) 48 (1.78%) 909 (34.28%)
All a0 891 (33.00%) 1,730 (64.07%) 27 (1.00%) 52 (1.93%) 0 (0.00%)
All a1 888 (32.90%) 1,733 (64.21%) 33 (1.22%) 45 (1.67%) 1 (0.04%)
All a2 914 (33.86%) 1,703 (63.10%) 27 (1.00%) 55 (2.04%) 1 (0.04%)
All a3 888 (32.90%) 1,657 (61.39%) 32 (1.19%) 122 (4.52%) 1 (0.04%)
Correct a0 660 (24.44%) 1,651 (61.15%) 258 (9.56%) 131 (4.85%) 0 (0.00%) 660 (24.44%)
Correct a0 Shuffled 642 (23.78%) 1,620 (60.00%) 293 (10.85%) 145 (5.37%) 0 (0.00%) 642 (23.78%)
Correct a1 384 (14.22%) 1,974 (73.11%) 241 (8.93%) 101 (3.74%) 0 (0.00%) 1,974 (73.11%)
Correct a1 Shuffled 364 (13.48%) 1,964 (72.74%) 240 (8.89%) 132 (4.89%) 0 (0.00%) 1,964 (72.74%)
Correct a2 386 (14.30%) 1,606 (59.48%) 566 (20.96%) 142 (5.26%) 0 (0.00%) 566 (20.96%)
Correct a2 Shuffled 352 (13.04%) 1,623 (60.11%) 589 (21.81%) 136 (5.04%) 0 (0.00%) 589 (21.81%)
Correct a3 416 (15.41%) 1,642 (60.81%) 311 (11.52%) 331 (12.26%) 0 (0.00%) 331 (12.26%)
Correct a3 Shuffled 357 (13.23%) 1,682 (62.32%) 345 (12.78%) 315 (11.67%) 1 (0.04%) 315 (11.67%)
All Correct Answers 928 (34.38%) 1,702 (63.06%) 29 (1.07%) 40 (1.48%) 1 (0.04%)
Empty Frames 267 (9.89%) 1,828 (67.70%) 437 (16.19%) 168 (6.22%) 0 (0.00%) 764 (28.30%)
Empty Questions 1,545 (57.24%) 960 (35.57%) 119 (4.41%) 75 (2.78%) 1 (0.04%) 790 (29.27%)
Empty Answers 359 (13.32%) 2,082 (77.23%) 165 (6.12%) 90 (3.34%) 4 (0.15%)
Table 12: SeViLA Performance on All the Settings of NExT-QA (including NExT-GQA). Each cell shows the count of selections for each option (a0a5) along with the percentage in parentheses. Accuracy represents the number of correct answers where applicable.
Setting a0 a1 a2 a3 a4 a5 Accuracy
Target 1,721 (20.10%) 1,671 (19.51%) 1,767 (20.63%) 1,719 (20.07%) 1,686 (19.69%)
Default 1,812 (21.16%) 1,606 (18.75%) 1,712 (19.99%) 1,805 (21.08%) 1,629 (19.02%) 5,462 (63.78%)
Answer Shuffling 1,740 (20.32%) 1,598 (18.66%) 1,737 (20.28%) 1,824 (21.30%) 1,665 (19.44%) 5,481 (64.00%)
Correct Frames 1,108 (22.33%) 939 (18.92%) 961 (19.37%) 1,069 (21.54%) 885 (17.84%) 2,932 (59.09%)
Rephrased Questions 1,825 (21.31%) 1,571 (18.34%) 1,657 (19.35%) 1,793 (20.94%) 1,718 (20.06%) 5,160 (60.25%)
Additional Empty Option 1,669 (19.49%) 1,401 (16.36%) 1,457 (17.01%) 1,563 (18.25%) 1,758 (20.53%) 716 (8.36%) 5,054 (59.01%)
All a0 4,972 (58.06%) 381 (4.45%) 631 (7.37%) 1,865 (21.78%) 715 (8.35%)
All a1 4,906 (57.29%) 381 (4.45%) 675 (7.88%) 1,954 (22.82%) 648 (7.57%)
All a2 4,887 (57.06%) 359 (4.19%) 681 (7.95%) 1,968 (22.98%) 669 (7.81%)
All a3 4,961 (57.93%) 337 (3.94%) 601 (7.02%) 1,981 (23.13%) 684 (7.99%)
All a4 4,904 (57.26%) 370 (4.32%) 615 (7.18%) 1,987 (23.20%) 688 (8.03%)
Correct a0 5,434 (63.45%) 732 (8.55%) 761 (8.89%) 844 (9.86%) 793 (9.26%) 5,434 (63.45%)
Correct a0 Shuffled 5,443 (63.56%) 736 (8.59%) 745 (8.70%) 814 (9.50%) 826 (9.65%) 5,443 (63.56%)
Correct a1 870 (10.16%) 5,399 (63.04%) 752 (8.78%) 825 (9.63%) 718 (8.38%) 5,399 (63.04%)
Correct a1 Shuffled 809 (9.45%) 5,379 (62.81%) 740 (8.64%) 866 (10.11%) 770 (8.99%) 5,379 (62.81%)
Correct a2 813 (9.49%) 741 (8.65%) 5,495 (64.16%) 772 (9.01%) 743 (8.68%) 5,495 (64.16%)
Correct a2 Shuffled 777 (9.07%) 734 (8.57%) 5,504 (64.27%) 796 (9.29%) 753 (8.79%) 5,504 (64.27%)
Correct a3 827 (9.66%) 704 (8.22%) 778 (9.08%) 5,591 (65.28%) 664 (7.75%) 5,591 (65.28%)
Correct a3 Shuffled 782 (9.13%) 684 (7.99%) 783 (9.14%) 5,566 (64.99%) 749 (8.75%) 5,566 (64.99%)
Correct a4 860 (10.04%) 730 (8.52%) 708 (8.27%) 853 (9.96%) 5,413 (63.21%) 5,413 (63.21%)
Correct a4 Shuffled 810 (9.46%) 727 (8.49%) 760 (8.87%) 831 (9.70%) 5,436 (63.48%) 5,436 (63.48%)
All Correct Answers 7,300 (85.24%) 209 (2.44%) 237 (2.77%) 448 (5.23%) 370 (4.32%)
Empty Frames 1,833 (21.40%) 1,495 (17.46%) 1,626 (18.99%) 1,820 (21.25%) 1,790 (20.90%) 3,921 (45.78%)
Empty Questions 1,671 (19.51%) 1,513 (17.67%) 1,685 (19.68%) 1,851 (21.61%) 1,844 (21.53%) 4,485 (52.37%)
Empty Answers 2,026 (23.66%) 52 (0.61%) 5 (0.06%) 0 (0.00%) 6,481 (75.68%)
Table 13: SeViLA Performance on All the Settings of STAR. Each cell shows the count of selections for each option (a0a4) along with the percentage in parentheses. Accuracy represents the number of correct answers where applicable.
Setting a0 a1 a2 a3 a4 Accuracy
Target 1,755 (24.73%) 1,758 (24.77%) 1,742 (24.54%) 1,843 (25.97%)
Default 1,501 (21.15%) 1,739 (24.50%) 1,782 (25.11%) 2,076 (29.25%) 3,285 (46.28%)
Answer Shuffling 1,473 (20.75%) 1,713 (24.13%) 1,807 (25.46%) 2,105 (29.66%) 3,260 (45.93%)
Correct Frames 1,526 (21.50%) 1,757 (24.75%) 1,759 (24.78%) 2,056 (28.97%) 3,248 (45.76%)
Rephrased Questions 1,507 (21.23%) 1,736 (24.46%) 1,771 (24.95%) 2,084 (29.36%) 3,153 (44.42%)
Additional Empty Option 1,003 (14.13%) 1,077 (15.17%) 1,214 (17.10%) 1,395 (19.65%) 2,409 (33.94%) 2,351 (33.12%)
All a0 2,951 (41.58%) 510 (7.19%) 480 (6.76%) 3,157 (44.48%)
All a1 2,909 (40.98%) 492 (6.93%) 480 (6.76%) 3,217 (45.32%)
All a2 2,935 (41.35%) 572 (8.06%) 442 (6.23%) 3,149 (44.36%)
All a3 2,892 (40.74%) 550 (7.75%) 472 (6.65%) 3,184 (44.86%)
Correct a0 2,806 (39.53%) 1,342 (18.91%) 1,351 (19.03%) 1,599 (22.53%) 2,806 (39.53%)
Correct a0 Shuffled 2,864 (40.35%) 1,274 (17.95%) 1,312 (18.48%) 1,648 (23.22%) 2,864 (40.35%)
Correct a1 1,149 (16.19%) 3,216 (45.31%) 1,231 (17.34%) 1,502 (21.16%) 3,216 (45.31%)
Correct a1 Shuffled 1,113 (15.68%) 3,181 (44.82%) 1,335 (18.81%) 1,469 (20.70%) 3,181 (44.82%)
Correct a2 1,141 (16.07%) 1,223 (17.23%) 3,253 (45.83%) 1,481 (20.87%) 3,253 (45.83%)
Correct a2 Shuffled 1,090 (15.36%) 1,252 (17.64%) 3,301 (46.51%) 1,455 (20.50%) 3,301 (46.51%)
Correct a3 1,088 (15.33%) 1,215 (17.12%) 1,189 (16.75%) 3,606 (50.80%) 3,606 (50.80%)
Correct a3 Shuffled 1,074 (15.13%) 1,194 (16.82%) 1,242 (17.50%) 3,588 (50.55%) 3,588 (50.55%)
All Correct Answers 3,363 (47.38%) 391 (5.51%) 171 (2.41%) 3,173 (44.70%)
Empty Frames 1,501 (21.15%) 1,739 (24.50%) 1,782 (25.11%) 2,076 (29.25%) 3,285 (46.28%)
Empty Questions 1,563 (22.02%) 1,608 (22.65%) 1,700 (23.95%) 2,227 (31.38%) 3,496 (49.25%)
Empty Answers 1,031 (14.53%) 25 (0.35%) 0 (0.00%) 6,042 (85.12%)
Table 14: SeViLA Performance on All the Settings of Perception Test. Each cell shows the count of selections for each option (a0a3) along with the percentage in parentheses. Accuracy represents the number of correct answers where applicable.
Setting a0 a1 a2 a3 Accuracy
Target 2,549 (33.29%) 2,561 (33.45%) 2,546 (33.25%)
Default 2,256 (29.47%) 2,331 (30.45%) 3,069 (40.09%) 3,468 (45.30%)
Answer Shuffling 2,323 (30.34%) 2,300 (30.04%) 3,033 (39.62%) 3,519 (45.96%)
Rephrased Questions 2,286 (29.86%) 2,231 (29.14%) 3,139 (41.00%) 3,369 (44.00%)
Additional Empty Option 1,674 (21.87%) 1,246 (16.27%) 1,436 (18.76%) 3,300 (43.10%) 2,223 (29.04%)
All a0 3,380 (44.15%) 464 (6.06%) 3,812 (49.79%)
All a1 3,427 (44.76%) 449 (5.86%) 3,780 (49.37%)
All a2 3,377 (44.11%) 449 (5.86%) 3,830 (50.03%)
Correct a0 3,128 (40.86%) 1,978 (25.84%) 2,550 (33.31%) 3,128 (40.86%)
Correct a0 Shuffled 3,112 (40.65%) 2,022 (26.41%) 2,522 (32.94%) 3,112 (40.65%)
Correct a1 1,909 (24.93%) 3,249 (42.44%) 2,498 (32.63%) 3,249 (42.44%)
Correct a1 Shuffled 1,869 (24.41%) 3,268 (42.69%) 2,519 (32.90%) 3,268 (42.69%)
Correct a2 1,858 (24.27%) 1,714 (22.39%) 4,084 (53.34%) 4,084 (53.34%)
Correct a2 Shuffled 1,846 (24.11%) 1,750 (22.86%) 4,060 (53.03%) 4,060 (53.03%)
All Correct Answers 3,531 (46.12%) 353 (4.61%) 3,772 (49.27%)
Empty Frames 2,684 (35.06%) 2,058 (26.88%) 2,914 (38.06%) 2,766 (36.13%)
Empty Questions 2,574 (33.62%) 2,040 (26.65%) 3,042 (39.73%) 3,139 (41.00%)
Empty Answers 985 (12.87%) 0 (0.00%) 6,671 (87.13%)
Table 15: SeViLA Performance on All the Settings of Video-MME. Each cell shows the count of selections for each option (a0a4) along with the percentage in parentheses. Accuracy represents the number of correct answers where applicable.
Setting a0 a1 a2 a3 a4 Accuracy
Target 675 (25.00%) 675 (25.00%) 675 (25.00%) 675 (25.00%)
Default 911 (33.74%) 592 (21.93%) 568 (21.04%) 629 (23.30%) 1,076 (39.85%)
Answer Shuffling 887 (32.85%) 566 (20.96%) 591 (21.89%) 656 (24.30%) 1,041 (38.56%)
Rephrased Questions 947 (35.07%) 556 (20.59%) 552 (20.44%) 645 (23.89%) 1,041 (38.56%)
Additional Empty Option 817 (30.26%) 413 (15.30%) 428 (15.85%) 472 (17.48%) 570 (21.11%) 875 (32.41%)
All a0 1,631 (60.41%) 97 (3.59%) 140 (5.19%) 832 (30.81%)
All a1 1,681 (62.26%) 76 (2.81%) 164 (6.07%) 779 (28.85%)
All a2 1,637 (60.63%) 100 (3.70%) 148 (5.48%) 815 (30.19%)
All a3 1,573 (58.26%) 117 (4.33%) 153 (5.67%) 857 (31.74%)
Correct a0 1,271 (47.07%) 469 (17.37%) 451 (16.70%) 509 (18.85%) 1,271 (47.07%)
Correct a0 Shuffled 1,229 (45.52%) 464 (17.19%) 469 (17.37%) 538 (19.93%) 1,229 (45.52%)
Correct a1 752 (27.85%) 977 (36.19%) 460 (17.04%) 511 (18.93%) 977 (36.19%)
Correct a1 Shuffled 714 (26.44%) 951 (35.22%) 474 (17.56%) 561 (20.78%) 951 (35.22%)
Correct a2 793 (29.37%) 443 (16.41%) 990 (36.67%) 474 (17.56%) 990 (36.67%)
Correct a2 Shuffled 755 (27.96%) 430 (15.93%) 984 (36.44%) 531 (19.67%) 984 (36.44%)
Correct a3 725 (26.85%) 451 (16.70%) 463 (17.15%) 1,061 (39.30%) 1,061 (39.30%)
Correct a3 Shuffled 746 (27.63%) 433 (16.04%) 475 (17.59%) 1,046 (38.74%) 1,046 (38.74%)
All Correct Answers 1,840 (68.15%) 75 (2.78%) 93 (3.44%) 692 (25.63%)
Empty Frames 912 (33.78%) 569 (21.07%) 555 (20.56%) 664 (24.59%) 773 (28.63%)
Empty Questions 876 (32.44%) 523 (19.37%) 571 (21.15%) 730 (27.04%) 1,020 (37.78%)
Empty Answers 727 (26.93%) 29 (1.07%) 37 (1.37%) 1,907 (70.63%)
Appendix D Debiasing Algorithms and Results
The BOLD algorithm 1 describes estimating and adjusting for prior probabilities derived from decomposed inputs of ill-defined tasks. The Weighted_BOLD algorithm  2 is an extension of Algorithm  1, this method introduces weighting to the prior estimation to enhance debiasing by better capturing critical priors.
We fixed the random seed to 1 and conducted a series of experiments varying the weights and the sampling parameter $k$. All results including the default values are presented in Tables  1619.
When considering only positive weights ($w_{i}\geq 0$), we assume that the debiasing vector is oriented correctly and use the weights to fine-tune the magnitude of its decomposed constituents. Therefore, positive weights generally enhance improvements and aggravate deteriorations. Specifically, in this case, all the primary debiasing metrics improve along with Recall_std, albeit at different rates across models and datasets.
When we optimize a single bias metric, i. e. Recall_std, by allowing negative weights, we search the latent subspace for the most optimal vector to improve that specific metric. However, this optimization may not correlate with an improvement in the overall bias, as it focuses solely on that metric. Consequently, using negative weights can lead to deterioration in other bias and performance metrics. This observation underscores the importance of evaluating bias using multiple metrics.
We provide the results for both cases, when weights are constrained to $0\leq w_{i}\leq 1$ and when $|w_{i}|\leq 1$, as we believe that the observation of the direction of the bias vectors is crucial for understanding the overall debiasing effect.
Algorithm 1 BOLD: Debiasing with Prior Estimation by Decomposition
1:Video-Language model $M$, dataset $\mathcal{D}=\{T_{i}\}$, number of options $n$, estimation budget $K$
2:Model predictions $\mathcal{Y}$
3:Initialize prediction set $\mathcal{Y}\leftarrow\emptyset$ and prior set $\mathcal{P}\leftarrow\emptyset$triangleright Initialization
4:Sample estimation set $\mathcal{D}_{k}$ of size $K$ from $\mathcal{D}$, set $\mathcal{D}_{r}\leftarrow\mathcal{D}\setminus\mathcal{D}_{k}$
5:for $T\in\mathcal{D}_{k}$ do
6:     Apply attacks to get observed predictions:
7:$P_{o}(d_{i}\mid A|_{v=0}(T))$, $P_{o}(d_{i}\mid A|_{q=0}(T))$, $P_{o}(d_{i}\mid A|_{\text{opts}=0}(T))$
8:     Compute decomposed priors using Equation 4:
9:$P_{\text{pri}}(d_{i}\mid A|_{v=0}(T))$, $P_{\text{pri}}(d_{i}\mid A|_{q=0}(T))$, $P_{\text{pri}}(d_{i}\mid A|_{\text{opts}=0}(T))$
10:     Estimate sample-specific prior $\tilde{P}_{\text{prior}}(d_{i}\mid T)$ using Equation 3; add to $\mathcal{P}$
11:end for
12:Estimate global prior $\tilde{P}_{\text{prior}}(d_{i})$ by averaging $\mathcal{P}$triangleright Prior Estimation
13:for $T\in\mathcal{D}$ do
14:     Debias model prediction using the global prior:
15:$P_{\text{deb}}(d_{i}\mid T)=\operatorname{softmax}\left(\log P_{o}(d_{i}\mid T)-\log\tilde{P}_{\text{prior}}(d_{i})\right)$
16:     Add the predicted answer to $\mathcal{Y}$
17:end for
18:return $\mathcal{Y}$
Algorithm 2 Weighted_BOLD: Debiasing with Prior Estimation by Decomposition with Weights
1:Model $M$, dataset $\mathcal{D}=\{T_{i}\}$, estimation budget $K$, number of folds $F=5$
2:Debiased model predictions $\mathcal{Y}$
3:Initialize prediction set $\mathcal{Y}\leftarrow\emptyset$ and prior set $\mathcal{P}\leftarrow\emptyset$triangleright Initialization
4:Sample estimation set $\mathcal{D}_{k}$ of size $K$ from $\mathcal{D}$; set remaining samples $\mathcal{D}_{r}\leftarrow\mathcal{D}\setminus\mathcal{D}_{k}$
5:for each fold $f=1$ to $F$ do $\triangleright$ Cross-validation
6:     Split $\mathcal{D}_{k}$ into test set $\mathcal{D}_{\text{test}}^{f}$ and validation set $\mathcal{D}_{\text{val}}^{f}$
7:     for $T\in\mathcal{D}_{\text{test}}^{f}$ do
8:         Apply attacks to obtain observed predictions: $P_{o}(d_{i}\mid A(T))$
9:         Compute decomposed priors for each attack using Equation 4: $P_{\text{pri}}(d_{i}\mid A(T))$
10:         Estimate sample-specific prior $\tilde{P}_{\text{pri}}(d_{i}\mid T)$ using Equation 6; add to $\mathcal{P}^{f}$
11:     end for
12:     Optimize weights $\{w_{i}\}$ using COBYLA, subject to constraints $0\leq w_{i}\leq 1$ or $|w_{i}|\leq 1$, to minimize recall_std bias on $\mathcal{D}_{\text{test}}^{f}$
13:end for
14:Compute global prior by averaging over folds:
$\tilde{P}_{\text{prior}}(d_{i})=\frac{1}{F}\sum_{f=1}^{F}\tilde{P}_{\text{pri}}^{f}(d_{i})$
$\triangleright$ Global Prior Estimation
15:for $T\in\mathcal{D}$ do $\triangleright$ Debiasing
16:     Debias the model prediction using the global prior:
$P_{\text{deb}}(d_{i}\mid T)=\operatorname{softmax}\left(\log P_{o}(d_{i}\mid T)-\log\tilde{P}_{\text{prior}}(d_{i})\right)$
17:     Add the predicted answer to $\mathcal{Y}$
18:end for
19:return $\mathcal{Y}$
Table 16: Comparison of BOLD and Weighted_BOLD bias mitigation approaches across models and datasets for performance and bias monitoring metrics with $k$ = 0.25. Green arrows indicate improvements: upward for Accuracy and F1_mean, and downward for standard deviation metrics. Red arrows represent deterioration, respectively.
Model Dataset Configuration Performance Metrics Bias Monitoring Metrics
Accuracy F1_mean Recall_std F1_std JS_std
NExT-QA Default 49.95 49.81 16.31 3.27 5.4
BOLD 51.81 (3.72%) 51.71 (3.82%) 13.29 (18.54%) 2.66 (18.59%) 4.58 (15.12%)
+Weighted_BOLD 51.89 (3.88%) 51.77 (3.95%) 13.28 (18.57%) 2.73 (16.57%) 4.56 (15.39%)
-/+Weighted_BOLD 52.28 (4.65%) 52.25 (4.9%) 10.9 (33.18%) 2.23 (31.89%) 3.9 (27.63%)
Video-LLaVA STAR Default 34.77 31.83 24.99 6.82 5.78
BOLD 37.45 (7.7%) 36.14 (13.54%) 17.43 (30.23%) 4.84 (29.02%) 4.52 (21.74%)
+Weighted_BOLD 37.81 (8.75%) 36.67 (15.22%) 16.15 (35.38%) 4.71 (30.94%) 4.15 (28.15%)
-/+Weighted_BOLD 39.87 (14.67%) 39.65 (24.57%) 6.85 (72.56%) 2.34 (65.67%) 2.08 (63.98%)
Perception Test Default 40.72 35.69 27.5 14.75 5.28
BOLD 41.62 (2.22%) 38.68 (8.38%) 21.77 (20.85%) 10.93 (25.88%) 3.77 (28.54%)
+Weighted_BOLD 41.95 (3.02%) 39.46 (10.58%) 19.73 (28.26%) 10.24 (30.53%) 3.39 (35.79%)
-/+Weighted_BOLD 42.01 (3.18%) 40.48 (13.43%) 14.69 (46.59%) 8.48 (42.52%) 2.91 (44.9%)
Video-MME Default 34.3 30.99 24.09 8.06 4.79
BOLD 34.63 (0.97%) 32.71 (5.54%) 18.21 (24.4%) 6.18 (23.36%) 3.86 (19.47%)
+Weighted_BOLD 34.59 (0.86%) 33.01 (6.51%) 15.92 (33.93%) 5.91 (26.68%) 3.38 (29.56%)
-/+Weighted_BOLD 33.93 (1.08%) 33.21 (7.16%) 10.1 (58.07%) 4.09 (49.26%) 2.33 (51.39%)
NExT-QA Default 44.79 40.85 23.83 15.86 16.09
BOLD 45.88 (2.43%) 42.15 (3.18%) 22.98 (3.53%) 15.53 (2.11%) 15.56 (3.3%)
+Weighted_BOLD 45.96 (2.61%) 42.27 (3.46%) 22.78 (4.4%) 15.46 (2.55%) 15.52 (3.55%)
-/+Weighted_BOLD 47.56 (6.19%) 46.63 (14.14%) 13.79 (42.11%) 8.47 (46.62%) 12.14 (24.53%)
Video-LLaMA STAR Default 36.52 31.86 24.1 15.44 15.29
BOLD 37.19 (1.82%) 33.02 (3.65%) 22.29 (7.5%) 14.55 (5.75%) 14.58 (4.6%)
+Weighted_BOLD 37.21 (1.89%) 33.32 (4.59%) 21.45 (10.97%) 14.03 (9.1%) 14.22 (7.0%)
-/+Weighted_BOLD 38.63 (5.76%) 38.07 (19.52%) 10.16 (57.84%) 4.78 (69.05%) 8.93 (41.61%)
Perception Test Default 41.39 37.19 27.06 11.4 16.06
BOLD 41.92 (1.27%) 38.7 (4.06%) 24.04 (11.16%) 9.85 (13.61%) 14.44 (10.12%)
+Weighted_BOLD 42.06 (1.62%) 39.36 (5.86%) 21.8 (19.45%) 9.11 (20.06%) 13.05 (18.75%)
-/+Weighted_BOLD 42.05 (1.59%) 41.35 (11.2%) 11.13 (58.86%) 4.71 (58.65%) 8.43 (47.51%)
Video-MME Default 32.05 28.15 20.62 12.52 13.75
BOLD 32.69 (2.01%) 29.2 (3.73%) 19.28 (6.5%) 11.95 (4.54%) 13.13 (4.53%)
+Weighted_BOLD 32.2 (0.47%) 28.98 (2.94%) 17.65 (14.38%) 11.69 (6.56%) 12.6 (8.39%)
-/+Weighted_BOLD 32.92 (2.72%) 30.39 (7.97%) 15.65 (24.08%) 10.21 (18.39%) 11.68 (15.04%)
NExT-QA Default 63.91 63.88 2.15 1.32 1.99
BOLD 63.92 (0.02%) 63.89 (0.03%) 2.0 (6.81%) 1.18 (10.43%) 1.97 (0.91%)
+Weighted_BOLD 63.88 (0.04%) 63.86 (0.03%) 2.08 (3.47%) 1.28 (2.87%) 1.9 (4.64%)
-/+Weighted_BOLD 63.95 (0.07%) 63.95 (0.11%) 1.44 (32.8%) 1.16 (12.17%) 1.63 (18.3%)
SeViLA STAR Default 46.28 46.14 4.47 2.3 2.25
BOLD 46.17 (0.24%) 46.05 (0.2%) 4.09 (8.42%) 2.24 (2.93%) 2.11 (6.25%)
+Weighted_BOLD 46.07 (0.46%) 45.95 (0.42%) 4.12 (7.63%) 2.28 (1.24%) 2.13 (5.17%)
-/+Weighted_BOLD 46.2 (0.18%) 46.08 (0.13%) 4.04 (9.49%) 2.23 (3.15%) 2.16 (3.9%)
Perception Test Default 45.3 45.11 6.21 3.02 1.59
BOLD 45.28 (0.03%) 45.15 (0.09%) 5.1 (17.85%) 2.93 (3.09%) 1.26 (21.04%)
+Weighted_BOLD 45.34 (0.09%) 45.24 (0.29%) 4.26 (31.39%) 2.65 (12.43%) 1.03 (35.41%)
-/+Weighted_BOLD 45.22 (0.17%) 45.15 (0.09%) 3.69 (40.51%) 2.38 (21.06%) 0.85 (46.67%)
Video-MME Default 39.85 39.82 4.67 1.68 0.84
BOLD 40.04 (0.46%) 40.04 (0.56%) 3.79 (18.89%) 1.48 (12.03%) 0.74 (11.46%)
+Weighted_BOLD 40.07 (0.56%) 40.08 (0.66%) 3.64 (22.05%) 1.48 (11.82%) 0.67 (20.11%)
-/+Weighted_BOLD 40.04 (0.46%) 40.06 (0.63%) 2.31 (50.5%) 1.36 (18.83%) 0.3 (63.83%)
Table 17: Comparison of BOLD and Weighted_BOLD bias mitigation approaches across models and datasets for performance and bias monitoring metrics with $k$ = 0.5. Green arrows indicate improvements: upward for Accuracy and F1_mean, and downward for standard deviation metrics. Red arrows represent deterioration, respectively.
Model Dataset Configuration Performance Metrics Bias Monitoring Metrics
Accuracy F1_mean Recall_std F1_std JS_std
NExT-QA Default 44.79 40.85 23.83 15.86 16.09
BOLD 45.88 (2.43%) 42.15 (3.18%) 22.98 (3.53%) 15.53 (2.11%) 15.55 (3.31%)
+Weighted_BOLD 45.91 (2.51%) 42.2 (3.29%) 22.83 (4.19%) 15.51 (2.2%) 15.56 (3.27%)
-/+Weighted_BOLD 48.25 (7.73%) 46.52 (13.86%) 16.6 (30.32%) 10.6 (33.17%) 13.27 (17.48%)
Video-LLaMA STAR Default 36.52 31.86 24.1 15.44 15.29
BOLD 37.19 (1.82%) 33.02 (3.65%) 22.29 (7.5%) 14.55 (5.75%) 14.58 (4.66%)
+Weighted_BOLD 37.34 (2.24%) 33.5 (5.16%) 21.13 (12.3%) 14.05 (8.98%) 14.14 (7.54%)
-/+Weighted_BOLD 38.61 (5.72%) 38.06 (19.47%) 10.18 (57.75%) 4.79 (68.99%) 8.58 (43.89%)
Perception Test Default 41.39 37.19 27.06 11.4 16.06
BOLD 41.9 (1.24%) 38.68 (4.01%) 24.05 (11.11%) 9.87 (13.45%) 14.45 (10.01%)
+Weighted_BOLD 42.06 (1.62%) 39.36 (5.86%) 21.8 (19.45%) 9.11 (20.06%) 13.02 (18.94%)
-/+Weighted_BOLD 42.27 (2.13%) 41.24 (10.9%) 13.55 (49.93%) 5.64 (50.54%) 9.73 (39.41%)
Video-MME Default 32.05 28.15 20.62 12.52 13.75
BOLD 32.73 (2.13%) 29.23 (3.85%) 19.29 (6.45%) 11.96 (4.46%) 13.1 (4.73%)
+Weighted_BOLD 32.2 (0.47%) 28.98 (2.94%) 17.65 (14.38%) 11.69 (6.56%) 12.64 (8.04%)
-/+Weighted_BOLD 32.88 (2.6%) 31.21 (10.87%) 12.77 (38.06%) 8.02 (35.91%) 10.42 (24.25%)
NExT-QA Default 49.95 49.81 16.31 3.27 5.4
BOLD 51.81 (3.72%) 51.71 (3.82%) 13.27 (18.63%) 2.67 (18.42%) 4.58 (15.06%)
+Weighted_BOLD 52.15 (4.39%) 52.04 (4.49%) 12.72 (22.02%) 2.62 (19.83%) 4.49 (16.83%)
-/+Weighted_BOLD 53.65 (7.41%) 53.53 (7.48%) 8.49 (47.97%) 3.15 (3.61%) 3.78 (29.87%)
Video-LLaVA STAR Default 34.77 31.83 24.99 6.82 5.78
BOLD 37.21 (7.01%) 35.76 (12.37%) 18.28 (26.85%) 4.97 (27.15%) 4.54 (21.39%)
+Weighted_BOLD 37.53 (7.94%) 36.18 (13.69%) 17.45 (30.15%) 4.91 (28.03%) 4.26 (26.24%)
-/+Weighted_BOLD 38.31 (10.17%) 37.39 (17.47%) 13.26 (46.92%) 4.59 (32.59%) 3.25 (43.67%)
Perception Test Default 40.72 35.69 27.5 14.75 5.28
BOLD 41.62 (2.22%) 38.69 (8.4%) 21.75 (20.9%) 10.92 (25.97%) 3.78 (28.4%)
+Weighted_BOLD 41.95 (3.02%) 39.46 (10.58%) 19.73 (28.26%) 10.24 (30.53%) 3.4 (35.51%)
-/+Weighted_BOLD 41.62 (2.22%) 40.88 (14.56%) 11.01 (59.96%) 5.44 (63.11%) 2.59 (50.97%)
Video-MME Default 34.3 30.99 24.09 8.06 4.79
BOLD 34.7 (1.19%) 32.79 (5.81%) 18.19 (24.51%) 6.16 (23.63%) 3.84 (19.93%)
+Weighted_BOLD 34.63 (0.97%) 32.97 (6.38%) 16.36 (32.07%) 6.01 (25.43%) 3.43 (28.5%)
-/+Weighted_BOLD 34.0 (0.86%) 32.6 (5.21%) 14.11 (41.43%) 5.79 (28.18%) 2.92 (39.01%)
NExT-QA Default 63.91 63.88 2.15 1.32 1.99
BOLD 63.92 (0.02%) 63.89 (0.02%) 2.03 (5.47%) 1.18 (10.4%) 1.99 (0.17%)
+Weighted_BOLD 63.93 (0.04%) 63.91 (0.04%) 1.99 (7.64%) 1.19 (9.83%) 1.99 (0.04%)
-/+Weighted_BOLD 64.01 (0.16%) 64.0 (0.19%) 1.14 (47.18%) 1.17 (11.43%) 2.01 (1.1%)
SeViLA STAR Default 46.28 46.14 4.47 2.3 2.25
BOLD 46.22 (0.12%) 46.1 (0.08%) 4.13 (7.46%) 2.26 (1.78%) 2.1 (6.64%)
+Weighted_BOLD 46.2 (0.18%) 46.08 (0.13%) 4.01 (10.17%) 2.2 (4.36%) 2.07 (8.2%)
-/+Weighted_BOLD 46.38 (0.21%) 46.3 (0.34%) 3.41 (23.66%) 1.94 (15.89%) 1.93 (14.33%)
Perception Test Default 45.3 45.11 6.21 3.02 1.59
BOLD 45.32 (0.06%) 45.18 (0.16%) 5.18 (16.61%) 2.95 (2.43%) 1.3 (18.42%)
+Weighted_BOLD 45.31 (0.03%) 45.2 (0.21%) 4.56 (26.58%) 2.83 (6.43%) 1.08 (31.92%)
-/+Weighted_BOLD 45.25 (0.12%) 45.2 (0.2%) 3.25 (47.69%) 2.28 (24.52%) 0.71 (55.13%)
Video-MME Default 39.85 39.82 4.67 1.68 0.84
BOLD 40.19 (0.84%) 40.17 (0.88%) 4.11 (12.03%) 1.41 (16.19%) 0.73 (12.52%)
+Weighted_BOLD 40.04 (0.46%) 40.03 (0.54%) 3.78 (19.14%) 1.44 (14.6%) 0.69 (17.22%)
-/+Weighted_BOLD 39.81 (0.09%) 39.83 (0.04%) 2.99 (35.93%) 1.43 (14.97%) 0.45 (45.74%)
Table 18: Comparison of BOLD and Weighted_BOLD bias mitigation approaches across models and datasets for performance and bias monitoring metrics with $k$ = 0.75. Green arrows indicate improvements: upward for Accuracy and F1_mean, and downward for standard deviation metrics. Red arrows represent deterioration for Accuracy and F1_mean.
Model Dataset Configuration Performance Metrics Bias Monitoring Metrics
Accuracy F1_mean Recall_std F1_std JS_std
NExT-QA Default 44.79 40.85 23.83 15.86 16.09
BOLD 45.87 (2.4%) 42.16 (3.19%) 22.96 (3.63%) 15.49 (2.37%) 15.56 (3.28%)
+Weighted_BOLD 45.91 (2.51%) 42.2 (3.29%) 22.83 (4.19%) 15.51 (2.2%) 15.57 (3.23%)
-/+Weighted_BOLD 47.56 (6.19%) 46.63 (14.14%) 13.36 (43.91%) 8.42 (46.91%) 11.79 (26.73%)
Video-LLaMA STAR Default 36.52 31.86 24.1 15.44 15.29
BOLD 37.19 (1.82%) 33.02 (3.65%) 22.29 (7.5%) 14.55 (5.75%) 14.59 (4.59%)
+Weighted_BOLD 37.34 (2.24%) 33.5 (5.16%) 21.13 (12.3%) 14.05 (8.98%) 14.13 (7.56%)
-/+Weighted_BOLD 38.74 (6.07%) 38.14 (19.74%) 10.58 (56.11%) 4.94 (68.03%) 9.64 (36.94%)
Perception Test Default 41.39 37.19 27.06 11.4 16.06
BOLD 41.9 (1.24%) 38.67 (3.98%) 24.09 (11.0%) 9.88 (13.29%) 14.46 (9.97%)
+Weighted_BOLD 42.06 (1.62%) 39.36 (5.86%) 21.8 (19.45%) 9.11 (20.06%) 13.12 (18.33%)
-/+Weighted_BOLD 42.29 (2.16%) 41.32 (11.13%) 12.96 (52.13%) 5.53 (51.52%) 9.56 (40.45%)
Video-MME Default 32.05 28.15 20.62 12.52 13.75
BOLD 32.69 (2.01%) 29.2 (3.75%) 19.24 (6.7%) 11.94 (4.63%) 13.09 (4.8%)
+Weighted_BOLD 32.2 (0.47%) 28.97 (2.94%) 17.65 (14.38%) 11.7 (6.54%) 12.56 (8.62%)
-/+Weighted_BOLD 32.65 (1.89%) 30.55 (8.55%) 13.94 (32.38%) 9.24 (26.17%) 11.04 (19.73%)
NExT-QA Default 49.95 49.81 16.31 3.27 5.4
BOLD 51.81 (3.72%) 51.71 (3.82%) 13.27 (18.63%) 2.67 (18.42%) 4.58 (15.14%)
+Weighted_BOLD 51.96 (4.02%) 51.86 (4.12%) 12.92 (20.76%) 2.64 (19.44%) 4.48 (17.02%)
-/+Weighted_BOLD 53.75 (7.6%) 53.66 (7.73%) 8.51 (47.82%) 2.58 (21.07%) 3.6 (33.28%)
Video-LLaVA STAR Default 34.77 31.83 24.99 6.82 5.78
BOLD 37.21 (7.01%) 35.76 (12.37%) 18.28 (26.85%) 4.97 (27.15%) 4.62 (20.03%)
+Weighted_BOLD 37.52 (7.9%) 36.17 (13.64%) 17.45 (30.14%) 4.91 (28.02%) 4.27 (26.1%)
-/+Weighted_BOLD 37.86 (8.87%) 36.87 (15.85%) 13.12 (47.51%) 4.68 (31.37%) 3.09 (46.45%)
Perception Test Default 40.72 35.69 27.5 14.75 5.28
BOLD 41.62 (2.22%) 38.68 (8.38%) 21.77 (20.85%) 10.93 (25.88%) 3.79 (28.21%)
+Weighted_BOLD 41.95 (3.02%) 39.46 (10.58%) 19.73 (28.26%) 10.24 (30.53%) 3.41 (35.4%)
-/+Weighted_BOLD 42.21 (3.66%) 41.64 (16.67%) 9.47 (65.56%) 5.14 (65.14%) 1.8 (65.84%)
Video-MME Default 34.3 30.99 24.09 8.06 4.79
BOLD 34.63 (0.97%) 32.64 (5.33%) 18.59 (22.84%) 6.19 (23.2%) 3.85 (19.65%)
+Weighted_BOLD 34.48 (0.54%) 32.89 (6.13%) 15.94 (33.83%) 5.86 (27.32%) 3.37 (29.66%)
-/+Weighted_BOLD 33.78 (1.51%) 33.2 (7.12%) 9.23 (61.7%) 3.35 (58.46%) 1.42 (70.35%)
NExT-QA Default 63.91 63.88 2.15 1.32 1.99
BOLD 63.91 (0.0%) 63.88 (0.01%) 2.03 (5.45%) 1.18 (10.77%) 1.98 (0.34%)
+Weighted_BOLD 63.93 (0.04%) 63.91 (0.04%) 2.0 (6.83%) 1.2 (8.97%) 1.98 (0.23%)
-/+Weighted_BOLD 64.11 (0.31%) 64.09 (0.34%) 1.2 (44.36%) 1.09 (17.41%) 2.15 (8.2%)
SeViLA STAR Default 46.28 46.14 4.47 2.3 2.25
BOLD 46.27 (0.03%) 46.15 (0.02%) 4.07 (8.91%) 2.22 (3.53%) 2.08 (7.72%)
+Weighted_BOLD 46.17 (0.24%) 46.05 (0.19%) 4.01 (10.17%) 2.2 (4.43%) 2.04 (9.14%)
-/+Weighted_BOLD 46.17 (0.24%) 46.1 (0.09%) 2.96 (33.63%) 1.78 (22.67%) 1.64 (27.25%)
Perception Test Default 45.3 45.11 6.21 3.02 1.59
BOLD 45.31 (0.03%) 45.16 (0.12%) 5.25 (15.42%) 2.98 (1.26%) 1.3 (18.44%)
+Weighted_BOLD 45.35 (0.12%) 45.25 (0.32%) 4.27 (31.19%) 2.65 (12.18%) 1.03 (34.92%)
-/+Weighted_BOLD 45.15 (0.32%) 45.09 (0.03%) 3.48 (43.88%) 2.3 (24.02%) 0.78 (51.02%)
Video-MME Default 39.85 39.82 4.67 1.68 0.84
BOLD 40.04 (0.46%) 40.03 (0.53%) 3.84 (17.85%) 1.46 (13.42%) 0.69 (18.0%)
+Weighted_BOLD 39.89 (0.09%) 39.87 (0.15%) 3.7 (20.9%) 1.43 (14.96%) 0.58 (31.13%)
-/+Weighted_BOLD 39.7 (0.37%) 39.71 (0.26%) 2.69 (42.53%) 1.42 (15.75%) 0.15 (81.72%)
Table 19: Comparison of BOLD and Weighted_BOLD bias mitigation approaches across models and datasets for performance and bias monitoring metrics $k$ = 1 (equal to the entire dataset). Green arrows indicate improvements: upward for Accuracy and F1_mean, and downward for standard deviation metrics. Red arrows represent deterioration for Accuracy and F1_mean.
Model Dataset Configuration Performance Metrics Bias Monitoring Metrics
Accuracy F1_mean Recall_std F1_std JS_std
NExT-QA Default 44.79 40.85 23.83 15.86 16.09
BOLD 45.87 (2.4%) 42.16 (3.19%) 22.96 (3.63%) 15.49 (2.37%) 15.56 (3.29%)
+Weighted_BOLD 45.91 (2.51%) 42.2 (3.29%) 22.83 (4.19%) 15.51 (2.2%) 15.57 (3.23%)
-/+Weighted_BOLD 47.53 (6.11%) 46.62 (14.1%) 13.25 (44.39%) 8.46 (46.68%) 11.78 (26.76%)
Video-LLaMA STAR Default 36.52 31.86 24.1 15.44 15.29
BOLD 37.19 (1.82%) 33.02 (3.65%) 22.29 (7.5%) 14.55 (5.75%) 14.59 (4.58%)
+Weighted_BOLD 37.34 (2.24%) 33.5 (5.16%) 21.13 (12.3%) 14.05 (8.98%) 14.13 (7.55%)
-/+Weighted_BOLD 38.72 (6.03%) 38.18 (19.85%) 10.0 (58.49%) 4.85 (68.6%) 8.53 (44.23%)
Perception Test Default 41.39 37.19 27.06 11.4 16.06
BOLD 41.89 (1.21%) 38.67 (3.98%) 24.04 (11.18%) 9.86 (13.52%) 14.45 (10.04%)
+Weighted_BOLD 42.08 (1.65%) 39.38 (5.89%) 21.8 (19.46%) 9.11 (20.03%) 12.93 (19.5%)
-/+Weighted_BOLD 42.25 (2.07%) 41.53 (11.69%) 11.21 (58.59%) 4.76 (58.2%) 8.83 (45.01%)
Video-MME Default 32.05 28.15 20.62 12.52 13.75
BOLD 32.65 (1.89%) 29.13 (3.49%) 19.29 (6.44%) 12.03 (3.86%) 13.09 (4.82%)
+Weighted_BOLD 32.2 (0.47%) 28.97 (2.94%) 17.65 (14.38%) 11.7 (6.54%) 12.57 (8.56%)
-/+Weighted_BOLD 32.92 (2.72%) 31.32 (11.27%) 12.6 (38.89%) 7.85 (37.25%) 10.36 (24.67%)
NExT-QA Default 49.95 49.81 16.31 3.27 5.4
BOLD 51.81 (3.72%) 51.71 (3.82%) 13.27 (18.63%) 2.67 (18.42%) 4.58 (15.18%)
+Weighted_BOLD 51.83 (3.76%) 51.77 (3.93%) 12.79 (21.58%) 2.5 (23.66%) 4.43 (17.81%)
-/+Weighted_BOLD 52.94 (5.98%) 53.0 (6.41%) 8.49 (47.93%) 1.98 (39.48%) 3.02 (43.96%)
Video-LLaVA STAR Default 34.77 31.83 24.99 6.82 5.78
BOLD 37.14 (6.81%) 35.65 (12.02%) 18.53 (25.84%) 5.01 (26.55%) 4.6 (20.38%)
+Weighted_BOLD 37.33 (7.37%) 35.98 (13.04%) 17.62 (29.49%) 4.83 (29.09%) 4.35 (24.73%)
-/+Weighted_BOLD 38.56 (10.9%) 37.86 (18.96%) 11.35 (54.56%) 3.69 (45.92%) 2.91 (49.67%)
Perception Test Default 40.72 35.69 27.5 14.75 5.28
BOLD 41.62 (2.22%) 38.69 (8.4%) 21.75 (20.9%) 10.92 (25.97%) 3.79 (28.14%)
+Weighted_BOLD 41.95 (3.02%) 39.46 (10.58%) 19.73 (28.26%) 10.24 (30.53%) 3.41 (35.37%)
-/+Weighted_BOLD 41.57 (2.09%) 41.11 (15.2%) 8.95 (67.47%) 4.3 (70.81%) 2.11 (59.94%)
Video-MME Default 34.3 30.99 24.09 8.06 4.79
BOLD 34.56 (0.76%) 32.63 (5.3%) 18.24 (24.28%) 6.08 (24.59%) 3.86 (19.41%)
+Weighted_BOLD 34.41 (0.32%) 32.78 (5.77%) 16.1 (33.17%) 5.89 (26.93%) 3.37 (29.69%)
-/+Weighted_BOLD 34.22 (0.22%) 33.56 (8.31%) 9.8 (59.31%) 3.71 (54.0%) 1.84 (61.61%)
NExT-QA Default 63.91 63.88 2.15 1.32 1.99
BOLD 63.91 (0.0%) 63.88 (0.01%) 2.03 (5.45%) 1.18 (10.77%) 1.98 (0.66%)
+Weighted_BOLD 64.0 (0.15%) 63.98 (0.15%) 1.95 (9.2%) 1.24 (6.37%) 1.97 (0.96%)
-/+Weighted_BOLD 64.07 (0.26%) 64.06 (0.29%) 0.84 (60.74%) 1.07 (18.78%) 2.08 (4.64%)
SeViLA STAR Default 46.28 46.14 4.47 2.3 2.25
BOLD 46.25 (0.06%) 46.14 (0.01%) 4.05 (9.25%) 2.22 (3.75%) 2.06 (8.29%)
+Weighted_BOLD 46.15 (0.27%) 46.04 (0.22%) 4.01 (10.23%) 2.2 (4.52%) 2.05 (9.07%)
-/+Weighted_BOLD 46.18 (0.21%) 46.1 (0.1%) 3.48 (22.16%) 2.04 (11.64%) 1.85 (17.64%)
Perception Test Default 45.3 45.11 6.21 3.02 1.59
BOLD 45.4 (0.23%) 45.25 (0.33%) 5.24 (15.53%) 3.04 (0.72%) 1.3 (18.45%)
+Weighted_BOLD 45.34 (0.09%) 45.24 (0.3%) 4.16 (33.06%) 2.62 (13.42%) 1.03 (34.99%)
-/+Weighted_BOLD 45.22 (0.17%) 45.14 (0.08%) 3.85 (38.0%) 2.43 (19.68%) 0.9 (43.13%)
Video-MME Default 39.85 39.82 4.67 1.68 0.84
BOLD 40.0 (0.37%) 39.99 (0.44%) 3.81 (18.5%) 1.48 (11.8%) 0.66 (21.49%)
+Weighted_BOLD 39.81 (0.09%) 39.81 (0.02%) 3.57 (23.49%) 1.45 (13.53%) 0.53 (36.44%)
-/+Weighted_BOLD 39.85 (0.0%) 39.85 (0.08%) 2.66 (43.14%) 1.47 (12.42%) 0.16 (80.65%)