### v2

# Addressing Blind Guessing: Calibration of Selection Bias in Multiple-Choice Question Answering by Video Language Models

Abstract
Evaluating Video Language Models (VLMs) is a challenging task. Due to its transparency, Multiple-Choice Question Answering (MCQA) is widely used to measure the performance of these models through accuracy. However, existing MCQA benchmarks fail to capture the full reasoning capabilities of VLMs due to selection bias, when models disproportionately favor certain answer options based on positional patterns observed during training. In this work, we conduct a comprehensive empirical analysis of several VLM architectures across major datasets designed to assess complex video-focused reasoning. We identify where the bias is most pronounced and demonstrate to what extent model responses reflect genuine understanding of video content and related questions, as opposed to reliance on arbitrary patterns or superficial cues, such as answer position. By decomposing the MCQA task and adapting fairness bias metrics to VLMs, we introduce a post-processing calibration technique BOLD to balance this bias. Our results show that reducing selection bias improves not only debiasing metrics but also overall model performance, including Accuracy and F1 Mean score. Our method, by suppressing "blind guessing", offers a more cost- and time-effective approach to mitigating selection bias compared to existing techniques. This study represents the first focused investigation of selection bias in video-to-text LLM-powered models. [^1] The codes for the settings, models and calibration technique are avaialable on Github: https://github.com/ologin/BOLD
Addressing Blind Guessing: Calibration of Selection Bias in Multiple-Choice Question Answering by Video Language Models
Olga Loginova University of Trento olga.loginova@unitn.it                        Oleksandr Bezrukov Gran Sasso Science Institute oleksandr.bezrukov@gssi.it
Ravi Shekhar University of Essex r.shekhar@essex.ac.uk                        Alexey Kravets University of Bath ak3095@bath.ac.uk
1 Introduction
Multiple-choice question answering (MCQA) is a transparent and convenient method to assess language models’ performance. Each MCQA instance presents a question and several answer options, requiring the selection of the correct one. However, large language models (LLMs) often exhibit selection bias, i.e., a tendency to favor certain answer positions irrespective of content Zheng et al. (2024a); Wang et al. (2024a). Such bias compromises fairness, reliability, and the true reasoning capabilities of these models. Although selection bias in LLMs has received increasing attention, it remains largely unexplored for video-language models (VLMs).
VLMs integrate textual and visual inputs by processing sequences of frames and capturing both spatial and temporal features. Recent advances align pre-trained LLMs with video encoders Bordes et al. (2024). Although VLMs inherit many properties from LLMs, their reliance on complex visual inputs introduces unique challenges, such as higher computational overhead and the need for robust spatiotemporal reasoning. The growing role of VLM-based MCQA tasks thus calls for a tailored investigation of selection bias in VLMs.
We address this gap with the first study of video-MCQA selection bias, focusing on video-specific aspects such as dynamic unfolding of events, spatial object tracking, and camera behavior, rather than general plot understanding.
Figure 1: Decomposition approach in 3 steps: key-component decomposition, rearrangement into pairs, applying probability debiasing to the aggregated data.
Through the empirical analysis, we locate where the model’s predictions are skewed by inherent option preferences. We find that, while each model demonstrates specific biases, accuracy-relevant settings yield option distributions similar to those in the default setting. This suggests that the models rely heavily on answer position, regardless of content. Conversely, in accuracy-irrelevant settings, the predicted distributions deviate significantly from uniformity, and this is where the bias is most apparent. Subsequently, we adapt the metrics from fairness bias studies (e.g., regarding race or age) to video-MCQA and deliver a calibration technique.
Figure 1 illustrates our decomposition strategy. We find that depriving the MCQA task of its key constituent components (video, question, answer options) reveals latent bias factors. We systematically isolate the three components: we remove one component at a time and rearrange the remaining pairs, obtaining three “ill-defined” variants of the original data. These decompositions allow us to identify the underlying bias patterns across different sub-spaces of the task. We then aggregate these insights and apply a probability-based debiasing procedure that adjusts the final predictions.
By experimenting with three VLM architectures and four video-specific MCQA datasets, we show that our post-training practical solution enhances interpretability, reduces computational complexity [^2] Traditionally, selection bias is mitigated through shuffling Zheng et al. (2024a); Liu et al. (2023), where answer options are permuted multiple times to produce more impartial results. Our method requires far fewer resources compared to all the exhaustive answer-option permutations., and improves bias metrics. Intriguingly, while our method is designed to mitigate bias, it generally improves performance with better accuracy and F1 Mean scores.
In summary, our contributions are:
  • A thorough investigation of selection bias in VLM-based MCQA with 11 dataset modifications to target video-specific reasoning challenges.
  • Adaptation of fairness metrics to assess selection bias in VLMs on video MCQA benchmarks.
  • A novel decomposition post-processing calibration technique that leverages the rearrangement of MCQA, mitigates selection bias and improves performance.
Through the refined MCQA benchmarks, we uncover deeper reasoning issues in contemporary VLMs. Our work helps move toward more robust, fair, reliable and interpretable VLMs by bridging a gap in the study of selection bias in video understanding models and offering a cost-effective solution.
2 Related Works
Selection bias in LLMs has been extensively studied Pezeshkpour and Hruschka (2023); Zheng et al. (2023); Wang et al. (2024a); Li et al. (2024); Liusie et al. (2024); Wang et al. (2024b). Fewer works are dedicated to image-to-text vision LLMs Zong et al. (2024); Zhang et al. (2024); Xue et al. (2024).
Zheng et al. (2024a) conducted an empirical analysis of both position and token biases and proposed PriDe, a debiasing method for LLMs that separates prior bias from unbiased results using shuffling.
We also treat the selection bias as subtraction from observed results in order to obtain the debiased outputs, but we move away from shuffling and regard bias as a vector projected onto three planes. Thus, our calibration method aligns more closely with Zhang et al. (2024), who cut off the visual input for the image-to-text MCQA. Differently, we operate on video-to-text, and, going further, decompose the bias vector into several projections rather than one. This way we eliminate the unimodal bias highlighted by Zhang et al. (2023b).
Wang et al. (2024a) and Liu et al. (2023) resolve the selection bias by augmenting datasets with re-ordered and additional options, increasing number of question-answer pairs in the dataset. We, in contrast, manipulate answer options, as well as videos and questions, not as a solution but as a means of analysis: we identify where the bias is most pronounced and use these modifications as a shortcut for debiasing.
Lastly, Pezeshkpour and Hruschka (2023) propose calibrating LLM predictions by redistributing probabilities to empirically determined options based on the model’s uncertainty. However, we find this approach insufficiently robust, as, according to our analysis, the bias depends on both the model and the dataset, and there is no consistent correlation between the model confidence and the presence of bias.
3 Empirical Analysis
3.1 Datasets
Table 1: Summary of Question-Answer (QA) pairs, Videos, and Number of Options for each dataset.
Dataset QA pairs Videos Options
NExT-QA 8564 1000 5
NExT-GQA 4962 971 5
STAR 7098 914 4
Video-MME 2700 900 4
Perception Test 7656 3926 3
Our goal is to investigate VLMs’ bais under different conditions. To this end, we chose datasets specifically to cover a wide variety of question types, including what, when, why, how, and yes/no questions, that represent a broad spectrum of video-specific reasoning categories such as causal, temporal, descriptive, situational reasoning, as well as memory, abstraction, physics, and semantics.
Table 2: Inference Accuracy (Acc) and F1 Mean Score by Models and Datasets.
Model NExT-QA STAR Perception Test Video-MME
Acc F1 Mean Acc F1 Mean Acc F1 Mean Acc F1 Mean
Video-LLaMA 40.85 40.85 36.59 31.86 41.59 37.19 32.67 28.15
Video-LLaVA 49.96 49.81 34.71 31.83 40.73 35.69 34.22 30.99
SeViLA 63.78 63.88 46.28 46.14 45.30 45.11 39.85 39.82
We selected the following four video MCQA datasets: NExT-QA Xiao et al. (2021) (and its subset NExT-GQA Xiao et al. (2023) with timestamp annotations to locate the answer moment), STAR Wu and Yu (2021), Perception Test Puatruaucean et al. (2023), and Video-MME Fu et al. (2024). These datasets consist of videos of varying lengths, from a few seconds to over an hour. The questions are either manually crafted or generated via functional programs, and the video content range from short everyday activities to animations.
The datasets also differ in the number of answer options per question. Table 1 gives the overview of the number of QA pairs, videos, and answer options for each dataset. Further details on each dataset, including QA examples, are provided in Appendix A.
For experimenting, we used the test splits of NExT-QA/NExT-GQA and Video-MME and the publicly available validation splits with correct answer annotations of STAR and Perception Test.
3.2 Models
To explore selection bias in a broader context, we chose to examine it with respect to diverse model architectures. Despite the differences, the selected models share key characteristics: they sample frames to extract visual features, either by keyframe selection or uniform sampling, and align visual features with textual representations using pre-trained LLMs. Each model leverages pre-trained visual encoders and fine-tunes pre-trained components with video-text datasets. The models differ substantially in methods of integrating video and language modalities Bordes et al. (2024):
Figure 2: Video-LLaMA option distribution for accuracy-irrelevant settings across tested datasets. All Each represents the aggregated distribution under the All Identical Answers setting.
Video-LLaMA Zhang et al. (2023a) has an additional audio channel and uses three query transformers (Q-formers) trained to align the video, audio, and language modalities.
Video-LLaVA (Lin et al., 2023a) aligns videos and LLMs within a shared representation space and has improved multimodal instruction-following capabilities.
SeViLA Yu et al. (2023), specifically designed for question answering, implements a self-chaining approach with Localizer and Answerer: the first selects relevant frames, which the second uses to generate or choose answers.
Table 2 presents the zero-shot inference results of each model on the chosen datasets. Further details on the models, their settings, and prompts can be found in Appendix B.
3.3 Experimental Settings
Eleven dataset modifications are used to examine the role of each component in MCQA benchmarks: video, question, and answer options. The outmost purpose of the modifications is to identify 1) if bias is universal across models and datasets, 2) in what option(s) bias is, 3) how much each component contribute to the bias, and 4) in what settings it is the strongest with respect to each video-MCQA component.
Video Modifications
alter the video input:
1. Correct Frames: Only frames containing the answer are given as video input. This tests if reducing extraneous visual information improves performance. The modification applies only to NExT-QA (NExT-GQA) and STAR due to the availability of timestamp annotations.
2. Empty Frames: Black frames are passed instead of meaningful video input. This tests the model’s response to the absence of visual information within the multimodal setting.
Question Modifications
adjust the questions in two ways:
3. Rephrased Questions: Llama3 Dubey et al. (2024) rephrases the original questions five times, then one rephrasing is randomly selected. [^3] A manual check of 10% of each rephrased dataset confirmed that meaning was retained, with minimal impact on accuracy due to meaning loss. Notably, very few hallucinations were spotted in the NExT-QA rephrased copy. By changing the wording, we check if the selection bias is linked to the token bias.
4. Empty Questions: Following Balepur et al. (2024a), we replace the question with an empty string to test if models can infer the question from answer choices and video. [^4] The previous research for LLMs Balepur et al. (2024b) has shown surprisingly high accuracy in this scenario. We investigate if the video modality contributes to even more meaningful predictions.
Answer Option Modifications
examine the role of answer and its position in the emergence of bias:
5. Answer Shuffling: Options are randomly shuffled once. Previous experiments Zheng et al. (2023); Zong et al. (2023) showed that the models are not robust against changing the option order. Regardless of accuracy, we check if the distribution of answers in each option is the same, which reveals the bias consistency.
6. Correct Answer in Each Option: The correct answer is placed in a fixed position; its content is swapped with another option. By changing the position of correct answer we test to what extend each option contribute to the models’ choice.
7. Correct Answer with Shuffling: The correct answer is fixed, and the remaining options are shuffled. This fusion of the previous two approaches tests how stable a particular option choice is given that it is correct.
8. Additional Empty Option: A new empty option is added in the end to check for overfitting to specific options. If the added blank line is predicted, this indicates a biased distribution of predictions among a certain number of options.
9. All Identical Answers: All options are identical testing for deviation from uniform distribution. When the answer from each option becomes the answer of all options, the answer bias with respect to tokens is eliminated. Moreover, since there is no correct answer in the case of distractors, we can check the answer choice that happens without meaningful reasoning.
10. All Correct Answers: Every option is correct. In addition to (9), where any answer is equal, in this setting any answer is valid. The expectation is to get the consistent uniform distribution.
11. Empty Answers: Each answer option is an empty string. With a non-existent answer, the option name and its position in the question are the only sources for choosing an answer.
To reduce token bias, we use less common option names a0, a1, a2, etc. instead of the conventional A, B, C, etc.
Modifications can also be categorized as accuracy-relevant (e.g., correct frames, shuffling, or rephrased questions) and accuracy-irrelevant (e.g., identical answers or empty parts). Further details, including accuracy results, confusion matrices and examples for each modification are provided in the Appendix C.
Figure 3: Video-LLaVA option distribution for accuracy-irrelevant settings across tested datasets. All Each represents the aggregated distribution under the All Identical Answers setting.
3.4 Observations
Accuracy-relevant settings (Correct Frames, Rephrased Questions, and Answer Shuffling) vary in accuracy but yield almost the same option distributions as the default setting. This suggests that models rely more on answer position than on visual or textual content.
When the correct answer is placed in each option (Correct Answer in Each Option) or combined with shuffling (Correct Answer with Shuffling), the selection rate for that option increases, which indicates adequate reasoning. Moving the correct answer into less favored positions increases their selection proportion, and adding shuffling to it has minimal additional effect.
In contrast, accuracy-irrelevant settings clearly expose bias. Under identical options (All Identical Answers and All Correct Answers) or incomplete conditions (Empty Frames, Empty Questions, Empty Answers), the predicted distributions deviate significantly from uniformity. Figures 2-4 highlight the patterns of the accuracy-agnostic settings as they are indicative for the study of bias, while full performance details are in Appendix C.
In case of All Identical Answers each-option answer showed consistently a very close distribution, so in the figures the result of All Each is obtained by averaging the outcomes across every option experiments within this setting. Tall bars concentrated in certain options indicate bias, while persistently low bars reveal neglected options. Each model exhibits pronounced biases toward specific options.
Video-LLaMA (Figure 2) strongly favors a1 and tends to ignore options beyond the third one. In the Perception Test (three options), it already shows bias against the third option. With identical answers, a0 sometimes competes with a1, and when answers are empty, a2 surprisingly surpasses a1 in Video-MME. Empty settings cover the behavior of the bias to the greatest extent.
Video-LLaVA (Figure 3) behaves similarly but exhibits even stronger bias toward a1, largely ignoring higher-numbered options. Unlike Video-LLaMA, it overwhelmingly favors a1 when answers are empty. With empty questions, a0 begins to compete with a1, especially in Video-MME.
Figure 4: SeViLA option distribution for accuracy-irrelevant settings across tested datasets. All Each represents the aggregated distribution under the All Identical Answers setting.
SeViLA [^5] SeViLA uniquely becomes distracted when an empty option is added, suggesting it may have been trained with fixed option counts. Notably, it shows reduced accuracy for all-correct frames in STAR, which aligns with prior findings Goyal et al. (2016); Lin et al. (2023b), where ”blind” language-only models were able to answer image-based questions. demonstrates greater robustness on NExT-QA, maintaining about uniform 20% distribution under harsh attacks. It only becomes unstable when all correct answers are identical (favoring a0) and with empty answers (favoring the last option). Figure 4 shows slightly higher saturation for these conditions. Unlike Video-LLaMA and Video-LLaVA, which rely on multinomial probabilities, SeViLA uses the argmax function, always selecting the first option if probabilities are equal. This explains why it often favors the first option, a0, in identical-answer scenarios.
4 Methodology
Our bias calibration approach is grounded in established guidelines for developing and validating multiple-choice test items Haladyna (2004), which emphasize balanced, independent, and logically plausible options. This way we ensure that the datasets isolate positional bias and allow our calibration method to be applied effectively. Specifically, we require the following conditions to hold in the MCQA benchmark:
1. Balanced options: All answer options are presented equally without any positional or presentation bias. This is crucial and straightforward to achieve via minimal preprocessing.
2. Independent options: Each option for a given question is independent of others. There should be no overlapping content or logical dependencies among options that influence selection.
3. No logically better or worse options except the correct one: All distractors should be plausible and not trivially inferior or superior to the correct answer. This ensures that the model must rely on genuine reasoning.
4.1 Evaluation Criteria
Our framework both detects bias and evaluates performance. Bias metrics assess model’s consistency in treating all answer options fairly, while performance metrics evaluate accuracy and reliability.
We monitor the standard deviations of the F1 score (F1_std) and the recall (Recall_std). A high F1_std indicates inconsistent performance across options, signaling bias. A high Recall_std suggests the model is better at identifying correct answers for certain options over others.
For option probabilities, we use the standard deviation of the symmetric Jensen-Shannon distance (JS_std) to detect inconsistencies between the predicted and the true probability distributions. This metric, bounded between 0 and 1, is more interpretable than Kullback-Leibler (KL) divergence.
By tracking these metrics, we address group fairness, ensuring the model treats all options equitably. Large standard deviations in any metric suggest bias, which we aim to minimize.
4.2 Debiasing Approach
Figure 5: Bias vector and its projections on the decomposition planes.
Every MCQA benchmark comprises essential components that make the answering task feasible. In the case of video-MCQA, these components include the video, the question, and the answer options. We construct decompositions where the task becomes impossible by design, as one of the key components is removed. This allows us to examine the projections of model’s behavioral bias on each decomposition plane as intuitively shown in Figure 5. In a fair and unbiased model, projecting the dataset by omitting one of the key components would result in a uniform distribution. However, as shown in Section 3, this is far from the case. In the decomposition planes, the bias is more clearly manifested, and since we expect uniformity, the bias becomes easier to detect and mitigate.
4.2.1 Decomposition
Let a Multiple-Choice Question Answering task $T$ be well-defined if there is one and only one correct option to answer the question based on the provided context. Otherwise, the $T$ is ill-defined. We wish to decompose $T$ into the following independent components: video context, question, answer options. Omitting any of these components makes $T$ ill-defined. In general, there could be $n$ independent components $C_{1},C_{2},\dots,C_{n}$. Let $P(T)$ be a well-defining property (correctness) of $T$ such that $P(T)=1$ if $T$ is well-defined, otherwise - $P(T)=0$. We assume that all tasks are initially well-defined.
A model $M$ takes a task $T$ as input and outputs an option ID corresponding to the correct answer for a video-based question: $M(T)=ID_{opt}$.
For an ill-defined task, a model that is both reasonable and unbiased should uniformly select any option ID. Formally, if $P(T)==0$: $M(T)=\text{RandomChoice}(\{ID_{opt}\})$.
The key insight from the analysis in Section 3 is if the model has bias on a given dataset, this bias is likely to be more ubiquitous, hence evident, in the ill-defined versions of the tasks.
Let $A$ be an attack on the well-defined task that removes one of the task component $C_{i}$ that makes this task defined insufficiently, i.e. $A(T)=\hat{T}$ and $P(\hat{T})=0$ while initially $P(T)=1$. Let $A(D)$ be an attack on dataset $D$ which means applying element-wise the same attack $A$ to each $T$ from $D$. Subsequently, we consider three types of attacks that make the task ill-defined: $A_{v=0}$ for all zeroed frames in the video, $A_{q=0}$ for all questions as an empty string, and $A_{o=0}$ for all options as IDs only.
We investigate the importance of the following components of the bias $Bias_{M}$ : $Bias_{M}|_{A_{v=0}(D)}$, $Bias_{M}|_{A_{q=0}(D)}$, $Bias_{M}|_{A_{o=0}(D)}$, and how efficiently we can handle the general bias by balancing through these components.
4.2.2 Debiasing
[…]
$P_{o}(d_{i}|T)=\frac{1}{Z_{T}}P_{p}(d_{i}|T)\times P_{d}(d_{i}|T)$ (1)
where $Z_{T}$ is a normalization factor.
The baseline method BOLD (Bias Optimisation Leveraging Decomposition) estimates the prior bias as following:
$\tilde{P}_{p}(d_{i}|T)=softmax\left(\sum_{j}{P_{p}(d_{i}|A_{j}(T)})\right)$
First, we estimate the model bias on the dataset after attack A. We derive the $P_{p}$ knowing that $P_{d}$ is uniformly distributed from Equation  1:
$P_{p}(d_{i}\mid A(T))=P_{o}(d_{i}\mid A(T))$
Table 3: Comparison of BOLD and Weighted_BOLD bias mitigation approaches across models and datasets for performance and bias monitoring metrics with $k$ = 0.5. Green arrows indicate improvements: upward for Accuracy and F1_mean, and downward for standard deviation metrics. Red arrows represent deterioration, respectively.
Model Dataset Configuration Performance Metrics Bias Monitoring Metrics
Accuracy F1_mean Recall_std F1_std JS_std
NExT-QA BOLD 45.88 (2.43%) 42.15 (3.18%) 22.98 (3.53%) 15.53 (2.11%) 15.55 (3.31%)
Weighted_BOLD 45.91 (2.51%) 42.2 (3.29%) 22.83 (4.19%) 15.51 (2.2%) 15.56 (3.27%)
Video-LLaMA STAR BOLD 37.19 (1.82%) 33.02 (3.65%) 22.29 (7.5%) 14.55 (5.75%) 14.58 (4.66%)
Weighted_BOLD 37.34 (2.24%) 33.5 (5.16%) 21.13 (12.3%) 14.05 (8.98%) 14.14 (7.54%)
Perception Test BOLD 41.9 (1.24%) 38.68 (4.01%) 24.05 (11.11%) 9.87 (13.45%) 14.45 (10.01%)
Weighted_BOLD 42.06 (1.62%) 39.36 (5.86%) 21.8 (19.45%) 9.11 (20.06%) 13.02 (18.94%)
Video-MME BOLD 32.73 (2.13%) 29.23 (3.85%) 19.29 (6.45%) 11.96 (4.46%) 13.1 (4.73%)
Weighted_BOLD 32.2 (0.47%) 28.98 (2.94%) 17.65 (14.38%) 11.69 (6.56%) 12.64 (8.04%)
NExT-QA BOLD 51.81 (3.72%) 51.71 (3.82%) 13.27 (18.63%) 2.67 (18.42%) 4.58 (15.06%)
Weighted_BOLD 52.15 (4.39%) 52.04 (4.49%) 12.72 (22.02%) 2.62 (19.83%) 4.49 (16.83%)
Video-LLaVA STAR BOLD 37.21 (7.01%) 35.76 (12.37%) 18.28 (26.85%) 4.97 (27.15%) 4.54 (21.39%)
Weighted_BOLD 37.53 (7.94%) 36.18 (13.69%) 17.45 (30.15%) 4.91 (28.03%) 4.26 (26.24%)
Perception Test BOLD 41.62 (2.22%) 38.69 (8.4%) 21.75 (20.9%) 10.92 (25.97%) 3.78 (28.4%)
Weighted_BOLD 41.95 (3.02%) 39.46 (10.58%) 19.73 (28.26%) 10.24 (30.53%) 3.4 (35.51%)
Video-MME BOLD 34.7 (1.19%) 32.79 (5.81%) 18.19 (24.51%) 6.16 (23.63%) 3.84 (19.93%)
Weighted_BOLD 34.63 (0.97%) 32.97 (6.38%) 16.36 (32.07%) 6.01 (25.43%) 3.43 (28.5%)
NExT-QA BOLD 63.92 (0.02%) 63.89 (0.02%) 2.03 (5.47%) 1.18 (10.4%) 1.99 (0.17%)
Weighted_BOLD 63.93 (0.04%) 63.91 (0.04%) 1.99 (7.64%) 1.19 (9.83%) 1.99 (0.04%)
SeViLA STAR BOLD 46.22 (0.12%) 46.1 (0.08%) 4.13 (7.46%) 2.26 (1.78%) 2.1 (6.64%)
Weighted_BOLD 46.2 (0.18%) 46.08 (0.13%) 4.01 (10.17%) 2.2 (4.36%) 2.07 (8.2%)
Perception Test BOLD 45.32 (0.06%) 45.18 (0.16%) 5.18 (16.61%) 2.95 (2.43%) 1.3 (18.42%)
Weighted_BOLD 45.31 (0.03%) 45.2 (0.21%) 4.56 (26.58%) 2.83 (6.43%) 1.08 (31.92%)
Video-MME BOLD 40.19 (0.84%) 40.17 (0.88%) 4.11 (12.03%) 1.41 (16.19%) 0.73 (12.52%)
Weighted_BOLD 40.04 (0.46%) 40.03 (0.54%) 3.78 (19.14%) 1.44 (14.6%) 0.69 (17.22%)
To estimate the bias under $A_{v=0}$, $A_{q=0}$ and $A_{o=0}$ attacks, we first compute $P_{p}(d_{i}|A_{v=0}(T))$, $P_{p}(d_{i}|A_{q=0}(T))$ and $P_{p}(d_{i}|A_{o=0})$, and then the prior estimation, respectively: $\tilde{P}_{p}(d_{i}|T)=softmax[P_{p}(d_{i}|A_{v=0}(T))+P_{p}(d_{i}|A_{q=0}(T))+P_{p}(d_{i}|A_{o=0}(T))]$.
According to Zheng et al. (2024b), biases related to option IDs may transfer effectively across different samples and domains. Therefore, finding the global prior with $K$ samples, we can mitigate the bias for the entire dataset with minimal efforts.
From the dataset $D$, we take $K=k*||D||$ samples, denoted as $D_{k}$, where $k$ is adjusted coefficient based on estimation budget. Each sample in $D_{k}$ is subjected to the set of attacks $A$. For each, we calculate the sample-specific prior $P_{p}(d_{i}|A(T))$ by Equation 3 and estimate the adjusted prior $\tilde{P}_{p}(d_{i}|T)$ by Equation 2. The global prior, $\tilde{P}_{\text{p}}(d_{i})$, is then obtained by averaging the priors for each sample.
Ultimately, the posterior debiased through the global prior is calculated as follows:
$P_{d}(d_{i}|T)=softmax\left(\log P_{o}(d_{i}|T)-\log\tilde{P}_{p}(d_{i})\right)$
In the extended version, Weighted_BOLD, we use weights for $P_{p}(d_{i}|A(T))$:
$\tilde{P}_{p}(d_{i}|T)=softmax\left(\sum_{j}{w_{j}P_{p}(d_{i}|A_{j}(T)})\right)$
This allows for a more optimal direction in the latent space of the priors, as we can prioritize certain priors by significance and refine the prior estimation $\tilde{P}_{p}(d_{i}|T)$.
To gain the optimal weights, we use a 5-fold cross-validation procedure. In each fold, we optimize weights $\{w_{i}\}$ using the COBYLA algorithm Powell (1994), subject to the constraints $0\leq w_{i}\leq 1$ or $|w_{i}|\leq 1$.
The restriction $0\leq w_{i}\leq 1$ implies that the resulting debiasing vector is formed as a convex combination of the prior directions, remaining within the positive span of the original vectors. With this easy-to-interpret setup, we analyze how much each prior contributes to the final debiasing vector. Notably, a weight distribution with $w_{i}=0$ enables ablation-style analysis by turning off the corresponding debiasing component. Uninformative or detrimental priors are automatically suppressed, resulting in a more streamlined and potentially more generalizable debiasing vector. When we allow the more relaxed constraint $|w_{i}|\leq 1$, the COBYLA optimizer explores the full linear subspace spanned by the prior vectors, and the solution is no longer restricted to the positive orthant. This provides additional flexibility to correct for complex biases that may not be adequately addressed by additive combinations alone.
While we minimize the standard deviation of Recall_std across the options on the test set, we monitor the same bias metrics on the validation set. After completing the cross-validation we average the estimated global priors across the folds and apply this averaged global prior. Both algorithms, BOLD and Weighted_BOLD, are outlined in Appendix D.
5 Results
We conducted tests with the $k$ sample parameter set to 0.25, 0.5, 0.75, and 1 (the entire dataset) in combination with positive and negative weights. The configuration with $k$=0.5 and $w_{i}\geq 0$ provided the optimal balance between sample size and performance. As shown in Table 3 with results for positive weights and $k$=0.5, [^6] Additional results for $k$ values of 0.25, 0.75, and 1 combined with different weights are provided in Appendix D. the overall significant improvement in bias metrics, including Recall_std, indicates that our method effectively reduces option disparity.
Weighted_BOLD generally outperforms BOLD because the weights capture critical priors in the latent space. Thus, they prove to be effective for fine-tuning bias correction within the decomposed priors.
Both BOLD and Weighted_BOLD consistently improve performance metrics for Video-LLaMA and Video-LLaVA. SeViLA generally shows more resilience to bias calibration, with minimal effect on performance on NExT-QA and marginal deterioration on STAR. This may be attributed to the model’s training for statistically close to uniform distribution of results over a fixed options number. Despite decreases in Accuracy and F1 Mean on STAR, the reduction in Recall_std across all datasets for this model suggests that its debiased responses are more truthful, reflecting a fairer distribution of correct answers across options.
Video-LLaVA, which exhibits the most pronounced selection bias, benefits the most from bias reduction via prior decomposition across all datasets, achieving a maximum accuracy gain of 7.94% on STAR with the weighted configuration.
6 Conclusion
We presented the first comprehensive analysis of selection bias in video-based MCQA tasks, using modified datasets to pinpoint how bias emerges within specific answer options and across different VLM architectures. Our findings reveal that while all models share common tendencies in answer distribution, they also possess distinct bias profiles both across and within datasets. Notably, the bias vector is consistently more pronounced when projecting onto the decomposed planes where one key MCQA component (video, question, or answer options) is removed.
Building on these insights, we proposed the BOLD calibration method using a global prior bias vector, derived from the decomposed versions of the task, to mitigate selection bias. Our technique requires only three decompositions per question and is thus resource-efficient. Our BOLD bias removal not only enhances fairness and interpretability but also improves performance metrics. Extending the method to Weighted_BOLD refines the balance between debiasing and performance gains.
This easily scalable, post-processing, automatic, and unsupervised method is applicable to any existing VLM and allows for a sharper focus on its true reasoning capabilities. By making MCQA benchmarks more robust, transparent, and impartial, we enable a clearer assessment of the genuine reasoning cues in VLMs and pave the way for future improvements in model design and training strategies.
Furthermore, elevating bias suppression to an abstract level, where tasks are systematically decomposed and key components are removed to identify bias, we offer a framework beyond the current scope to address other bias-related challenges in machine learning models.
7 Limitations
Our approach rests on the assumption that when models lack sufficient direct information, they rely on indirect cues, manifesting as selection bias. While we have shown that bias can be decomposed by selectively removing the video, question, or answer component, this decomposition may not be exhaustive. There could be more optimal directions in the latent space of decomposed priors. Similarly, although our method identifies pronounced bias when all answers are correct, there remain scenarios where further supervised debiasing (e.g., applying COBYLA optimization with true labels) might yield more optimal results. We aimed for an easy-to-implement, unsupervised solution without relying on ground-truth answers.
We also assume the validity of uniform distribution as a baseline: Jensen-Shannon Divergence operates under the assumption that, in the absence of prior information, the distribution should be uniform. If the test MCQA data or task suggest an uneven distribution of answer options, this may lead to underestimating the significance of certain factors.
Another limitation is that we primarily assessed our post-processing calibration on the original datasets rather than systematically verifying performance improvements across all modifications. While our current evaluation covers four challenging video datasets and three model architectures, more granular, scenario-specific testing would provide stronger evidence of robustness. However, performing comprehensive evaluations on every modification poses practical challenges. These limitations point toward future work: exploring more nuanced testing scenarios, investigating methods that do not rely on uniform reference distributions, and potentially leveraging additional computational resources or innovative techniques for even more thorough bias and performance assessments.
8 Ethics Statement
We based our experiments on existing datasets, ensuring compliance with their respective copyrights. Our research uses open-source VLMs, which, while offering accessibility and flexibility, also carry the inherent risks associated with open-ended text generation.
Acknowledgements
Olga Loginova thanks Amazon Alexa for supporting her research through a generous donation to Raffaella Bernardi. Ravi Shekhar has been partially funded by the UKRI and the European Union. The views and opinions expressed are, however, those of the authors only and do not necessarily reflect those of the European Union or Research Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. All authors are deeply grateful to Giovanni Bonetta for his invaluable contributions and insightful consultations at earlier stages of the project.
Appendix A Datasets
A.1 NExT-QA and NExT-GQA
NExT-QA, released in 2021, is one of the first datasets designed to assess a VLM’s ability to capture causal, temporal, and descriptive topics through what/when/why/how and before/after questions. It comprises videos sourced from the VidOR dataset Shang et al. (2019), with an average length of 35.73 seconds. The test split primarily contains causal questions (52.57%) addressing why and how, temporal reasoning questions (31.03%) about preceeding, following, or concurrent actions, and descriptive tasks (16.04%). Videos are manually annotated with questions and correct answers. Distractors are selected from 50 candidates based on a cosine similarity of less than 0.9 to the correct answer, ensuring they are close but not identical. This approach sometimes allows questions to be meaningfully answered without viewing the video (see Figures  68).
Since our models were released after 2021, they might have included NExT-QA in their training or fine-tuning data.
NExT-GQA, released in 2024 as an extension of NExT-QA, focuses on evaluating the localization abilities of VLMs. It includes manually annotated timestamp labels that are key for answering the questions. Descriptive questions are excluded because they pertain to global video content or have answers available throughout the video with no specific temporal grounding. We also excluded QA pairs with multiple annotated time intervals. Consequently, the test set of NExT-GQA comprises 56.69% causal questions and 43.31% temporal questions.
Figure 6: Example of a QA pair from NExT-QA in the Causal category. The correct answer is in the box.
Figure 7: Example of a QA pair from NExT-QA in the Temporal category. The correct answer is in the box.
Figure 8: Example of a QA pair from NExT-QA in the Descriptive category. The correct answer is in the box.
A.2 STAR
The STAR dataset is derived from the Action Genome dataset Ji et al. (2019), which itself is based on Charades Yuan et al. (2017). Charades comprises indoor activity videos with an average length of 30 seconds. Action Genome provides annotations of objects and their relationships within these videos.
The question-answer pairs are generated using functional programs based on situation hyper-graphs: the graphs represent the extraction of abstract representations from the videos, while the functional programs provide multiple types of questions that cover different levels of difficulty in situated reasoning. There are four types of questions focusing on occurred facts, temporal order, future probability, and feasibility in specific situations:
  • Sequence (50.53%): What did the person do before/after …?”
  • Interaction (33.78%): “What did a person do …?”
  • Prediction (8.79%): “What will the person do next with …?”
  • Feasibility (6.9%): “What is the person able to do?” or “Which object is possible to be …?”
All QA pairs are annotated with a timestamp label. To avoid answer distribution bias, each question type is balanced through resampling. To prevent reasoning shortcuts due to frequent action or entity co-occurrences and to ensure models cannot easily answer without genuine understanding, the compositionality of verbs and nouns is controlled.
Examples are shown in Figures  912.
Figure 9: Example of a QA pair from STAR in the Sequence category. The correct answer is in the box.
Figure 10: Example of a QA pair from STAR in the Interaction category. The correct answer is in the box.
Figure 11: Example of a QA pair from STAR in the Prediction category. The correct answer is in the box.
Figure 12: Example of a QA pair from STAR in the Feasibility category. The correct answer is in the box.
A.3 Perception Test
The Perception Test dataset consists of four categories designed to evaluate VLMs on different types of reasoning:
  • Physics (36.26%)
  • Semantics (30.54%)
  • Abstraction (29.28%)
  • Memory (3.92%)
These categories test descriptive, explanatory, predictive, and counterfactual reasoning abilities. The video scripts are based on human perception screening tests, with an average length of 23 seconds. All answers are crowd-sourced. Distractors are both manually created and automatically generated, particularly for standard closed yes/no questions.
We use a 40% split of the correct-answer-annotated validation set to save the resources. Sampling is stratified to maintain the original balance of question types and the correct answer distribution. The split contains approximately 15% closed questions, which poses an additional challenge for models that tend to exhibit affirmation bias Pezeshkpour and Hruschka (2023). Notably, for closed questions in this dataset, Video-LLaMA and Video-LLaVA often respond with "yes" or "no" rather than selecting an answer option.
Examples of QA pairs from the Perception Test are shown in Figures  1316.
Figure 13: Example of a QA pair from Perception Test in the Descriptive Physics category. The correct answer is in the box.
Figure 14: Example of a QA pair from Perception Test in the Descriptive Semantics category. The correct answer is in the box.
Figure 15: Example of a QA pair from Perception Test in the Counterfactual Abstraction category. The correct answer is in the box.
Figure 16: Example of a QA pair from Perception Test in the Explanatory Memory category. The correct answer is in the box.
A.4 Video-MME
Video-MME was developed to address the lack of diversity in video MCQA benchmarks. It consists of 900 YouTube videos ranging from 11 seconds to 1 hour in length. The dataset covers 6 primary visual domains with 30 subfields and 12 task types including questions on temporal and spatial perception, object and action reasoning, OCR, and counting problems. All QA pairs and distractors are manually created with three questions per video. Additionally, the dataset features videos in languages other than English. The videos additionally are supplied with subtitles.
The released test split has an imbalanced option distribution, so we equalized it to properly assess option bias.
Examples are shown in Figures  1720.
Figure 17: Example of a QA pair from Video-MME in the Temporal Reasoning category. The correct answer is in the box.
Figure 18: Example of a QA pair from Video-MME in the OCR Problems category. The correct answer is in the box.
Figure 19: Example of a QA pair from Video-MME in the Counting Problem category. The correct answer is in the box.
Figure 20: Example of a QA pair from Video-MME in the Spatial Perception category. The correct answer is in the box.
Appendix B Models’ Settings and Prompts
B.1 Video-LLaMA
Video-LLaMA aligns a BLIP-2-based visual encoder Li et al. (2023) with LLaMA Touvron et al. (2023). It incorporates audio signals using Imagebind Girdhar et al. (2023). Both the video and audio Q-formers are trained on the Webvid-2M dataset Bain et al. (2021) to align the modalities. LLaMA is fine-tuned with visual instructional data.
Video-LLaMA is designed as a conversational model and was likely not specifically trained on MCQA, but rather on open-ended VQA datasets to generate captions, descriptions, and explanatory answers. Consequently, throughout our experiments with different prompt options, it tends to engage in dialogue: it repeats the question, responds with the general phrase "Hello, how can I help you today?", describes the video content, or even asks follow-up questions such as "What do you think about the video?" rather than providing a direct answer from a predefined list of options or option IDs. [^7] It is noteworthy that the model often uses the following template to explain its answer choice: ”The ID of the correct answer: $a_{i}$. Explanation: …” Making it elicit a clear and concise answer option is quite challenging.
To overcome this, we experimented with various prompts, temperature settings, and token limits, and developed the following strategy:
System prompt:
Analyse the video, choose the correct answer in multiple choice question answering. Give the ID of the correct answer option from the given list of options. Be concise.
User prompt:
<question> Options: <options>. The ID of the correct answer:
We set the number of beams to 2 and the temperature to 0.8. Additionally, we allowed the model up to 30 attempts to produce an answer that includes an answer ID. We used the 7B parameter model due to resource constraints. For answering, the model selects eight uniformly sampled frames.
B.2 Video-LLaVA
Video-LLaVA (Lin et al., 2023a) uses the pre-trained CLIP ViT-L/14 vision encoder Radford et al. (2021) and the Vicuna language model Chiang et al. (2023). Visual features from the encoder are projected into the same-dimensional space as the textual embeddings via a linear projector.
Video-LLaVA effectively delivers concise answers as just a single option based on a short, general prompt. However, in some cases, the model outputs as an answer option only "A" or responds with "yes" or "no" to a closed question.
For this model, we used the following prompt:
<video_tag> Question: <question> Options: <options> The ID of the correct answer is: Respond with only the correct answer ID. ASSISTANT:
Notably, on the completely unfamiliar datasets Perception Test and Video-MME adding an empty option increases the number of responses that cannot be parsed as an answer ID (although this increase is statistically insignificant). Combined with the observation that the model tends to disregard options presented later in the prompt, this suggests that the length of the question affects the model’s performance, and tokens closer to the end of the prompt have a lower probability.
The model also selects eight frames from the video by uniform sampling. To conserve computational resources, we used Video-LLaVA-7B with half-precision set by Torch’s autocast. The temperature is equal to 1.
B.3 SeViLA
SeViLA adopts a selfchaining approach with BLIP-2 Li et al. (2023) for keyframe localization and answering question. Its Localizer uniformly samples 32 frames and selects the 4 most relevant key ones, while the Answerer generates responses based on those frames. The Localizer and Answerer both have Flan-T5 Chung et al. (2022) for LLM.
SeViLA does not have any unanswered or unparsable responses because it was specifically tailored for the multiple-choice VQA task: internally, it maps any answer option IDs to {A, B, C, D, E}. However, it is limited to observing at most six answer options. In a side experiment, we shifted all answers to options G, H, I, J, K leaving the first six options empty, and obtained identical results to our other experimental setting, Empty Answers. This indicates that the model is completely blind to options beyond F and is likely overfitted to the options {A, B, C, D, E}.
Furthermore, the shape of the confusion matrices for NExT-QA and STAR in Figures 21 and 22 shows an unusually accurate understanding of these datasets, unlike those for Perception Test and Video-MME. Since NExT-QA and STAR were released before SeViLA, the model might have been fine-tuned on them. On the other hand, this suggests that the positional bias in this model is much less pronounced.
Another vulnerability of SeViLA is that it uses argmax calculated on logit probabilities. However, logits without numerical stabilization can produce NaN values when activated via softmax. In such cases, according to argmax, the model’s answer defaults to $a_{0}$. Among the settings relevant for debiasing, such behavior was observed in 22 cases on NExT-QA.
For this model, we used the original prompts: Localizer prompt (when the model selects 32 frames):
Does the information within the frame provide necessary details to accurately answer the given question?
Answerer prompt (when the model selects the answer based on 4 key frames):
Considering information in frames, select the correct answer from the options.
The confusion matrices for all models in the default (unmodified) setting in Figures  2123 provide a quick insight into positional bias: darker blue cells indicate a greater accumulation of answers.
Appendix C Models’ Performance on All Settings
C.1 Experimental Settings Examples
As the modifications to the video component do not affect the questions and answers, Table 4 presents examples of adjustments applied only to the question-answer pairs, along with their original versions. The example provided for the Video-MME dataset illustrates analogous changes implemented across all four datasets.
C.2 Models’ Performance on the Settings
Tables 5 through 16 present the distribution of selections across answer options and the accuracy (where applicable) for all settings, along with the target distributions. Notably, as soon as the correct answer aligns with the model’s biased option, the performance boosts.
The full data on distribution answers in each option is given in Figures 25 to 27.
Appendix D Debiasing Algorithms and Results
The BOLD algorithm 1 describes estimating and adjusting for prior probabilities derived from decomposed inputs of ill-defined tasks. The Weighted_BOLD algorithm  2 is an extension of Algorithm  1, this method introduces weighting to the prior estimation to enhance debiasing by better capturing critical priors.
We fixed the random seed to 1 and conducted a series of experiments varying the weights and the sampling parameter $k$. All results including the default values are presented in Tables  1720.
When considering only positive weights ($w_{i}\geq 0$), we assume that the debiasing vector is oriented correctly and use the weights to fine-tune the magnitude of its decomposed constituents. Therefore, positive weights generally enhance improvements and aggravate deteriorations. Specifically, in this case, all the primary debiasing metrics improve along with Recall_std, albeit at different rates across models and datasets.
When we optimize a single bias metric, i. e. Recall_std, by allowing negative weights, we search the latent subspace for the most optimal vector to improve that specific metric. However, this optimization may not correlate with an improvement in the overall bias, as it focuses solely on that metric. Consequently, using negative weights can lead to deterioration in other bias and performance metrics. This observation underscores the importance of evaluating bias using multiple metrics.
We provide the results for both cases, when weights are constrained to $0\leq w_{i}\leq 1$ and when $|w_{i}|\leq 1$, as we believe that the observation of the direction of the bias vectors is crucial for understanding the overall debiasing effect.
Table 4: Questions and answers modifications example from MME dataset. Correct answer is highlighted in bold in the “Default” modification type.
Modification Type Question Answers
Default How many national flags appear in the video? (a0) 3. (a1) 5. (a2) 2. (a3) 4.
Empty answers Default question (a0) – (a1) – (a2) – (a3) –
Answer shuffling Default question (a0) 4. (a1) 5. (a2) 2. (a3) 3.
Additional empty option Default question (a0) 3. (a1) 5. (a2) 2. (a3) 4. (a4) –
Correct answer in each option Default question (a0) 4. (a1) 4. (a2) 4. (a3) 4.
All identical answers (a0) Default question (a0) 3. (a1) 3. (a2) 3. (a3) 3.
All identical answers (a1) Default question (a0) 5. (a1) 5. (a2) 5. (a3) 5.
All identical answers (a2) Default question (a0) 2. (a1) 2. (a2) 2. (a3) 2.
All identical answers (a3) Default question (a0) 4. (a1) 4. (a2) 4. (a3) 4.
Correct answer in (a0) Default question (a0) 4. (a1) 5. (a2) 2. (a3) 3.
Correct answer in (a1) Default question (a0) 3. (a1) 4. (a2) 2. (a3) 5.
Correct answer in (a2) Default question (a0) 3. (a1) 5. (a2) 4. (a3) 2.
Correct answer in (a3) Default question (a0) 3. (a1) 5. (a2) 2. (a3) 4.
Correct answer (a0) with shuffling Default question (a0) 4. (a1) 5. (a2) 3. (a3) 2.
Correct answer (a1) with shuffling Default question (a0) 5. (a1) 4. (a2) 2. (a3) 3.
Correct answer (a2) with shuffling Default question (a0) 3. (a1) 2. (a2) 4. (a3) 5.
Correct answer (a3) with shuffling Default question (a0) 5. (a1) 3. (a2) 2. (a3) 4.
Rephrased question What is the total number of national flags that appear in the video? (a0) 3. (a1) 5. (a2) 2. (a3) 4.
Empty questions (a0) 3. (a1) 5. (a2) 2. (a3) 4.
Figure 21: All Models’ Confusion Matrices for NeXT-QA.
Figure 22: All Models’ Confusion Matrices for STAR.
Figure 23: All Models’ Confusion Matrices for Video-MME.
Figure 24: All Models’ Confusion Matrices for Perception Test.
Figure 25: Video-Llama option distribution for all settings of NeXT-QA, NExT-GQA, STAR, Perception Test and Video-MME. Correct ai, Correct ai Shuffled and All ai represent Correct Answer, Correct Answer with Shuffling and All Correct Answers in each option, respectively.
Figure 26: Video-Llava option distribution for all settings of NeXT-QA, NExT-GQA, STAR, Perception Test and Video-MME. Correct ai, Correct ai Shuffled and All ai represent Correct Answer, Correct Answer with Shuffling and All Correct Answers in each option, respectively.
Figure 27: SeViLA option distribution for all settings of NeXT-QA, NExT-GQA, STAR, Perception Test and Video-MME. Correct ai, Correct ai Shuffled and All ai represent Correct Answer, Correct Answer with Shuffling and All Correct Answers in each option, respectively.
Table 5: Video-LLaMA Performance on All Settings of NExT-QA (including NExT-GQA). Each cell shows the count of selections for each option (a0a5), along with the percentage in parentheses. The Accuracy column represents the number of correct answers where applicable. The N/A column indicates the number of cases where the model did not output an answer option after 30 attempts.
Setting a0 a1 a2 a3 a4 a5 N/A Accuracy
Target 1,721 (20.10%) 1,671 (19.51%) 1,767 (20.63%) 1,719 (20.07%) 1,686 (19.69%) 0 (0.00%)
Default 1,430 (16.70%) 3,285 (38.37%) 2,727 (31.85%) 1,002 (11.70%) 117 (1.37%) 3 (0.04%) 3,837 (44.82%)
Answer Shuffling 1,443 (16.97%) 3,295 (38.75%) 2,654 (31.21%) 993 (11.68%) 119 (1.40%) 60 (0.70%) 3,799 (44.67%)
Correct Frames 869 (17.61%) 1,836 (37.20%) 1,588 (32.18%) 587 (11.89%) 55 (1.11%) 27 (0.54%) 2,155 (43.67%)
Rephrased Questions 1,305 (15.36%) 3,270 (38.50%) 2,704 (31.83%) 1,062 (12.50%) 153 (1.80%) 70 (0.82%) 3,762 (44.29%)
Additional Empty Option 1,432 (17.17%) 2,987 (35.82%) 2,067 (24.78%) 1,090 (13.07%) 725 (8.69%) 39 (0.47%) 224 (2.62%) 4,086 (48.99%)
All a1 3,537 (42.71%) 3,891 (46.98%) 737 (8.90%) 32 (0.39%) 85 (1.03%) 282 (3.29%)
All a2 3,614 (43.82%) 3,780 (45.83%) 748 (9.07%) 32 (0.39%) 73 (0.89%) 317 (3.70%)
All a3 3,585 (43.43%) 3,831 (46.41%) 722 (8.75%) 33 (0.40%) 83 (1.01%) 310 (3.62%)
All a4 3,590 (43.35%) 3,837 (46.33%) 733 (8.85%) 34 (0.41%) 87 (1.05%) 283 (3.30%)
Correct a0 4,136 (48.98%) 2,303 (27.27%) 1,489 (17.63%) 453 (5.36%) 64 (0.76%) 119 (1.39%) 4,136 (48.98%)
Correct a0 Shuffled 4,136 (49.00%) 2,304 (27.30%) 1,525 (18.07%) 415 (4.92%) 61 (0.72%) 123 (1.44%) 4,136 (49.00%)
Correct a1 730 (8.61%) 6,323 (74.61%) 1,102 (13.00%) 292 (3.45%) 28 (0.33%) 89 (1.04%) 6,323 (74.61%)
Correct a1 Shuffled 704 (8.30%) 6,348 (74.88%) 1,098 (12.95%) 297 (3.50%) 31 (0.37%) 86 (1.00%) 6,348 (74.88%)
Correct a2 622 (7.32%) 2,325 (27.35%) 5,240 (61.64%) 280 (3.29%) 34 (0.40%) 63 (0.74%) 5,240 (61.64%)
Correct a2 Shuffled 583 (6.88%) 2,356 (27.81%) 5,240 (61.84%) 254 (3.00%) 40 (0.47%) 91 (1.06%) 5,240 (61.84%)
Correct a3 762 (8.99%) 2,315 (27.32%) 2,321 (27.39%) 3,029 (35.74%) 47 (0.55%) 90 (1.05%) 3,029 (35.74%)
Correct a3 Shuffled 766 (9.07%) 2,335 (27.65%) 2,235 (26.47%) 3,049 (36.11%) 59 (0.70%) 120 (1.40%) 3,049 (36.11%)
Correct a4 1,194 (14.11%) 3,014 (35.63%) 2,857 (33.77%) 909 (10.74%) 486 (5.74%) 104 (1.21%) 486 (5.74%)
Correct a4 Shuffled 1,164 (13.77%) 2,912 (34.45%) 2,949 (34.88%) 912 (10.79%) 517 (6.12%) 110 (1.28%) 517 (6.12%)
All Correct Answers 5,247 (63.17%) 2,730 (32.87%) 295 (3.55%) 9 (0.11%) 25 (0.30%) 258 (3.01%)
Empty Frames 1,092 (12.77%) 3,494 (40.87%) 3,344 (39.11%) 544 (6.36%) 76 (0.89%) 14 (0.16%) 2,620 (30.64%)
Empty Questions 1,686 (20.31%) 3,975 (47.89%) 1,676 (20.19%) 786 (9.47%) 178 (2.14%) 263 (3.07%) 3,082 (37.13%)
Empty Answers 987 (11.66%) 4,543 (53.68%) 2,837 (33.52%) 37 (0.44%) 59 (0.70%) 101 (1.18%)
Table 6: Video-LLaMA Performance on All Settings of STAR. Each cell shows the count of selections for each option (a0a4), along with the percentage in parentheses. The Accuracy column represents the number of correct answers where applicable. The N/A column indicates the number of cases where the model did not output an answer option after 30 attempts.
Setting a0 a1 a2 a3 a4 N/A Accuracy
Target 1,755 (24.73%) 1,758 (24.77%) 1,742 (24.54%) 1,843 (25.97%) 0 (0.00%)
Default 1,357 (19.13%) 3,645 (51.38%) 1,982 (27.94%) 110 (1.55%) 4 (0.06%) 2,596 (36.59%)
Answer Shuffling 1,255 (18.72%) 3,458 (51.57%) 1,893 (28.23%) 99 (1.48%) 393 (5.54%) 2,467 (36.79%)
Correct Frames 1,322 (19.43%) 3,455 (50.79%) 1,899 (27.91%) 127 (1.87%) 295 (4.16%) 2,524 (37.10%)
Rephrased Questions 1,376 (20.03%) 3,462 (50.39%) 1,891 (27.53%) 141 (2.05%) 228 (3.21%) 2,520 (36.68%)
Additional Empty Option 1,314 (20.59%) 2,892 (45.31%) 1,494 (23.41%) 630 (9.87%) 52 (0.81%) 716 (10.09%) 2,528 (39.61%)
All a0 2,484 (38.04%) 3,949 (60.47%) 88 (1.35%) 9 (0.14%) 568 (8.00%)
All a1 2,695 (39.68%) 3,976 (58.54%) 104 (1.53%) 17 (0.25%) 306 (4.31%)
All a2 2,711 (39.93%) 3,957 (58.28%) 113 (1.66%) 9 (0.13%) 308 (4.34%)
All a3 2,512 (38.52%) 3,909 (59.94%) 94 (1.44%) 7 (0.11%) 576 (8.11%)
Correct a0 2,340 (34.53%) 2,996 (44.21%) 1,378 (20.33%) 63 (0.93%) 321 (4.52%) 2,340 (34.53%)
Correct a0 Shuffled 2,313 (34.36%) 2,960 (43.97%) 1,392 (20.68%) 67 (1.00%) 366 (5.16%) 2,313 (34.36%)
Correct a1 966 (14.16%) 4,677 (68.54%) 1,124 (16.47%) 57 (0.84%) 274 (3.86%) 4,677 (68.54%)
Correct a1 Shuffled 985 (14.48%) 4,622 (67.93%) 1,136 (16.70%) 61 (0.90%) 294 (4.14%) 4,622 (67.93%)
Correct a2 918 (13.53%) 2,913 (42.93%) 2,898 (42.71%) 56 (0.83%) 313 (4.41%) 2,898 (42.71%)
Correct a2 Shuffled 844 (12.57%) 2,908 (43.32%) 2,901 (43.21%) 60 (0.89%) 385 (5.42%) 2,901 (43.21%)
Correct a3 1,226 (18.03%) 3,298 (48.50%) 2,002 (29.44%) 274 (4.03%) 298 (4.20%) 274 (4.03%)
Correct a3 Shuffled 1,153 (17.21%) 3,190 (47.60%) 2,127 (31.74%) 231 (3.45%) 397 (5.59%) 231 (3.45%)
All Correct Answers 3,129 (47.60%) 3,376 (51.35%) 63 (0.96%) 6 (0.09%) 524 (7.38%)
Empty Frames 1,323 (18.64%) 3,721 (52.42%) 2,047 (28.84%) 7 (0.10%) 0 (0.00%) 1,854 (26.12%)
Empty Questions 2,365 (34.84%) 2,973 (43.80%) 1,275 (18.78%) 175 (2.58%) 310 (4.37%) 2,427 (35.75%)
Empty Answers 1,121 (15.84%) 3,093 (43.72%) 2,778 (39.27%) 83 (1.17%) 23 (0.32%)
Table 7: Video-LLaMA Performance on All Settings of Perception Test. Each cell shows the count of selections for each option (a0a3), along with the percentage in parentheses. The Accuracy column represents the number of correct answers where applicable. The N/A column indicates the number of cases where the model did not output an answer option after 30 attempts.
Setting a0 a1 a2 a3 N/A Accuracy
Target 2,549 (33.29%) 2,561 (33.45%) 2,546 (33.25%) 0 (0.00%)
Default 1,546 (20.34%) 5,281 (69.47%) 775 (10.19%) 54 (0.71%) 3,162 (41.59%)
Answer Shuffling 1,593 (21.10%) 5,165 (68.42%) 791 (10.48%) 107 (1.40%) 3,146 (41.67%)
Rephrased Questions 1,652 (21.83%) 5,004 (66.13%) 911 (12.04%) 89 (1.16%) 3,161 (41.77%)
Additional Empty Option 1,448 (19.56%) 4,356 (58.86%) 1,549 (20.93%) 48 (0.65%) 255 (3.33%) 3,004 (40.59%)
All a0 2,768 (36.60%) 4,414 (58.36%) 381 (5.04%) 93 (1.21%)
All a1 2,842 (37.56%) 4,334 (57.28%) 390 (5.15%) 90 (1.18%)
All a2 2,806 (37.15%) 4,326 (57.28%) 421 (5.57%) 103 (1.35%)
Correct a0 2,245 (29.76%) 4,675 (61.98%) 623 (8.26%) 113 (1.48%) 2,245 (29.76%)
Correct a0 Shuffled 2,271 (30.12%) 4,654 (61.73%) 614 (8.14%) 117 (1.53%) 2,271 (30.12%)
Correct a1 1,204 (15.95%) 5,833 (77.29%) 510 (6.76%) 109 (1.42%) 5,833 (77.29%)
Correct a1 Shuffled 1,233 (16.32%) 5,820 (77.06%) 500 (6.62%) 103 (1.35%) 5,820 (77.06%)
Correct a2 1,239 (16.40%) 5,077 (67.18%) 1,241 (16.42%) 99 (1.29%) 1,241 (16.42%)
Correct a2 Shuffled 1,235 (16.36%) 5,082 (67.30%) 1,234 (16.34%) 105 (1.37%) 1,234 (16.34%)
All Correct Answers 3,330 (43.87%) 4,053 (53.40%) 207 (2.73%) 66 (0.86%)
Empty Frames 990 (12.93%) 5,641 (73.68%) 1,025 (13.39%) 0 (0.00%) 3,005 (39.25%)
Empty Questions 2,564 (34.61%) 3,707 (50.03%) 1,138 (15.36%) 247 (3.23%) 3,021 (40.77%)
Empty Answers 2,969 (39.11%) 2,876 (37.88%) 1,747 (23.01%) 64 (0.84%)
Table 8: Video-LLaMA Performance on All Settings of Video-MME. Each cell shows the count of selections for each option (a0a4), along with the percentage in parentheses. The Accuracy column represents the number of correct answers where applicable. The N/A column indicates the number of cases where the model did not output an answer option after 30 attempts.
Setting a0 a1 a2 a3 a4 N/A Accuracy
Target 675 (25.00%) 675 (25.00%) 675 (25.00%) 675 (25.00%) 0 (0.00%)
Default 474 (17.92%) 1,344 (50.81%) 757 (28.62%) 70 (2.65%) 55 (2.04%) 864 (32.67%)
Answer Shuffling 451 (18.11%) 1,242 (49.86%) 733 (29.43%) 65 (2.61%) 209 (7.74%) 785 (31.51%)
Rephrased Questions 1,367 (50.65%) 979 (36.27%) 198 (7.34%) 155 (5.74%) 1 (0.04%) 873 (32.35%)
Additional Empty Option 391 (17.29%) 1,040 (45.98%) 624 (27.59%) 178 (7.87%) 29 (1.28%) 438 (16.22%) 746 (32.98%)
All a0 1,042 (42.81%) 1,244 (51.11%) 120 (4.93%) 28 (1.15%) 266 (9.85%)
All a1 1,027 (42.00%) 1,295 (52.97%) 91 (3.72%) 32 (1.31%) 255 (9.44%)
All a2 1,061 (43.36%) 1,232 (50.35%) 116 (4.74%) 38 (1.55%) 253 (9.37%)
All a3 1,001 (41.06%) 1,223 (50.16%) 149 (6.11%) 65 (2.67%) 262 (9.70%)
Correct a0 678 (27.19%) 1,141 (45.75%) 624 (25.02%) 51 (2.04%) 206 (7.63%) 678 (27.19%)
Correct a0 Shuffled 699 (27.90%) 1,085 (43.31%) 664 (26.51%) 57 (2.28%) 195 (7.22%) 699 (27.90%)
Correct a1 376 (14.98%) 1,532 (61.04%) 547 (21.79%) 55 (2.19%) 190 (7.04%) 1,532 (61.04%)
Correct a1 Shuffled 344 (13.78%) 1,519 (60.83%) 584 (23.39%) 50 (2.00%) 203 (7.52%) 1,519 (60.83%)
Correct a2 366 (14.63%) 1,147 (45.86%) 948 (37.90%) 40 (1.60%) 199 (7.37%) 948 (37.90%)
Correct a2 Shuffled 355 (14.27%) 1,144 (45.98%) 942 (37.86%) 47 (1.89%) 212 (7.85%) 942 (37.86%)
Correct a3 408 (16.47%) 1,178 (47.56%) 777 (31.37%) 114 (4.60%) 223 (8.26%) 114 (4.60%)
Correct a3 Shuffled 386 (15.55%) 1,199 (48.31%) 786 (31.67%) 111 (4.47%) 218 (8.07%) 111 (4.47%)
All Correct Answers 1,195 (46.46%) 1,226 (47.67%) 114 (4.43%) 37 (1.44%) 128 (4.74%)
Empty Frames 188 (6.98%) 1,612 (59.81%) 843 (31.28%) 52 (1.93%) 5 (0.19%) 776 (28.79%)
Empty Questions 753 (29.89%) 1,180 (46.84%) 527 (20.92%) 59 (2.34%) 181 (6.70%) 798 (31.68%)
Empty Answers 440 (17.07%) 725 (28.13%) 1,352 (52.46%) 60 (2.33%) 123 (4.56%)
Table 9: Video-LLaVA Performance on All the Settings of NExT-QA (including NExT-GQA). Each cell shows the count of selections for each option (a0a5) along with the percentage in parentheses. Accuracy represents the number of correct answers where applicable.
Setting a0 a1 a2 a3 a4 a5 N/A Accuracy
Target 1,721 (20.10%) 1,671 (19.51%) 1,767 (20.63%) 1,719 (20.07%) 1,686 (19.69%) 0 (0.00%)
Default 874 (10.21%) 3,733 (43.59%) 890 (10.39%) 1,505 (17.57%) 1,562 (18.24%) 0 (0.00%) 4,279 (49.96%)
Answer Shuffling 1,507 (17.60%) 4,366 (50.98%) 1,630 (19.03%) 1,061 (12.39%) 0 (0.00%) 0 (0.00%) 3,653 (42.66%)
Correct Frames 440 (8.87%) 2,194 (44.22%) 482 (9.71%) 791 (15.94%) 1,055 (21.26%) 0 (0.00%) 2,407 (48.51%)
Rephrased Questions 836 (9.79%) 3,531 (41.36%) 1,054 (12.34%) 1,613 (18.89%) 1,504 (17.62%) 26 (0.30%) 4,287 (50.21%)
Additional Empty Option 418 (4.88%) 3,449 (40.27%) 847 (9.89%) 1,208 (14.11%) 2,641 (30.84%) 1 (0.01%) 0 (0.00%) 4,264 (49.79%)
All a0 1,232 (14.39%) 5,839 (68.18%) 9 (0.11%) 175 (2.04%) 1,309 (15.28%) 0 (0.00%)
All a1 1,254 (14.64%) 5,839 (68.18%) 9 (0.11%) 187 (2.18%) 1,275 (14.89%) 0 (0.00%)
All a2 1,238 (14.46%) 5,755 (67.21%) 12 (0.14%) 158 (1.85%) 1,400 (16.35%) 1 (0.01%)
All a3 966 (11.28%) 7,063 (82.47%) 14 (0.16%) 178 (2.08%) 343 (4.01%) 0 (0.00%)
All a4 996 (11.63%) 7,021 (81.98%) 20 (0.23%) 176 (2.06%) 351 (4.10%) 0 (0.00%)
Correct a0 2,966 (34.87%) 3,054 (35.91%) 484 (5.69%) 975 (11.46%) 1,026 (12.06%) 59 (0.69%) 2,966 (34.87%)
Correct a0 Shuffled 2,702 (31.55%) 3,680 (42.97%) 341 (3.98%) 759 (8.86%) 1,082 (12.63%) 0 (0.00%) 2,702 (31.55%)
Correct a1 251 (2.93%) 6,993 (81.66%) 209 (2.44%) 431 (5.03%) 680 (7.94%) 0 (0.00%) 6,993 (81.66%)
Correct a1 Shuffled 270 (3.15%) 6,981 (81.52%) 203 (2.37%) 424 (4.95%) 686 (8.01%) 0 (0.00%) 6,981 (81.52%)
Correct a2 274 (3.20%) 3,619 (42.26%) 3,032 (35.41%) 669 (7.81%) 969 (11.32%) 1 (0.01%) 3,032 (35.41%)
Correct a2 Shuffled 282 (3.29%) 3,561 (41.58%) 3,056 (35.68%) 707 (8.26%) 958 (11.19%) 0 (0.00%) 3,056 (35.68%)
Correct a3 258 (3.01%) 2,765 (32.29%) 319 (3.73%) 4,161 (48.59%) 1,060 (12.38%) 1 (0.01%) 4,161 (48.59%)
Correct a3 Shuffled 278 (3.25%) 2,746 (32.06%) 271 (3.16%) 4,182 (48.83%) 1,087 (12.69%) 0 (0.00%) 4,182 (48.83%)
Correct a4 324 (3.78%) 2,374 (27.72%) 300 (3.50%) 557 (6.50%) 5,009 (58.49%) 0 (0.00%) 5,009 (58.49%)
Correct a4 Shuffled 336 (3.92%) 2,305 (26.91%) 313 (3.65%) 579 (6.76%) 5,031 (58.75%) 0 (0.00%) 5,031 (58.75%)
All Correct Answers 1,536 (17.94%) 6,531 (76.27%) 4 (0.05%) 59 (0.69%) 433 (5.06%) 1 (0.01%)
Empty Frames 634 (7.40%) 4,625 (54.01%) 740 (8.64%) 1,499 (17.50%) 1,066 (12.45%) 0 (0.00%) 3,025 (35.32%)
Empty Questions 1,976 (23.07%) 4,925 (57.51%) 228 (2.66%) 696 (8.13%) 739 (8.63%) 0 (0.00%) 2,979 (34.79%)
Empty Answers 45 (0.53%) 7,590 (88.63%) 12 (0.14%) 484 (5.65%) 433 (5.06%) 0 (0.00%)
Table 10: Video-LLaVA Performance on All the Settings of STAR. Each cell shows the count of selections for each option (a0a4) along with the percentage in parentheses. Accuracy represents the number of correct answers where applicable.
Setting a0 a1 a2 a3 a4 N/A Accuracy
Target 1,755 (24.73%) 1,758 (24.77%) 1,742 (24.54%) 1,843 (25.97%) 0 (0.00%)
Default 909 (12.81%) 4,748 (66.89%) 854 (12.03%) 587 (8.27%) 0 (0.00%) 2,464 (34.71%)
Answer Shuffling 825 (11.62%) 4,688 (66.05%) 675 (9.51%) 910 (12.82%) 0 (0.00%) 2,521 (35.52%)
Correct Frames 1,115 (15.71%) 4,611 (64.96%) 743 (10.47%) 629 (8.86%) 0 (0.00%) 1,881 (26.50%)
Rephrased Questions 759 (10.70%) 4,945 (69.70%) 799 (11.26%) 592 (8.34%) 3 (0.04%) 2,455 (34.60%)
Additional Empty Option 346 (4.87%) 4,472 (63.00%) 545 (7.68%) 1,541 (21.71%) 194 (2.73%) 0 (0.00%) 2,490 (35.08%)
All a0 499 (7.03%) 6,596 (92.93%) 1 (0.01%) 2 (0.03%) 0 (0.00%)
All a1 525 (7.40%) 6,573 (92.60%) 0 (0.00%) 0 (0.00%) 0 (0.00%)
All a2 523 (7.37%) 6,571 (92.58%) 1 (0.01%) 3 (0.04%) 0 (0.00%)
All a3 485 (6.83%) 6,610 (93.12%) 2 (0.03%) 1 (0.01%) 0 (0.00%)
Correct a0 1,438 (20.26%) 4,598 (64.78%) 459 (6.47%) 603 (8.50%) 0 (0.00%) 1,438 (20.26%)
Correct a0 Shuffled 1,424 (20.06%) 4,578 (64.51%) 453 (6.38%) 642 (9.05%) 1 (0.01%) 1,424 (20.06%)
Correct a1 533 (7.51%) 5,621 (79.19%) 426 (6.00%) 518 (7.30%) 0 (0.00%) 5,621 (79.19%)
Correct a1 Shuffled 566 (7.98%) 5,560 (78.34%) 449 (6.33%) 522 (7.36%) 1 (0.01%) 5,560 (78.34%)
Correct a2 1,055 (14.86%) 4,561 (64.26%) 849 (11.96%) 633 (8.92%) 0 (0.00%) 849 (11.96%)
Correct a2 Shuffled 586 (8.26%) 4,486 (63.20%) 1,381 (19.46%) 645 (9.09%) 0 (0.00%) 1,381 (19.46%)
Correct a3 588 (8.28%) 4,387 (61.81%) 458 (6.45%) 1,665 (23.46%) 0 (0.00%) 1,665 (23.46%)
Correct a3 Shuffled 1,061 (14.95%) 4,503 (63.44%) 741 (10.44%) 793 (11.17%) 0 (0.00%) 793 (11.17%)
All Correct Answers 374 (5.27%) 6,710 (94.53%) 1 (0.01%) 13 (0.18%) 0 (0.00%)
Empty Frames 709 (9.99%) 4,772 (67.23%) 1,048 (14.76%) 569 (8.02%) 0 (0.00%) 1,891 (26.64%)
Empty Questions 3,013 (42.45%) 3,748 (52.80%) 194 (2.73%) 143 (2.01%) 0 (0.00%) 2,119 (29.85%)
Empty Answers 63 (0.89%) 7,033 (99.08%) 0 (0.00%) 2 (0.03%) 0 (0.00%)
Table 11: Video-LLaVA Performance on All the Settings of Perception Test. Each cell shows the count of selections for each option (a0a3) along with the percentage in parentheses. Accuracy represents the number of correct answers where applicable.
Setting a0 a1 a2 a3 N/A Accuracy
Target 2,549 (33.29%) 2,561 (33.45%) 2,546 (33.25%) 0 (0.00%)
Default 1,944 (25.43%) 5,174 (67.68%) 527 (6.89%) 11 (0.14%) 3,114 (40.73%)
Answer Shuffling 2,164 (28.28%) 4,887 (63.86%) 602 (7.87%) 3 (0.04%) 3,259 (42.58%)
Rephrased Questions 1,836 (24.04%) 5,253 (68.79%) 547 (7.16%) 20 (0.26%) 3,154 (41.30%)
Additional Empty Option 1,133 (15.01%) 4,548 (60.24%) 1,634 (21.64%) 235 (3.11%) 106 (1.38%) 3,233 (42.82%)
All a0 2,942 (38.43%) 4,586 (59.90%) 128 (1.67%) 0 (0.00%)
All a1 3,001 (39.20%) 4,538 (59.27%) 117 (1.53%) 0 (0.00%)
All a2 3,684 (48.13%) 3,951 (51.61%) 20 (0.26%) 1 (0.01%)
Correct a0 3,086 (40.32%) 4,040 (52.79%) 527 (6.89%) 3 (0.04%) 3,086 (40.32%)
Correct a0 Shuffled 3,121 (40.78%) 4,038 (52.76%) 494 (6.45%) 3 (0.04%) 3,121 (40.78%)
Correct a1 1,524 (19.91%) 5,712 (74.62%) 419 (5.47%) 1 (0.01%) 5,712 (74.62%)
Correct a1 Shuffled 1,542 (20.15%) 5,712 (74.63%) 400 (5.23%) 2 (0.03%) 5,712 (74.63%)
Correct a2 3,085 (40.32%) 4,040 (52.80%) 527 (6.89%) 4 (0.05%) 527 (6.89%)
Correct a2 Shuffled 1,839 (24.03%) 4,944 (64.60%) 870 (11.37%) 3 (0.04%) 870 (11.37%)
All Correct Answers 3,493 (45.64%) 4,078 (53.28%) 83 (1.08%) 2 (0.03%)
Empty Frames 1,920 (25.08%) 5,265 (68.77%) 471 (6.15%) 0 (0.00%) 2,957 (38.62%)
Empty Questions 3,649 (47.66%) 3,767 (49.20%) 240 (3.13%) 0 (0.00%) 3,007 (39.28%)
Empty Answers 1,183 (15.45%) 6,284 (82.09%) 188 (2.46%) 1 (0.01%)
Table 12: Video-LLaVA Performance on All the Settings of Video-MME. Each cell shows the count of selections for each option (a0a4) along with the percentage in parentheses. Accuracy represents the number of correct answers where applicable.
Setting a0 a1 a2 a3 a4 N/A Accuracy
Target 675 (25.00%) 675 (25.00%) 675 (25.00%) 675 (25.00%) 0 (0.00%)
Default 476 (17.63%) 1,693 (62.70%) 353 (13.07%) 178 (6.59%) 0 (0.00%) 924 (34.22%)
Answer Shuffling 427 (15.81%) 1,728 (64.00%) 373 (13.81%) 172 (6.37%) 0 (0.00%) 841 (31.15%)
Rephrased Questions 478 (17.70%) 1,705 (63.15%) 352 (13.04%) 165 (6.11%) 0 (0.00%) 911 (33.74%)
Additional Empty Option 208 (7.84%) 1,833 (69.12%) 191 (7.20%) 420 (15.84%) 0 (0.00%) 48 (1.78%) 909 (34.28%)
All a0 891 (33.00%) 1,730 (64.07%) 27 (1.00%) 52 (1.93%) 0 (0.00%)
All a1 888 (32.90%) 1,733 (64.21%) 33 (1.22%) 45 (1.67%) 1 (0.04%)
All a2 914 (33.86%) 1,703 (63.10%) 27 (1.00%) 55 (2.04%) 1 (0.04%)
All a3 888 (32.90%) 1,657 (61.39%) 32 (1.19%) 122 (4.52%) 1 (0.04%)
Correct a0 660 (24.44%) 1,651 (61.15%) 258 (9.56%) 131 (4.85%) 0 (0.00%) 660 (24.44%)
Correct a0 Shuffled 642 (23.78%) 1,620 (60.00%) 293 (10.85%) 145 (5.37%) 0 (0.00%) 642 (23.78%)
Correct a1 384 (14.22%) 1,974 (73.11%) 241 (8.93%) 101 (3.74%) 0 (0.00%) 1,974 (73.11%)
Correct a1 Shuffled 364 (13.48%) 1,964 (72.74%) 240 (8.89%) 132 (4.89%) 0 (0.00%) 1,964 (72.74%)
Correct a2 386 (14.30%) 1,606 (59.48%) 566 (20.96%) 142 (5.26%) 0 (0.00%) 566 (20.96%)
Correct a2 Shuffled 352 (13.04%) 1,623 (60.11%) 589 (21.81%) 136 (5.04%) 0 (0.00%) 589 (21.81%)
Correct a3 416 (15.41%) 1,642 (60.81%) 311 (11.52%) 331 (12.26%) 0 (0.00%) 331 (12.26%)
Correct a3 Shuffled 357 (13.23%) 1,682 (62.32%) 345 (12.78%) 315 (11.67%) 1 (0.04%) 315 (11.67%)
All Correct Answers 928 (34.38%) 1,702 (63.06%) 29 (1.07%) 40 (1.48%) 1 (0.04%)
Empty Frames 267 (9.89%) 1,828 (67.70%) 437 (16.19%) 168 (6.22%) 0 (0.00%) 764 (28.30%)
Empty Questions 1,545 (57.24%) 960 (35.57%) 119 (4.41%) 75 (2.78%) 1 (0.04%) 790 (29.27%)
Empty Answers 359 (13.32%) 2,082 (77.23%) 165 (6.12%) 90 (3.34%) 4 (0.15%)
Table 13: SeViLA Performance on All the Settings of NExT-QA (including NExT-GQA). Each cell shows the count of selections for each option (a0a5) along with the percentage in parentheses. Accuracy represents the number of correct answers where applicable.
Setting a0 a1 a2 a3 a4 a5 Accuracy
Target 1,721 (20.10%) 1,671 (19.51%) 1,767 (20.63%) 1,719 (20.07%) 1,686 (19.69%)
Default 1,812 (21.16%) 1,606 (18.75%) 1,712 (19.99%) 1,805 (21.08%) 1,629 (19.02%) 5,462 (63.78%)
Answer Shuffling 1,740 (20.32%) 1,598 (18.66%) 1,737 (20.28%) 1,824 (21.30%) 1,665 (19.44%) 5,481 (64.00%)
Correct Frames 1,108 (22.33%) 939 (18.92%) 961 (19.37%) 1,069 (21.54%) 885 (17.84%) 2,932 (59.09%)
Rephrased Questions 1,825 (21.31%) 1,571 (18.34%) 1,657 (19.35%) 1,793 (20.94%) 1,718 (20.06%) 5,160 (60.25%)
Additional Empty Option 1,669 (19.49%) 1,401 (16.36%) 1,457 (17.01%) 1,563 (18.25%) 1,758 (20.53%) 716 (8.36%) 5,054 (59.01%)
All a0 4,972 (58.06%) 381 (4.45%) 631 (7.37%) 1,865 (21.78%) 715 (8.35%)
All a1 4,906 (57.29%) 381 (4.45%) 675 (7.88%) 1,954 (22.82%) 648 (7.57%)
All a2 4,887 (57.06%) 359 (4.19%) 681 (7.95%) 1,968 (22.98%) 669 (7.81%)
All a3 4,961 (57.93%) 337 (3.94%) 601 (7.02%) 1,981 (23.13%) 684 (7.99%)
All a4 4,904 (57.26%) 370 (4.32%) 615 (7.18%) 1,987 (23.20%) 688 (8.03%)
Correct a0 5,434 (63.45%) 732 (8.55%) 761 (8.89%) 844 (9.86%) 793 (9.26%) 5,434 (63.45%)
Correct a0 Shuffled 5,443 (63.56%) 736 (8.59%) 745 (8.70%) 814 (9.50%) 826 (9.65%) 5,443 (63.56%)
Correct a1 870 (10.16%) 5,399 (63.04%) 752 (8.78%) 825 (9.63%) 718 (8.38%) 5,399 (63.04%)
Correct a1 Shuffled 809 (9.45%) 5,379 (62.81%) 740 (8.64%) 866 (10.11%) 770 (8.99%) 5,379 (62.81%)
Correct a2 813 (9.49%) 741 (8.65%) 5,495 (64.16%) 772 (9.01%) 743 (8.68%) 5,495 (64.16%)
Correct a2 Shuffled 777 (9.07%) 734 (8.57%) 5,504 (64.27%) 796 (9.29%) 753 (8.79%) 5,504 (64.27%)
Correct a3 827 (9.66%) 704 (8.22%) 778 (9.08%) 5,591 (65.28%) 664 (7.75%) 5,591 (65.28%)
Correct a3 Shuffled 782 (9.13%) 684 (7.99%) 783 (9.14%) 5,566 (64.99%) 749 (8.75%) 5,566 (64.99%)
Correct a4 860 (10.04%) 730 (8.52%) 708 (8.27%) 853 (9.96%) 5,413 (63.21%) 5,413 (63.21%)
Correct a4 Shuffled 810 (9.46%) 727 (8.49%) 760 (8.87%) 831 (9.70%) 5,436 (63.48%) 5,436 (63.48%)
All Correct Answers 7,300 (85.24%) 209 (2.44%) 237 (2.77%) 448 (5.23%) 370 (4.32%)
Empty Frames 1,833 (21.40%) 1,495 (17.46%) 1,626 (18.99%) 1,820 (21.25%) 1,790 (20.90%) 3,921 (45.78%)
Empty Questions 1,671 (19.51%) 1,513 (17.67%) 1,685 (19.68%) 1,851 (21.61%) 1,844 (21.53%) 4,485 (52.37%)
Empty Answers 2,026 (23.66%) 52 (0.61%) 5 (0.06%) 0 (0.00%) 6,481 (75.68%)
Table 14: SeViLA Performance on All the Settings of STAR. Each cell shows the count of selections for each option (a0a4) along with the percentage in parentheses. Accuracy represents the number of correct answers where applicable.
Setting a0 a1 a2 a3 a4 Accuracy
Target 1,755 (24.73%) 1,758 (24.77%) 1,742 (24.54%) 1,843 (25.97%)
Default 1,501 (21.15%) 1,739 (24.50%) 1,782 (25.11%) 2,076 (29.25%) 3,285 (46.28%)
Answer Shuffling 1,473 (20.75%) 1,713 (24.13%) 1,807 (25.46%) 2,105 (29.66%) 3,260 (45.93%)
Correct Frames 1,526 (21.50%) 1,757 (24.75%) 1,759 (24.78%) 2,056 (28.97%) 3,248 (45.76%)
Rephrased Questions 1,507 (21.23%) 1,736 (24.46%) 1,771 (24.95%) 2,084 (29.36%) 3,153 (44.42%)
Additional Empty Option 1,003 (14.13%) 1,077 (15.17%) 1,214 (17.10%) 1,395 (19.65%) 2,409 (33.94%) 2,351 (33.12%)
All a0 2,951 (41.58%) 510 (7.19%) 480 (6.76%) 3,157 (44.48%)
All a1 2,909 (40.98%) 492 (6.93%) 480 (6.76%) 3,217 (45.32%)
All a2 2,935 (41.35%) 572 (8.06%) 442 (6.23%) 3,149 (44.36%)
All a3 2,892 (40.74%) 550 (7.75%) 472 (6.65%) 3,184 (44.86%)
Correct a0 2,806 (39.53%) 1,342 (18.91%) 1,351 (19.03%) 1,599 (22.53%) 2,806 (39.53%)
Correct a0 Shuffled 2,864 (40.35%) 1,274 (17.95%) 1,312 (18.48%) 1,648 (23.22%) 2,864 (40.35%)
Correct a1 1,149 (16.19%) 3,216 (45.31%) 1,231 (17.34%) 1,502 (21.16%) 3,216 (45.31%)
Correct a1 Shuffled 1,113 (15.68%) 3,181 (44.82%) 1,335 (18.81%) 1,469 (20.70%) 3,181 (44.82%)
Correct a2 1,141 (16.07%) 1,223 (17.23%) 3,253 (45.83%) 1,481 (20.87%) 3,253 (45.83%)
Correct a2 Shuffled 1,090 (15.36%) 1,252 (17.64%) 3,301 (46.51%) 1,455 (20.50%) 3,301 (46.51%)
Correct a3 1,088 (15.33%) 1,215 (17.12%) 1,189 (16.75%) 3,606 (50.80%) 3,606 (50.80%)
Correct a3 Shuffled 1,074 (15.13%) 1,194 (16.82%) 1,242 (17.50%) 3,588 (50.55%) 3,588 (50.55%)
All Correct Answers 3,363 (47.38%) 391 (5.51%) 171 (2.41%) 3,173 (44.70%)
Empty Frames 1,501 (21.15%) 1,739 (24.50%) 1,782 (25.11%) 2,076 (29.25%) 3,285 (46.28%)
Empty Questions 1,563 (22.02%) 1,608 (22.65%) 1,700 (23.95%) 2,227 (31.38%) 3,496 (49.25%)
Empty Answers 1,031 (14.53%) 25 (0.35%) 0 (0.00%) 6,042 (85.12%)
Table 15: SeViLA Performance on All the Settings of Perception Test. Each cell shows the count of selections for each option (a0a3) along with the percentage in parentheses. Accuracy represents the number of correct answers where applicable.
Setting a0 a1 a2 a3 Accuracy
Target 2,549 (33.29%) 2,561 (33.45%) 2,546 (33.25%)
Default 2,256 (29.47%) 2,331 (30.45%) 3,069 (40.09%) 3,468 (45.30%)
Answer Shuffling 2,323 (30.34%) 2,300 (30.04%) 3,033 (39.62%) 3,519 (45.96%)
Rephrased Questions 2,286 (29.86%) 2,231 (29.14%) 3,139 (41.00%) 3,369 (44.00%)
Additional Empty Option 1,674 (21.87%) 1,246 (16.27%) 1,436 (18.76%) 3,300 (43.10%) 2,223 (29.04%)
All a0 3,380 (44.15%) 464 (6.06%) 3,812 (49.79%)
All a1 3,427 (44.76%) 449 (5.86%) 3,780 (49.37%)
All a2 3,377 (44.11%) 449 (5.86%) 3,830 (50.03%)
Correct a0 3,128 (40.86%) 1,978 (25.84%) 2,550 (33.31%) 3,128 (40.86%)
Correct a0 Shuffled 3,112 (40.65%) 2,022 (26.41%) 2,522 (32.94%) 3,112 (40.65%)
Correct a1 1,909 (24.93%) 3,249 (42.44%) 2,498 (32.63%) 3,249 (42.44%)
Correct a1 Shuffled 1,869 (24.41%) 3,268 (42.69%) 2,519 (32.90%) 3,268 (42.69%)
Correct a2 1,858 (24.27%) 1,714 (22.39%) 4,084 (53.34%) 4,084 (53.34%)
Correct a2 Shuffled 1,846 (24.11%) 1,750 (22.86%) 4,060 (53.03%) 4,060 (53.03%)
All Correct Answers 3,531 (46.12%) 353 (4.61%) 3,772 (49.27%)
Empty Frames 2,684 (35.06%) 2,058 (26.88%) 2,914 (38.06%) 2,766 (36.13%)
Empty Questions 2,574 (33.62%) 2,040 (26.65%) 3,042 (39.73%) 3,139 (41.00%)
Empty Answers 985 (12.87%) 0 (0.00%) 6,671 (87.13%)
Table 16: SeViLA Performance on All the Settings of Video-MME. Each cell shows the count of selections for each option (a0a4) along with the percentage in parentheses. Accuracy represents the number of correct answers where applicable.
Setting a0 a1 a2 a3 a4 Accuracy
Target 675 (25.00%) 675 (25.00%) 675 (25.00%) 675 (25.00%)
Default 911 (33.74%) 592 (21.93%) 568 (21.04%) 629 (23.30%) 1,076 (39.85%)
Answer Shuffling 887 (32.85%) 566 (20.96%) 591 (21.89%) 656 (24.30%) 1,041 (38.56%)
Rephrased Questions 947 (35.07%) 556 (20.59%) 552 (20.44%) 645 (23.89%) 1,041 (38.56%)
Additional Empty Option 817 (30.26%) 413 (15.30%) 428 (15.85%) 472 (17.48%) 570 (21.11%) 875 (32.41%)
All a0 1,631 (60.41%) 97 (3.59%) 140 (5.19%) 832 (30.81%)
All a1 1,681 (62.26%) 76 (2.81%) 164 (6.07%) 779 (28.85%)
All a2 1,637 (60.63%) 100 (3.70%) 148 (5.48%) 815 (30.19%)
All a3 1,573 (58.26%) 117 (4.33%) 153 (5.67%) 857 (31.74%)
Correct a0 1,271 (47.07%) 469 (17.37%) 451 (16.70%) 509 (18.85%) 1,271 (47.07%)
Correct a0 Shuffled 1,229 (45.52%) 464 (17.19%) 469 (17.37%) 538 (19.93%) 1,229 (45.52%)
Correct a1 752 (27.85%) 977 (36.19%) 460 (17.04%) 511 (18.93%) 977 (36.19%)
Correct a1 Shuffled 714 (26.44%) 951 (35.22%) 474 (17.56%) 561 (20.78%) 951 (35.22%)
Correct a2 793 (29.37%) 443 (16.41%) 990 (36.67%) 474 (17.56%) 990 (36.67%)
Correct a2 Shuffled 755 (27.96%) 430 (15.93%) 984 (36.44%) 531 (19.67%) 984 (36.44%)
Correct a3 725 (26.85%) 451 (16.70%) 463 (17.15%) 1,061 (39.30%) 1,061 (39.30%)
Correct a3 Shuffled 746 (27.63%) 433 (16.04%) 475 (17.59%) 1,046 (38.74%) 1,046 (38.74%)
All Correct Answers 1,840 (68.15%) 75 (2.78%) 93 (3.44%) 692 (25.63%)
Empty Frames 912 (33.78%) 569 (21.07%) 555 (20.56%) 664 (24.59%) 773 (28.63%)
Empty Questions 876 (32.44%) 523 (19.37%) 571 (21.15%) 730 (27.04%) 1,020 (37.78%)
Empty Answers 727 (26.93%) 29 (1.07%) 37 (1.37%) 1,907 (70.63%)
Algorithm 1 BOLD: Debiasing with Prior Estimation by Decomposition
1:Video-Language model $M$, dataset $\mathcal{D}=\{T_{i}\}$, number of options $n$, estimation budget $K$
2:Model predictions $\mathcal{Y}$
3:Initialize prediction set $\mathcal{Y}\leftarrow\emptyset$ and prior set $\mathcal{P}\leftarrow\emptyset$triangleright Initialization
4:Sample estimation set $\mathcal{D}_{k}$ of size $K$ from $\mathcal{D}$, set $\mathcal{D}_{r}\leftarrow\mathcal{D}\setminus\mathcal{D}_{k}$
5:for $T\in\mathcal{D}_{k}$ do
6:     Apply attacks to get observed predictions:
7:$P_{o}(d_{i}\mid A|_{v=0}(T))$, $P_{o}(d_{i}\mid A|_{q=0}(T))$, $P_{o}(d_{i}\mid A|_{\text{opts}=0}(T))$
8:     Compute decomposed priors using Equation 3:
9:$P_{\text{pri}}(d_{i}\mid A|_{v=0}(T))$, $P_{\text{pri}}(d_{i}\mid A|_{q=0}(T))$, $P_{\text{pri}}(d_{i}\mid A|_{\text{opts}=0}(T))$
10:     Estimate sample-specific prior $\tilde{P}_{\text{prior}}(d_{i}\mid T)$ using Equation 2; add to $\mathcal{P}$
11:end for
12:Estimate global prior $\tilde{P}_{\text{prior}}(d_{i})$ by averaging $\mathcal{P}$triangleright Prior Estimation
13:for $T\in\mathcal{D}$ do
14:     Debias model prediction using the global prior:
15:$P_{\text{deb}}(d_{i}\mid T)=\operatorname{softmax}\left(\log P_{o}(d_{i}\mid T)-\log\tilde{P}_{\text{prior}}(d_{i})\right)$
16:     Add the predicted answer to $\mathcal{Y}$
17:end for
18:return $\mathcal{Y}$
Algorithm 2 Weighted_BOLD: Debiasing with Prior Estimation by Decomposition with Weights
1:Model $M$, dataset $\mathcal{D}=\{T_{i}\}$, estimation budget $K$, number of folds $F=5$
2:Debiased model predictions $\mathcal{Y}$
3:Initialize prediction set $\mathcal{Y}\leftarrow\emptyset$ and prior set $\mathcal{P}\leftarrow\emptyset$triangleright Initialization
4:Sample estimation set $\mathcal{D}_{k}$ of size $K$ from $\mathcal{D}$; set remaining samples $\mathcal{D}_{r}\leftarrow\mathcal{D}\setminus\mathcal{D}_{k}$
5:for each fold $f=1$ to $F$ do $\triangleright$ Cross-validation
6:     Split $\mathcal{D}_{k}$ into test set $\mathcal{D}_{\text{test}}^{f}$ and validation set $\mathcal{D}_{\text{val}}^{f}$
7:     for $T\in\mathcal{D}_{\text{test}}^{f}$ do
8:         Apply attacks to obtain observed predictions: $P_{o}(d_{i}\mid A(T))$
9:         Compute decomposed priors for each attack using Equation 3: $P_{\text{pri}}(d_{i}\mid A(T))$
10:         Estimate sample-specific prior $\tilde{P}_{\text{pri}}(d_{i}\mid T)$ using Equation 5; add to $\mathcal{P}^{f}$
11:     end for
12:     Optimize weights $\{w_{i}\}$ using COBYLA, subject to constraints $0\leq w_{i}\leq 1$ or $|w_{i}|\leq 1$, to minimize recall_std bias on $\mathcal{D}_{\text{test}}^{f}$
13:end for
14:Compute global prior by averaging over folds:
$\tilde{P}_{\text{prior}}(d_{i})=\frac{1}{F}\sum_{f=1}^{F}\tilde{P}_{\text{pri}}^{f}(d_{i})$
$\triangleright$ Global Prior Estimation
15:for $T\in\mathcal{D}$ do $\triangleright$ Debiasing
16:     Debias the model prediction using the global prior:
$P_{\text{deb}}(d_{i}\mid T)=\operatorname{softmax}\left(\log P_{o}(d_{i}\mid T)-\log\tilde{P}_{\text{prior}}(d_{i})\right)$
17:     Add the predicted answer to $\mathcal{Y}$
18:end for
19:return $\mathcal{Y}$
Table 17: Comparison of BOLD and Weighted_BOLD bias mitigation approaches across models and datasets for performance and bias monitoring metrics with $k$ = 0.25. Green arrows indicate improvements: upward for Accuracy and F1_mean, and downward for standard deviation metrics. Red arrows represent deterioration, respectively.
Model Dataset Configuration Performance Metrics Bias Monitoring Metrics
Accuracy F1_mean Recall_std F1_std JS_std
NExT-QA Default 49.95 49.81 16.31 3.27 5.4
BOLD 51.81 (3.72%) 51.71 (3.82%) 13.29 (18.54%) 2.66 (18.59%) 4.58 (15.12%)
+Weighted_BOLD 51.89 (3.88%) 51.77 (3.95%) 13.28 (18.57%) 2.73 (16.57%) 4.56 (15.39%)
-/+Weighted_BOLD 52.28 (4.65%) 52.25 (4.9%) 10.9 (33.18%) 2.23 (31.89%) 3.9 (27.63%)
Video-LLaVA STAR Default 34.77 31.83 24.99 6.82 5.78
BOLD 37.45 (7.7%) 36.14 (13.54%) 17.43 (30.23%) 4.84 (29.02%) 4.52 (21.74%)
+Weighted_BOLD 37.81 (8.75%) 36.67 (15.22%) 16.15 (35.38%) 4.71 (30.94%) 4.15 (28.15%)
-/+Weighted_BOLD 39.87 (14.67%) 39.65 (24.57%) 6.85 (72.56%) 2.34 (65.67%) 2.08 (63.98%)
Perception Test Default 40.72 35.69 27.5 14.75 5.28
BOLD 41.62 (2.22%) 38.68 (8.38%) 21.77 (20.85%) 10.93 (25.88%) 3.77 (28.54%)
+Weighted_BOLD 41.95 (3.02%) 39.46 (10.58%) 19.73 (28.26%) 10.24 (30.53%) 3.39 (35.79%)
-/+Weighted_BOLD 42.01 (3.18%) 40.48 (13.43%) 14.69 (46.59%) 8.48 (42.52%) 2.91 (44.9%)
Video-MME Default 34.3 30.99 24.09 8.06 4.79
BOLD 34.63 (0.97%) 32.71 (5.54%) 18.21 (24.4%) 6.18 (23.36%) 3.86 (19.47%)
+Weighted_BOLD 34.59 (0.86%) 33.01 (6.51%) 15.92 (33.93%) 5.91 (26.68%) 3.38 (29.56%)
-/+Weighted_BOLD 33.93 (1.08%) 33.21 (7.16%) 10.1 (58.07%) 4.09 (49.26%) 2.33 (51.39%)
NExT-QA Default 44.79 40.85 23.83 15.86 16.09
BOLD 45.88 (2.43%) 42.15 (3.18%) 22.98 (3.53%) 15.53 (2.11%) 15.56 (3.3%)
+Weighted_BOLD 45.96 (2.61%) 42.27 (3.46%) 22.78 (4.4%) 15.46 (2.55%) 15.52 (3.55%)
-/+Weighted_BOLD 47.56 (6.19%) 46.63 (14.14%) 13.79 (42.11%) 8.47 (46.62%) 12.14 (24.53%)
Video-LLaMA STAR Default 36.52 31.86 24.1 15.44 15.29
BOLD 37.19 (1.82%) 33.02 (3.65%) 22.29 (7.5%) 14.55 (5.75%) 14.58 (4.6%)
+Weighted_BOLD 37.21 (1.89%) 33.32 (4.59%) 21.45 (10.97%) 14.03 (9.1%) 14.22 (7.0%)
-/+Weighted_BOLD 38.63 (5.76%) 38.07 (19.52%) 10.16 (57.84%) 4.78 (69.05%) 8.93 (41.61%)
Perception Test Default 41.39 37.19 27.06 11.4 16.06
BOLD 41.92 (1.27%) 38.7 (4.06%) 24.04 (11.16%) 9.85 (13.61%) 14.44 (10.12%)
+Weighted_BOLD 42.06 (1.62%) 39.36 (5.86%) 21.8 (19.45%) 9.11 (20.06%) 13.05 (18.75%)
-/+Weighted_BOLD 42.05 (1.59%) 41.35 (11.2%) 11.13 (58.86%) 4.71 (58.65%) 8.43 (47.51%)
Video-MME Default 32.05 28.15 20.62 12.52 13.75
BOLD 32.69 (2.01%) 29.2 (3.73%) 19.28 (6.5%) 11.95 (4.54%) 13.13 (4.53%)
+Weighted_BOLD 32.2 (0.47%) 28.98 (2.94%) 17.65 (14.38%) 11.69 (6.56%) 12.6 (8.39%)
-/+Weighted_BOLD 32.92 (2.72%) 30.39 (7.97%) 15.65 (24.08%) 10.21 (18.39%) 11.68 (15.04%)
NExT-QA Default 63.91 63.88 2.15 1.32 1.99
BOLD 63.92 (0.02%) 63.89 (0.03%) 2.0 (6.81%) 1.18 (10.43%) 1.97 (0.91%)
+Weighted_BOLD 63.88 (0.04%) 63.86 (0.03%) 2.08 (3.47%) 1.28 (2.87%) 1.9 (4.64%)
-/+Weighted_BOLD 63.95 (0.07%) 63.95 (0.11%) 1.44 (32.8%) 1.16 (12.17%) 1.63 (18.3%)
SeViLA STAR Default 46.28 46.14 4.47 2.3 2.25
BOLD 46.17 (0.24%) 46.05 (0.2%) 4.09 (8.42%) 2.24 (2.93%) 2.11 (6.25%)
+Weighted_BOLD 46.07 (0.46%) 45.95 (0.42%) 4.12 (7.63%) 2.28 (1.24%) 2.13 (5.17%)
-/+Weighted_BOLD 46.2 (0.18%) 46.08 (0.13%) 4.04 (9.49%) 2.23 (3.15%) 2.16 (3.9%)
Perception Test Default 45.3 45.11 6.21 3.02 1.59
BOLD 45.28 (0.03%) 45.15 (0.09%) 5.1 (17.85%) 2.93 (3.09%) 1.26 (21.04%)
+Weighted_BOLD 45.34 (0.09%) 45.24 (0.29%) 4.26 (31.39%) 2.65 (12.43%) 1.03 (35.41%)
-/+Weighted_BOLD 45.22 (0.17%) 45.15 (0.09%) 3.69 (40.51%) 2.38 (21.06%) 0.85 (46.67%)
Video-MME Default 39.85 39.82 4.67 1.68 0.84
BOLD 40.04 (0.46%) 40.04 (0.56%) 3.79 (18.89%) 1.48 (12.03%) 0.74 (11.46%)
+Weighted_BOLD 40.07 (0.56%) 40.08 (0.66%) 3.64 (22.05%) 1.48 (11.82%) 0.67 (20.11%)
-/+Weighted_BOLD 40.04 (0.46%) 40.06 (0.63%) 2.31 (50.5%) 1.36 (18.83%) 0.3 (63.83%)
Table 18: Comparison of BOLD and Weighted_BOLD bias mitigation approaches across models and datasets for performance and bias monitoring metrics with $k$ = 0.5. Green arrows indicate improvements: upward for Accuracy and F1_mean, and downward for standard deviation metrics. Red arrows represent deterioration, respectively.
Model Dataset Configuration Performance Metrics Bias Monitoring Metrics
Accuracy F1_mean Recall_std F1_std JS_std
NExT-QA Default 44.79 40.85 23.83 15.86 16.09
BOLD 45.88 (2.43%) 42.15 (3.18%) 22.98 (3.53%) 15.53 (2.11%) 15.55 (3.31%)
+Weighted_BOLD 45.91 (2.51%) 42.2 (3.29%) 22.83 (4.19%) 15.51 (2.2%) 15.56 (3.27%)
-/+Weighted_BOLD 48.25 (7.73%) 46.52 (13.86%) 16.6 (30.32%) 10.6 (33.17%) 13.27 (17.48%)
Video-LLaMA STAR Default 36.52 31.86 24.1 15.44 15.29
BOLD 37.19 (1.82%) 33.02 (3.65%) 22.29 (7.5%) 14.55 (5.75%) 14.58 (4.66%)
+Weighted_BOLD 37.34 (2.24%) 33.5 (5.16%) 21.13 (12.3%) 14.05 (8.98%) 14.14 (7.54%)
-/+Weighted_BOLD 38.61 (5.72%) 38.06 (19.47%) 10.18 (57.75%) 4.79 (68.99%) 8.58 (43.89%)
Perception Test Default 41.39 37.19 27.06 11.4 16.06
BOLD 41.9 (1.24%) 38.68 (4.01%) 24.05 (11.11%) 9.87 (13.45%) 14.45 (10.01%)
+Weighted_BOLD 42.06 (1.62%) 39.36 (5.86%) 21.8 (19.45%) 9.11 (20.06%) 13.02 (18.94%)
-/+Weighted_BOLD 42.27 (2.13%) 41.24 (10.9%) 13.55 (49.93%) 5.64 (50.54%) 9.73 (39.41%)
Video-MME Default 32.05 28.15 20.62 12.52 13.75
BOLD 32.73 (2.13%) 29.23 (3.85%) 19.29 (6.45%) 11.96 (4.46%) 13.1 (4.73%)
+Weighted_BOLD 32.2 (0.47%) 28.98 (2.94%) 17.65 (14.38%) 11.69 (6.56%) 12.64 (8.04%)
-/+Weighted_BOLD 32.88 (2.6%) 31.21 (10.87%) 12.77 (38.06%) 8.02 (35.91%) 10.42 (24.25%)
NExT-QA Default 49.95 49.81 16.31 3.27 5.4
BOLD 51.81 (3.72%) 51.71 (3.82%) 13.27 (18.63%) 2.67 (18.42%) 4.58 (15.06%)
+Weighted_BOLD 52.15 (4.39%) 52.04 (4.49%) 12.72 (22.02%) 2.62 (19.83%) 4.49 (16.83%)
-/+Weighted_BOLD 53.65 (7.41%) 53.53 (7.48%) 8.49 (47.97%) 3.15 (3.61%) 3.78 (29.87%)
Video-LLaVA STAR Default 34.77 31.83 24.99 6.82 5.78
BOLD 37.21 (7.01%) 35.76 (12.37%) 18.28 (26.85%) 4.97 (27.15%) 4.54 (21.39%)
+Weighted_BOLD 37.53 (7.94%) 36.18 (13.69%) 17.45 (30.15%) 4.91 (28.03%) 4.26 (26.24%)
-/+Weighted_BOLD 38.31 (10.17%) 37.39 (17.47%) 13.26 (46.92%) 4.59 (32.59%) 3.25 (43.67%)
Perception Test Default 40.72 35.69 27.5 14.75 5.28
BOLD 41.62 (2.22%) 38.69 (8.4%) 21.75 (20.9%) 10.92 (25.97%) 3.78 (28.4%)
+Weighted_BOLD 41.95 (3.02%) 39.46 (10.58%) 19.73 (28.26%) 10.24 (30.53%) 3.4 (35.51%)
-/+Weighted_BOLD 41.62 (2.22%) 40.88 (14.56%) 11.01 (59.96%) 5.44 (63.11%) 2.59 (50.97%)
Video-MME Default 34.3 30.99 24.09 8.06 4.79
BOLD 34.7 (1.19%) 32.79 (5.81%) 18.19 (24.51%) 6.16 (23.63%) 3.84 (19.93%)
+Weighted_BOLD 34.63 (0.97%) 32.97 (6.38%) 16.36 (32.07%) 6.01 (25.43%) 3.43 (28.5%)
-/+Weighted_BOLD 34.0 (0.86%) 32.6 (5.21%) 14.11 (41.43%) 5.79 (28.18%) 2.92 (39.01%)
NExT-QA Default 63.91 63.88 2.15 1.32 1.99
BOLD 63.92 (0.02%) 63.89 (0.02%) 2.03 (5.47%) 1.18 (10.4%) 1.99 (0.17%)
+Weighted_BOLD 63.93 (0.04%) 63.91 (0.04%) 1.99 (7.64%) 1.19 (9.83%) 1.99 (0.04%)
-/+Weighted_BOLD 64.01 (0.16%) 64.0 (0.19%) 1.14 (47.18%) 1.17 (11.43%) 2.01 (1.1%)
SeViLA STAR Default 46.28 46.14 4.47 2.3 2.25
BOLD 46.22 (0.12%) 46.1 (0.08%) 4.13 (7.46%) 2.26 (1.78%) 2.1 (6.64%)
+Weighted_BOLD 46.2 (0.18%) 46.08 (0.13%) 4.01 (10.17%) 2.2 (4.36%) 2.07 (8.2%)
-/+Weighted_BOLD 46.38 (0.21%) 46.3 (0.34%) 3.41 (23.66%) 1.94 (15.89%) 1.93 (14.33%)
Perception Test Default 45.3 45.11 6.21 3.02 1.59
BOLD 45.32 (0.06%) 45.18 (0.16%) 5.18 (16.61%) 2.95 (2.43%) 1.3 (18.42%)
+Weighted_BOLD 45.31 (0.03%) 45.2 (0.21%) 4.56 (26.58%) 2.83 (6.43%) 1.08 (31.92%)
-/+Weighted_BOLD 45.25 (0.12%) 45.2 (0.2%) 3.25 (47.69%) 2.28 (24.52%) 0.71 (55.13%)
Video-MME Default 39.85 39.82 4.67 1.68 0.84
BOLD 40.19 (0.84%) 40.17 (0.88%) 4.11 (12.03%) 1.41 (16.19%) 0.73 (12.52%)
+Weighted_BOLD 40.04 (0.46%) 40.03 (0.54%) 3.78 (19.14%) 1.44 (14.6%) 0.69 (17.22%)
-/+Weighted_BOLD 39.81 (0.09%) 39.83 (0.04%) 2.99 (35.93%) 1.43 (14.97%) 0.45 (45.74%)
Table 19: Comparison of BOLD and Weighted_BOLD bias mitigation approaches across models and datasets for performance and bias monitoring metrics with $k$ = 0.75. Green arrows indicate improvements: upward for Accuracy and F1_mean, and downward for standard deviation metrics. Red arrows represent deterioration for Accuracy and F1_mean.
Model Dataset Configuration Performance Metrics Bias Monitoring Metrics
Accuracy F1_mean Recall_std F1_std JS_std
NExT-QA Default 44.79 40.85 23.83 15.86 16.09
BOLD 45.87 (2.4%) 42.16 (3.19%) 22.96 (3.63%) 15.49 (2.37%) 15.56 (3.28%)
+Weighted_BOLD 45.91 (2.51%) 42.2 (3.29%) 22.83 (4.19%) 15.51 (2.2%) 15.57 (3.23%)
-/+Weighted_BOLD 47.56 (6.19%) 46.63 (14.14%) 13.36 (43.91%) 8.42 (46.91%) 11.79 (26.73%)
Video-LLaMA STAR Default 36.52 31.86 24.1 15.44 15.29
BOLD 37.19 (1.82%) 33.02 (3.65%) 22.29 (7.5%) 14.55 (5.75%) 14.59 (4.59%)
+Weighted_BOLD 37.34 (2.24%) 33.5 (5.16%) 21.13 (12.3%) 14.05 (8.98%) 14.13 (7.56%)
-/+Weighted_BOLD 38.74 (6.07%) 38.14 (19.74%) 10.58 (56.11%) 4.94 (68.03%) 9.64 (36.94%)
Perception Test Default 41.39 37.19 27.06 11.4 16.06
BOLD 41.9 (1.24%) 38.67 (3.98%) 24.09 (11.0%) 9.88 (13.29%) 14.46 (9.97%)
+Weighted_BOLD 42.06 (1.62%) 39.36 (5.86%) 21.8 (19.45%) 9.11 (20.06%) 13.12 (18.33%)
-/+Weighted_BOLD 42.29 (2.16%) 41.32 (11.13%) 12.96 (52.13%) 5.53 (51.52%) 9.56 (40.45%)
Video-MME Default 32.05 28.15 20.62 12.52 13.75
BOLD 32.69 (2.01%) 29.2 (3.75%) 19.24 (6.7%) 11.94 (4.63%) 13.09 (4.8%)
+Weighted_BOLD 32.2 (0.47%) 28.97 (2.94%) 17.65 (14.38%) 11.7 (6.54%) 12.56 (8.62%)
-/+Weighted_BOLD 32.65 (1.89%) 30.55 (8.55%) 13.94 (32.38%) 9.24 (26.17%) 11.04 (19.73%)
NExT-QA Default 49.95 49.81 16.31 3.27 5.4
BOLD 51.81 (3.72%) 51.71 (3.82%) 13.27 (18.63%) 2.67 (18.42%) 4.58 (15.14%)
+Weighted_BOLD 51.96 (4.02%) 51.86 (4.12%) 12.92 (20.76%) 2.64 (19.44%) 4.48 (17.02%)
-/+Weighted_BOLD 53.75 (7.6%) 53.66 (7.73%) 8.51 (47.82%) 2.58 (21.07%) 3.6 (33.28%)
Video-LLaVA STAR Default 34.77 31.83 24.99 6.82 5.78
BOLD 37.21 (7.01%) 35.76 (12.37%) 18.28 (26.85%) 4.97 (27.15%) 4.62 (20.03%)
+Weighted_BOLD 37.52 (7.9%) 36.17 (13.64%) 17.45 (30.14%) 4.91 (28.02%) 4.27 (26.1%)
-/+Weighted_BOLD 37.86 (8.87%) 36.87 (15.85%) 13.12 (47.51%) 4.68 (31.37%) 3.09 (46.45%)
Perception Test Default 40.72 35.69 27.5 14.75 5.28
BOLD 41.62 (2.22%) 38.68 (8.38%) 21.77 (20.85%) 10.93 (25.88%) 3.79 (28.21%)
+Weighted_BOLD 41.95 (3.02%) 39.46 (10.58%) 19.73 (28.26%) 10.24 (30.53%) 3.41 (35.4%)
-/+Weighted_BOLD 42.21 (3.66%) 41.64 (16.67%) 9.47 (65.56%) 5.14 (65.14%) 1.8 (65.84%)
Video-MME Default 34.3 30.99 24.09 8.06 4.79
BOLD 34.63 (0.97%) 32.64 (5.33%) 18.59 (22.84%) 6.19 (23.2%) 3.85 (19.65%)
+Weighted_BOLD 34.48 (0.54%) 32.89 (6.13%) 15.94 (33.83%) 5.86 (27.32%) 3.37 (29.66%)
-/+Weighted_BOLD 33.78 (1.51%) 33.2 (7.12%) 9.23 (61.7%) 3.35 (58.46%) 1.42 (70.35%)
NExT-QA Default 63.91 63.88 2.15 1.32 1.99
BOLD 63.91 (0.0%) 63.88 (0.01%) 2.03 (5.45%) 1.18 (10.77%) 1.98 (0.34%)
+Weighted_BOLD 63.93 (0.04%) 63.91 (0.04%) 2.0 (6.83%) 1.2 (8.97%) 1.98 (0.23%)
-/+Weighted_BOLD 64.11 (0.31%) 64.09 (0.34%) 1.2 (44.36%) 1.09 (17.41%) 2.15 (8.2%)
SeViLA STAR Default 46.28 46.14 4.47 2.3 2.25
BOLD 46.27 (0.03%) 46.15 (0.02%) 4.07 (8.91%) 2.22 (3.53%) 2.08 (7.72%)
+Weighted_BOLD 46.17 (0.24%) 46.05 (0.19%) 4.01 (10.17%) 2.2 (4.43%) 2.04 (9.14%)
-/+Weighted_BOLD 46.17 (0.24%) 46.1 (0.09%) 2.96 (33.63%) 1.78 (22.67%) 1.64 (27.25%)
Perception Test Default 45.3 45.11 6.21 3.02 1.59
BOLD 45.31 (0.03%) 45.16 (0.12%) 5.25 (15.42%) 2.98 (1.26%) 1.3 (18.44%)
+Weighted_BOLD 45.35 (0.12%) 45.25 (0.32%) 4.27 (31.19%) 2.65 (12.18%) 1.03 (34.92%)
-/+Weighted_BOLD 45.15 (0.32%) 45.09 (0.03%) 3.48 (43.88%) 2.3 (24.02%) 0.78 (51.02%)
Video-MME Default 39.85 39.82 4.67 1.68 0.84
BOLD 40.04 (0.46%) 40.03 (0.53%) 3.84 (17.85%) 1.46 (13.42%) 0.69 (18.0%)
+Weighted_BOLD 39.89 (0.09%) 39.87 (0.15%) 3.7 (20.9%) 1.43 (14.96%) 0.58 (31.13%)
-/+Weighted_BOLD 39.7 (0.37%) 39.71 (0.26%) 2.69 (42.53%) 1.42 (15.75%) 0.15 (81.72%)
Table 20: Comparison of BOLD and Weighted_BOLD bias mitigation approaches across models and datasets for performance and bias monitoring metrics $k$ = 1 (equal to the entire dataset). Green arrows indicate improvements: upward for Accuracy and F1_mean, and downward for standard deviation metrics. Red arrows represent deterioration for Accuracy and F1_mean.
Model Dataset Configuration Performance Metrics Bias Monitoring Metrics
Accuracy F1_mean Recall_std F1_std JS_std
NExT-QA Default 44.79 40.85 23.83 15.86 16.09
BOLD 45.87 (2.4%) 42.16 (3.19%) 22.96 (3.63%) 15.49 (2.37%) 15.56 (3.29%)
+Weighted_BOLD 45.91 (2.51%) 42.2 (3.29%) 22.83 (4.19%) 15.51 (2.2%) 15.57 (3.23%)
-/+Weighted_BOLD 47.53 (6.11%) 46.62 (14.1%) 13.25 (44.39%) 8.46 (46.68%) 11.78 (26.76%)
Video-LLaMA STAR Default 36.52 31.86 24.1 15.44 15.29
BOLD 37.19 (1.82%) 33.02 (3.65%) 22.29 (7.5%) 14.55 (5.75%) 14.59 (4.58%)
+Weighted_BOLD 37.34 (2.24%) 33.5 (5.16%) 21.13 (12.3%) 14.05 (8.98%) 14.13 (7.55%)
-/+Weighted_BOLD 38.72 (6.03%) 38.18 (19.85%) 10.0 (58.49%) 4.85 (68.6%) 8.53 (44.23%)
Perception Test Default 41.39 37.19 27.06 11.4 16.06
BOLD 41.89 (1.21%) 38.67 (3.98%) 24.04 (11.18%) 9.86 (13.52%) 14.45 (10.04%)
+Weighted_BOLD 42.08 (1.65%) 39.38 (5.89%) 21.8 (19.46%) 9.11 (20.03%) 12.93 (19.5%)
-/+Weighted_BOLD 42.25 (2.07%) 41.53 (11.69%) 11.21 (58.59%) 4.76 (58.2%) 8.83 (45.01%)
Video-MME Default 32.05 28.15 20.62 12.52 13.75
BOLD 32.65 (1.89%) 29.13 (3.49%) 19.29 (6.44%) 12.03 (3.86%) 13.09 (4.82%)
+Weighted_BOLD 32.2 (0.47%) 28.97 (2.94%) 17.65 (14.38%) 11.7 (6.54%) 12.57 (8.56%)
-/+Weighted_BOLD 32.92 (2.72%) 31.32 (11.27%) 12.6 (38.89%) 7.85 (37.25%) 10.36 (24.67%)
NExT-QA Default 49.95 49.81 16.31 3.27 5.4
BOLD 51.81 (3.72%) 51.71 (3.82%) 13.27 (18.63%) 2.67 (18.42%) 4.58 (15.18%)
+Weighted_BOLD 51.83 (3.76%) 51.77 (3.93%) 12.79 (21.58%) 2.5 (23.66%) 4.43 (17.81%)
-/+Weighted_BOLD 52.94 (5.98%) 53.0 (6.41%) 8.49 (47.93%) 1.98 (39.48%) 3.02 (43.96%)
Video-LLaVA STAR Default 34.77 31.83 24.99 6.82 5.78
BOLD 37.14 (6.81%) 35.65 (12.02%) 18.53 (25.84%) 5.01 (26.55%) 4.6 (20.38%)
+Weighted_BOLD 37.33 (7.37%) 35.98 (13.04%) 17.62 (29.49%) 4.83 (29.09%) 4.35 (24.73%)
-/+Weighted_BOLD 38.56 (10.9%) 37.86 (18.96%) 11.35 (54.56%) 3.69 (45.92%) 2.91 (49.67%)
Perception Test Default 40.72 35.69 27.5 14.75 5.28
BOLD 41.62 (2.22%) 38.69 (8.4%) 21.75 (20.9%) 10.92 (25.97%) 3.79 (28.14%)
+Weighted_BOLD 41.95 (3.02%) 39.46 (10.58%) 19.73 (28.26%) 10.24 (30.53%) 3.41 (35.37%)
-/+Weighted_BOLD 41.57 (2.09%) 41.11 (15.2%) 8.95 (67.47%) 4.3 (70.81%) 2.11 (59.94%)
Video-MME Default 34.3 30.99 24.09 8.06 4.79
BOLD 34.56 (0.76%) 32.63 (5.3%) 18.24 (24.28%) 6.08 (24.59%) 3.86 (19.41%)
+Weighted_BOLD 34.41 (0.32%) 32.78 (5.77%) 16.1 (33.17%) 5.89 (26.93%) 3.37 (29.69%)
-/+Weighted_BOLD 34.22 (0.22%) 33.56 (8.31%) 9.8 (59.31%) 3.71 (54.0%) 1.84 (61.61%)
NExT-QA Default 63.91 63.88 2.15 1.32 1.99
BOLD 63.91 (0.0%) 63.88 (0.01%) 2.03 (5.45%) 1.18 (10.77%) 1.98 (0.66%)
+Weighted_BOLD 64.0 (0.15%) 63.98 (0.15%) 1.95 (9.2%) 1.24 (6.37%) 1.97 (0.96%)
-/+Weighted_BOLD 64.07 (0.26%) 64.06 (0.29%) 0.84 (60.74%) 1.07 (18.78%) 2.08 (4.64%)
SeViLA STAR Default 46.28 46.14 4.47 2.3 2.25
BOLD 46.25 (0.06%) 46.14 (0.01%) 4.05 (9.25%) 2.22 (3.75%) 2.06 (8.29%)
+Weighted_BOLD 46.15 (0.27%) 46.04 (0.22%) 4.01 (10.23%) 2.2 (4.52%) 2.05 (9.07%)
-/+Weighted_BOLD 46.18 (0.21%) 46.1 (0.1%) 3.48 (22.16%) 2.04 (11.64%) 1.85 (17.64%)
Perception Test Default 45.3 45.11 6.21 3.02 1.59
BOLD 45.4 (0.23%) 45.25 (0.33%) 5.24 (15.53%) 3.04 (0.72%) 1.3 (18.45%)
+Weighted_BOLD 45.34 (0.09%) 45.24 (0.3%) 4.16 (33.06%) 2.62 (13.42%) 1.03 (34.99%)
-/+Weighted_BOLD 45.22 (0.17%) 45.14 (0.08%) 3.85 (38.0%) 2.43 (19.68%) 0.9 (43.13%)
Video-MME Default 39.85 39.82 4.67 1.68 0.84
BOLD 40.0 (0.37%) 39.99 (0.44%) 3.81 (18.5%) 1.48 (11.8%) 0.66 (21.49%)
+Weighted_BOLD 39.81 (0.09%) 39.81 (0.02%) 3.57 (23.49%) 1.45 (13.53%) 0.53 (36.44%)
-/+Weighted_BOLD 39.85 (0.0%) 39.85 (0.08%) 2.66 (43.14%) 1.47 (12.42%) 0.16 (80.65%)